### v1

# Multi-Token Joint Speculative Decoding for Accelerating Large Language Model Inference

Abstract
Transformer-based Large language models (LLMs) have demonstrated their power in various tasks, but their inference incurs significant time and energy costs. To accelerate LLM inference, speculative decoding uses a smaller model to propose one sequence of tokens, which are subsequently validated in batch by the target large model. Compared with autoregressive decoding, speculative decoding generates the same number of tokens with fewer runs of the large model, hence accelerating the overall inference by $1$-2×2$\times$2 ×. However, greedy decoding is not the optimal decoding algorithm in terms of output perplexity, which is a direct measurement of the effectiveness of a decoding algorithm. An algorithm that has better output perplexity and even better efficiency than speculative decoding can be more useful in practice. To achieve this seemingly contradictory goal, we first introduce multi-token joint greedy decoding (MJGD), which greedily generates multiple tokens at each step based on their joint perplexity. We show that it leads to better perplexity for the whole output. But the computation cost of MJGD is infeasible in practice. So we further propose multi-token joint speculative decoding (MJSD), which approximates and accelerates the MJGD from two aspects: it approximates the joint distribution of the large model with that of a small model, and uses a verification step to guarantee the accuracy of approximation; then it uses beam decoding to accelerate the sequence generation from the joint distribution. Compared with vanilla speculative decoding, MJSD has two advantages: (1) it is an approximation of MJGD, thus achieving better output perplexity; (2) verification with joint likelihood allows it to accept the longest prefix sub-sequence of the draft tokens with valid perplexity, leading to better efficiency. In addition, we analyze energy consumption during LLM inference and provide theoretical and empirical evidence that, surprisingly, MJSD reduces energy consumption even when the number of operations increases. Experiment results show that our approach enhances perplexity by $21.2\%$ than greedy decoding. Moreover, MJSD achieves 2.21×2.21$\times$2.21 × speed-up and 2.84×2.84$\times$2.84 × less energy consumption than greedy decoding, and 1.49×1.49$\times$1.49 × speed-up and 1.62×1.62$\times$1.62 × less energy consumption than vanilla speculative decoding.
1 Introduction
Large Language Models (LLMs) have exhibited remarkable performance across real-world tasks spanning text and image domains [4, 7, 33, 34]. However, the substantial parameter size of these models leads to high computational cost during inference [27]. For example, ChatGPT is estimated to have an energy demand of 564 MWh per day [8]. Compounding this issue, the autoregressive generation of $K$ tokens necessitates calling the model $K$ times, involving the repeated loading of weight matrices and intermediate results from GPU global memory to computing units. As a result, LLM inference is often hampered by constraints related to memory bandwidth and communication [20], leading to diminished hardware utilization, heightened inference latency, and elevated energy cost.
To tackle this challenge, researchers have delved into non-autoregressive decoding approaches. Earlier methods [11, 12, 13] aimed at reducing inference latency by concurrently generating multiple tokens. But these methods usually require task-dependent techniques and information to achieve comparable performance to that of autoregressive decoding [17, 36]. More recently, speculative decoding has emerged [20, 5, 17, 32], exploiting the observation that, even when a small model is ten times smaller than a large model, most of the small model’s prediction aligns well with that of a large model. So it leverages a smaller auxiliary model to draft a few future tokens autoregressively, which are subsequently validated in parallel by the larger model. As the smaller model operates significantly faster and parallel token verification incurs a similar time cost as generating a single token, speculative decoding achieves an overall speed-up of $1$-2×2$\times$2 ×. Meanwhile, it always has an identical sampling distribution as greedy decoding.
However, there are some limitations of the existing studies on speculative decoding. First, they aim to have an identical sampling distribution as greedy sampling. But the goal of decoding algorithms is to generate sequences with optimal perplexity, which is defined as the exponentiated average negative log-likelihood of all tokens. An algorithm that does not follow the same sampling distribution as greedy decoding, but yields output with better perplexity could be more useful in the real world. Second, in these algorithms, if a draft token is rejected, all the draft tokens after it are discarded. But the overall draft sequences may have good quality despite there being a few low-quality tokens. Rejecting these sequences will shorten the average acceptance length for each iteration, lowering the decoding efficiency. Third, existing research does not investigate the impact of speculative decoding on inference energy consumption. While speculative decoding increases the number of FLOPs during inference, it simultaneously reduces the overall inference time, communication, and memory operations, which are crucial factors in determining energy consumption [1, 6]. Thus, it remains unclear if speculative decoding increases or decreases the energy cost of inference.
So we aim to design a new algorithm that is both more efficient and more effective than greedy decoding and vanilla speculative decoding. We first introduce multi-token joint greedy decoding (MJGD), which generates multiple tokens at each step greedily based on their joint likelihood. We empirically show that it has better overall perplexity than greedy decoding does. However, the computational cost of MJGD is infeasible in practice. So we further propose multi-token joint speculative decoding (MJSD), which approximates and accelerates MJGD from two aspects: (1) it approximates the joint distribution of the large model with that of a small model by generating draft tokens with the small model and validating them with the large model; (2) it uses beam decoding to accelerate the sequence generation from the joint distribution of the small model. Compared with vanilla speculative decoding, MJSD has two advantages: (1) it is more effective, as it approximates MJGD, which has better output perplexity than greedy decoding does; (2) it is more efficient because it uses joint likelihood as the verification criterion, which allows it to accept the longest prefix sub-sequence of the draft tokens with valid perplexity. In addition, we analyze the energy consumption of LLM inference. We give theoretical and empirical evidence that, despite that MJSD and other speculative decoding algorithms increase the number of FLOPs needed during LLM inference, they reduce the overall energy consumption by reducing the overhead induced by accessing GPU global memory.
We evaluate our method on text generation tasks with Llama-2 [34] and OPT models [38]. Experiment results show that our approach enhances perplexity by $21.2\%$ than the baselines with identical sampling distributions as greedy decoding. In addition, MJSD achieves 2.21×2.21$\times$2.21 × speed-up and 2.84×2.84$\times$2.84 × smaller energy consumption than greedy decoding, and 1.49×1.49$\times$1.49 × speed-up and 1.62×1.62$\times$1.62 × smaller energy consumption than vanilla speculative decoding. Our code is open-sourced [^1] https://anonymous.4open.science/r/LLMSpeculativeSampling-EE52.
2 Preliminaries
2.1 Decodings of LLMs
Let $p$ denote the distribution defined by LLM model $M_{p}$. Given an input prefix $prefix$, the optimal decoding algorithm is to generate a sequence of $N$ tokens with maximum likelihood $p(x_{1:N}|prefix)$. The likelihood of the sequence is directly linked to perplexity of the sequence, which is the exponentiated average negative log-likelihood of all tokens. Based on autoregressive decomposition $p(x_{1:N}|prefix)=\prod_{t=1}^{N}p(x_{t}|x_{1:t-1},prefix)$, the perplexity is defined as:
$PPL(x_{1:N})=\exp\left\{-\frac{1}{N}\sum_{t=1}^{N}\log p(x_{t}|x_{1:t-1})\right\}$ (1)
Now we introduce commonly used decoding approaches.
Greedy Decoding.
Greedy decoding samples the next single token $x_{t}$ based on $p(\cdot|x_{1:t-1},prefix)$. According to the autoregressive decomposition, greedy sampling leads to the desired joint distribution for the sequence. In practice, however, the objective of decoding algorithms is to generate sequences with high likelihood. Instead of sampling from the joint distribution, a warping operation $\mathcal{T}$ usually is applied to boost the high probability region, which transforms the original distribution $p$ to $\tilde{p}=\mathcal{T}\circ p$. A very popular warping operation is $argmax$, which only selects the sample with the highest probability. Other possible warping operations include top-k. For most of the warping operations, autoregressive decomposition no longer holds after the warping. For example, $\arg\max_{x_{1:N}}{p}(x_{1:N}|prefix)\neq\prod_{t=1}^{N}\arg\max_{x_{t}}{p}(x_{t}|x_{1:t-1},prefix)$.
Beam Decoding.
Beam decoding aims to do a better job than greedy decoding. For each position $t$ ($1\leq t\leq N$), it maintains $M>1$ candidate sequences, which are also called beams. Still take $argmax$ warping operator as an example and assume we have already kept the $M$ sequences $\mathcal{I}_{t-1}=\{x_{1:t-1}^{(1)},\ldots,x_{1:t-1}^{(M)}\}$ at position $t-1$ with highest likelihood, $M$ tokens are then sampled from $p(\cdot|x_{1:t-1}^{(m)},prefix)$ for each sequence $m$ and the top $M$ sequences with the highest likelihood $p(x_{1:t}|prefix)$ will be kept. It is widely recognized that beam decoding is a closer approximation to $\tilde{p}(x_{1:N}|prefix)$ than greedy decoding [18], with few exceptions under special construction. The computation for each beam can be parallelized, thus its run time is about the same as greedy decoding. The fact that beam decoding is more effective than greedy decoding also suggests that the output perplexity of greedy decoding still has room for improvement.
2.2 Vanilla Speculative Decoding
Besides effectiveness, speculative decoding is proposed by [20, 5] to accelerate the inference of LLMs. It utilizes a small model to generate the next $\gamma$ tokens and then uses the large model to verify the drafted tokens in parallel, which is summarized below:
  1. 1.
    Let $prefix$ be the input context, the small model samples $\gamma$ draft tokens $x_{1},\ldots,x_{\gamma}$ using greedy decoding based on the warped predicted conditional probability $\tilde{q}(x_{t}|x_{1:t-1},prefix))$ for $t=1,\ldots,\gamma$, where $\tilde{q}=\mathcal{T}\circ q$ and $q$ is the small model’s output distribution.
  2. 2.
    The large model verifies the draft tokens in parallel by computing the conditional probability $\tilde{p}(x_{t}|x_{1:t-1},prefix)$ for $t=1,\ldots,\gamma$.
  3. 3.
    Each draft token $x_{t}$ is accepted with probability $\min(1,\tilde{p}(x_{t})/\tilde{q}(x_{t}))$. The draft tokens before the first rejected token are kept as the decoding output. An additional token is sampled from a residual distribution as a correction to the first rejected token. Then the accepted tokens and the resampled token are appended to the context $prefix$ as the input to the next iteration.
  4. 4.
    Repeat step 1-3 until reaching the stopping criteria, e.g., reaching the length limit..
Because the large model verifies $\gamma$ tokens in parallel with one run, the time cost is smaller than calling it $\gamma$ times. Meanwhile, although the small model still runs in an autoregressive way, its inference speed is much faster than the large model. So speculative decoding can accelerate the inference process of LLMs. Additionally, it is proven that each token $x_{i}$ generated by speculative sampling follows the identical sampling distribution $\tilde{p}(x_{t}|x_{1:t-1},prefix)$ as greedy decoding [20]. However, since $\tilde{p}(x_{1:N}|prefix))\neq\prod_{t=1}^{N}\tilde{p}(x_{t}|x_{1:t-1},prefix)$, the effectiveness of speculative decoding also has room for improvement. Therefore, instead of having an identical sampling distributon as greedy decoding, our algorithm aims to achieve a better approximation to $\tilde{p}(x_{1:N}|prefix)$.
3 Methodology
The perplexity is a direct measurement of the effectiveness of a decoding algorithm. In practice, when the model is well-trained, a better perplexity usually leads to better downstream effectiveness. Although it is possible that improving perplexity does not improve the performance on downstream tasks, the problem is due to the model itself instead of the decoding algorithm. Therefore, the goal of this work is to design an algorithm that yields better output perplexity and better efficiency than greedy decoding and vanilla speculative decoding.
In this section, we first introduce multi-token joint greedy decoding (MJGD) to give a motivating example for why generating multiple tokens based on their joint likelihood can lead to better perplexity. Then we introduce multi-token joint speculative decoding (MJSD), which approximates and accelerates MJGD by exploiting a small model.
3.1 Multi-Token Joint Greedy Decoding
Let $M_{p}$ be the large target model, the goal is to generate a sequence of $N$ tokens with maximum likelihood $p(x_{1:N}|prefix)$. We now introduce a new decoding algorithm to improve greedy decoding in terms of perplexity.
Definition 3.1.
Multi-Token Joint Greedy Decoding. Let $M_{p}$ be the large target model with distribution $p$. Different from single-token greedy decoding, multi-token joint greedy decoding (MJGD) generates the next $\gamma_{i}$ tokens at step $i$ based on their joint conditional probability $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^2] In the paper, we omit $prefix$ when there is no ambiguity., where $\gamma_{i}$ is an integer no less than 1 and $t=\sum_{i^{\prime}=1}^{i-1}{\gamma_{i^{\prime}}}$, i.e., the total tokens generated in the previous $i-1$ steps.
Figure 1: Perplexity and rougeL score of the output when $\gamma_{i}=K$ for MJGD with OPT-125M and Llama-2-68M fined-tuned on ChatGPT-Prompts [26] dataset.
Greedy decoding is a special case of MJGD where $\gamma_{i}=1$, $\forall i$. When $\gamma_{1}=N$, MJGD generates the optimal sequence by returning the sequence with the highest likelihood. So intuitively, output perplexity should improve as $\gamma_{i}$ increases. Additionally, generating $\gamma_{i}$ tokens simultaneously allows MJGD to return the optimal $\gamma_{i}$ tokens for step $i$. In contrast, vanilla greedy decoding selects each token without considering any future tokens. So MJGD is less prone to choosing local optima than greedy decoding.
Empirical evidence supports our claim. Figure 1 shows the output perplexity and downstream ROUGE-L scores of MJGD with $\gamma_{i}$ set to a constant $K$, where $K=1,\ldots,5$. Notice that setting $K=1$ is equivalent to vanilla greedy decoding. We use beam decoding to approximate the $\arg\max$ sampling from the joint distribution $p(x_{t+1:t+K}|x+{1:t},prefix)$. We can see that when $K$ increases from 1 to 2, there is a significant improvement in perplexity. The perplexity continues to decrease when $K$ increases further. It confirms our claim that increasing $\gamma_{i}$ will increase the output perplexity. So $\prod_{i=1}^{n}\tilde{p}(x_{\Gamma_{i-1}+1:\Gamma_{i}}|prefix,x_{1:\Gamma_{i-1}})$ is a better approximation to $\tilde{p}(x_{1:N}|prefix)$ than $\prod_{t=1}^{N}\tilde{p}(x_{t}|prefix,x_{1:t-1})$. Here $\Gamma_{0}=0$, and $\Gamma_{i}=\sum_{j=1}^{i}\gamma_{i}$ for $i\geq 1$, and $n$ is the total number of steps of MJGD.
Moreover, the ROUGE-L score also improves with $K$, supporting our claim that better perplexity reflects enhanced performance in downstream tasks. This demonstrates that MJGD not only achieves lower perplexity but also yields higher quality text generation, making it a superior decoding strategy compared to vanilla greedy decoding [^3] Additional discussions comparing MJGD and greedy decoding are provided in the Appendix F..
3.2 Multi-Token Joint Speculative Decoding
Unfortunately, the computation cost of MJGD is infeasible in practice, since the time and space complexity to compute the joint distribution of $\gamma_{i}$ tokens is $|V|^{\gamma_{i}}$. Inspired by speculative decoding and the facts that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s prediction deviate from those of the large model” [20, 17], we propose multi-token joint speculative decoding (MJSD), which exploits a small model $M_{q}$ to accelerate MJGD approximately. The core idea is to (1) use the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$ output by $M_{q}$ to approximate $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^4] It is also valid to approximate $\tilde{p}$ with $\tilde{q}$. Without loss of generality, we consider non-warped distribution in the illustration of MJSD. and generate $\gamma$ draft tokens from $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$, then (2) use the large model to validate draft tokens in parallel and accept the longest draft sub-sequence that passes verification, and (3) sample an additional token from the distribution of the large model without extra overhead to ensure at least one token is generated at each iteration. However, it is still infeasible to directly generate draft tokens from the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$. So we propose to further approximate this process with beam decoding, which is an effective and efficient algorithm to generate sequences with high likelihood. In this way, MJSD reduces the number of runs of the large model to generate $N$ tokens, thus accelerating the inference in the same way as vanilla speculative decoding does. Algorithm 1 in the Appendix illustrates the pseudocode of MJSD algorithm.
Figure 2: An example of MJSD’s verification process. MJSD accepts the longest draft sub-sequence that passes verification based on joint likelihood.
Draft Tokens Verification
Figure 2 illustrates the verification process of MJSD. Let $x_{t+1},\ldots,x_{t+\gamma}$ be the draft tokens generated by beam decoding with the small model. Since beam decoding is a widely recognized algorithm to generate sequences with high overall likelihood [18], it is reasonable to assume $q(x_{t+1:t+\gamma}|x_{1:t})$ is large. Also, since beam decoding works in an autoregressive way, we can also assume that $\forall j\in\{1,\ldots,\gamma\}$, $q(x_{t+1:t+j}|x_{1:t})$ is large. To approximate MJGD, for each step $i$, MJSD needs to ensure the accepted tokens $x_{t+1:t+\gamma_{i}}$ ($0\leq\gamma_{i}\leq\gamma$) also have high joint likelihood with the large model $M_{p}$. So MJSD first computes the joint likelihood $p(x_{t+1:t+j}|x_{1:t})$ for $j=1,\ldots,\gamma$. Then for each prefix sub-sequence $x_{t+1:t+j}$, it passes verification if and only if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, where $\tau\in[0,1)$ is a pre-defined threshold. Notice that if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, we have $\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})}>\tau$, which means $\frac{q(x_{t+1:t+j}|x_{1:t})-p(x_{t+1:t+j}|x_{1:t})}{p(x_{t+1:t+j}|x_{1:t})}<\frac{1}{\tau}-1$. Therefore, our acceptance policy guarantees that when $q(x_{t+1:t+j}|x_{1:t})>p(x_{t+1:t+j}|x_{1:t})$, the relative error is bounded. And if $q(x_{t+1:t+j}|x_{1:t})\leq p(x_{t+1:t+j}|x_{1:t})$, it means the sub-sequence has higher likelihood in the large model, then it is reasonable to accept it. After verifying all the sub-sequences, MJSD accepts the longest draft sub-sequence that passes verification.
The verification step of MJSD ensures that the accepted tokens have a high joint likelihood with the large model. We have shown that selecting multiple tokens based on their joint likelihood lead to better output perplexity. Thus, MJSD is more effective than greedy decoding and vanilla speculative decoding. Furthermore, since MJSD accepts the longest draft sub-sequence with high likelihood, it can tolerate low-quality tokens as long as the joint likelihood is high. So at each iteration, MJSD can accept more draft tokens than vanilla speculative decoding, which results in better efficiency.
4 Energy Efficiency Analysis
Previous studies [20, 5, 17, 32] only focus on the speed of speculative decoding. However, an equally important consideration is energy consumption. To our knowledge, there is no existing work evaluating the impact of speculative decoding on inference energy consumption. Although MJSD and speculative decoding increase the number of FLOPs due to the involvement of an additional small model and the rollback operation, they concurrently reduce the inference time and memory operations, which are key factors of GPU energy consumption [1, 6]. Consequently, it poses an open question regarding whether speculative decoding increases or decreases overall energy consumption.
To understand the net effect of speculative decoding, we decompose the total energy consumption into two parts following [1]:
$E_{total}=PW_{flop}T_{flop}+PW_{mem}T_{mem}$ (2)
where $PW_{flop},PW_{mem}$ denote the power (energy/second) of FLOPs and memory operations, and $T_{flop},T_{mem}$ denote the time spent on these operations. When input batch size increases, $PW_{flop}$ increases until it reaches the power of maximum FLOPs, denoted as $PW^{*}_{flop}$. Meanwhile, $PW_{mem}$ is irrelevant to the input batch size, as it only depends on the memory hardware.
To determine the relative magnitude relationship between $PW_{flop}$ and $PW_{mem}$, we first point out the fact that GPU memory operations in LLM inference are dominated by accessing off-chip global memory, which consumes about 100×100$\times$100 × of energy compared to accessing on-chip shared memory [16]. It is because each multiprocessor on GPU usually has 64KB of on-chip memory shared by multiple threads, while storing a single layer of LLM, say T5-11b [25], requires about 1GB memory. Moreover, Allen and Ge showed that doing sequential read from off-chip memory consumes $20$-$30\%$ more power than running maximum FLOPs [1]. So we have $PW_{mem}>PW^{*}_{flop}\geq PW_{flop}$. Notice that $PW^{*}_{flop}=PW_{flop}$ only if the batch size reaches the maximum parallelization capacity of GPUs. During greedy decoding and speculative decoding, the batch size is usually small [20]. So most of the computing power is not utilized [20], which means $PW_{mem}\gg PW_{flop}$.
In addition, previous studies have shown that during LLM inference $T_{mem}\gg T_{flop}$ [20]. Therefore, the energy induced by memory operations, i.e., $PW_{mem}T_{mem}$ dominates $E_{total}$. Since speculative decoding reduces $T_{mem}$ by reducing the number of runs of the large model, it should reduce the inference energy consumption to a similar extent as it reduces time consumption.
To validate our hypothesis, we conducted an experiment to evaluate how batch size influences energy consumption during inference. We ran OPT-13b models on 2 Nvidia A10 GPUs with 24GB memory. Fixing the total number of runs of the large model while varying the input batch size $b\in\{1,2,4,8,16\}$ for each run, we measured time and energy cost. The details of energy measurement are illustrated in the Appendix B. Table 1 shows the results. As batch size doubles, although the number of FLOPs doubles, the energy consumption per run increases slightly. This observation demonstrates that $PW_{mem}T_{mem}$ dominates $E_{total}$. In the next section, we will show more direct evidence for the energy efficiency of MJSD and other speculative decoding algorithms.
Table 1: The effect of batch size to inference speed and energy consumption. The number of inputs is the product of the number of LLM runs and input batch size.
Batch Size Energy (J) Energy/run (J) Energy/Input (J) Time (s) Time/run (s) Time/input(s)
1 42,450 14.1 14.1 1,129 0.376 0.376
2 49,621 16.5 8.26 1,191 0.397 0.198
4 53,325 17.7 4.43 1,178 0.392 0.098
8 59,210 19.7 2.46 1,211 0.403 0.050
16 74,058 24.7 1.54 1,255 0.418 0.026
5 Experiments
Datasets and Models.We use three public datasets for evaluation: (1) ChatGPT-Prompt [26], (2) ChatAlpaca [3], and (3) CNN Dailymail [30]. The three datasets represent three common tasks for LLMs: instruction, multi-turn chat, and summarization. Table 4 in the Appendix shows more details of the datasets. We use two pairs of small and large models in our experiments: OPT-125m and OPT-13B [38], and Llama-68M [24] and Llama-2-13B [34].
Baselines. For each pair of small and large models, we compare our method with greedy decoding (greedy), and three speculative decoding methods: vanilla speculative decoding (speculative[19, 5], Spectr [32], and BiLD [17]. Our implementation of MJSD and all the baselines are based on a public implementation of speculative decoding [2]. We also tried to compare with SpecInfer [24] using their released implementation. But it runs out of memory on our machine. For each method, we let it generate at most 128 tokens for each input and run it for $1,000$ seconds. We open-sourced our code for reproduction.
More details of the hyper-parameters, warping operations, and machine configurations of the experiments can be found in the Appendix B, C, and D .
Table 2: Inference efficiency and output perplexity of different methods on ChatGPT-Prompt (CP), ChatAlpaca (CA), and CNNDailyMail (CD) datasets. Bold numbers mark the best result, underlined numbers mark the second best.
[…]
speed up $\uparrow$ 1.00$\pm$0.00 1.63$\pm$0.02 1.52$\pm$0.04 1.99$\pm$0.06 2.78$\pm$0.01
energy (J/token) $\downarrow$ 11.7$\pm$1.56 6.62$\pm$0.91 7.45$\pm$0.90 5.17$\pm$0.88 3.38$\pm$0.02
perplexity $\downarrow$ 3.74$\pm$0.14 3.64$\pm$0.11 3.15$\pm$0.06 3.64$\pm$0.08 2.06$\pm$0.06
OPT speed (token/s) $\uparrow$ 22.4$\pm$0.48 33.8$\pm$2.47 31.5$\pm$1.87 38.0$\pm$2.20 55.8$\pm$0.30
speed $\uparrow$ 1.00$\pm$0.00 1.51$\pm$0.08 1.41$\pm$0.06 1.70$\pm$0.06 2.50$\pm$0.05
energy (J/token) $\downarrow$ 13.2$\pm$0.28 7.48$\pm$0.07 8.75$\pm$0.13 6.08$\pm$0.11 3.61$\pm$0.03
perplexity $\downarrow$ 5.49$\pm$0.15 5.47$\pm$0.11 4.51$\pm$0.09 5.27$\pm$0.09 3.00$\pm$0.09
CA Llama-2 speed (token/s) $\uparrow$ 22.0$\pm$0.15 31.6$\pm$0.35 28.8$\pm$0.20 27.7$\pm$0.29 44.1$\pm$0.25
speed up $\uparrow$ 1.00$\pm$0.00 1.43$\pm$0.02 1.31$\pm$0.01 1.26$\pm$0.00 2.00$\pm$0.02
energy (J/token) $\downarrow$ 11.2$\pm$0.16 6.98$\pm$0.15 7.99$\pm$0.15 7.20$\pm$0.08 4.72$\pm$0.03
perplexity $\downarrow$ 2.11$\pm$0.01 2.13$\pm$0.03 1.95$\pm$0.03 2.15$\pm$0.01 1.88$\pm$0.05
OPT speed (token/s) $\uparrow$ 23.8$\pm$0.10 35.6$\pm$0.45 38.5$\pm$0.93 28.4$\pm$0.34 49.6$\pm$0.42
speed $\uparrow$ 1.00$\pm$0.00 1.49$\pm$0.01 1.62$\pm$0.03 1.19$\pm$0.01 2.08$\pm$0.03
energy (J/token) $\downarrow$ 11.3$\pm$0.22 5.74$\pm$0.11 5.12$\pm$0.06 6.24$\pm$0.11 4.03$\pm$0.02
perplexity $\downarrow$ 3.28$\pm$0.06 3.32$\pm$0.10 2.60$\pm$0.06 3.16$\pm$0.06 2.07$\pm$0.03
CD Llama-2 speed (token/s) $\uparrow$ 21.7$\pm$0.08 30.7$\pm$0.18 30.5$\pm$0.21 25.0$\pm$0.31 44.2$\pm$0.99
speed up $\uparrow$ 1.00$\pm$0.00 1.41$\pm$0.00 1.41$\pm$0.01 1.15$\pm$0.01 2.04$\pm$0.05
energy (J/token) $\downarrow$ 11.3$\pm$0.21 7.07$\pm$0.19 7.41$\pm$0.16 8.22$\pm$0.19 4.80$\pm$0.12
perplexity $\downarrow$ 2.88$\pm$0.04 2.87$\pm$0.08 2.93$\pm$0.03 3.06$\pm$0.11 2.63$\pm$0.10
OPT speed (token/s) $\uparrow$ 23.3$\pm$0.81 31.7$\pm$0.91 30.9$\pm$0.80 23.7$\pm$0.40 43.6$\pm$0.33
speed $\uparrow$ 1.00$\pm$0.00 1.36$\pm$0.04 1.32$\pm$0.04 1.01$\pm$0.02 1.87$\pm$0.02
energy (J/token) $\downarrow$ 11.5$\pm$0.14 6.37$\pm$0.11 6.71$\pm$0.17 7.31$\pm$0.17 4.86$\pm$0.03
perplexity $\downarrow$ 3.93$\pm$0.14 3.97$\pm$0.06 3.74$\pm$0.09 4.04$\pm$0.07 3.17$\pm$0.06
[…]
5.1 Comparison with Baselines
Table 2 shows the main results of our experiments. The standard deviations in the table are computed by repeating the experiment four times. First, we compare the output perplexity of different algorithms. We can see that the perplexity of greedy decoding, vanilla speculative decoding, and Spectr decoding are close because they have identical sampling distributions for each token. Meanwhile, BiLD approximates the sampling distribution of greedy decoding but yields better perplexity. It is because we set a strict acceptance threshold for BiLD decoding, which lowers the acceptance rate but ensures every token has a high probability in the large model. It can be viewed as having an additional warping operation towards the sampling distribution. Nevertheless, there is a significant gap between the perplexity of MJSD and other baselines. On average, the output perplexity of MJSD is $21.2\%$ lower than that of BiLD. It provides strong evidence that MJSD has better effectiveness than existing speculative decoding methods that aim to have identical sampling distribution as greedy decoding.
Next, we compare the speeds of different algorithms. While the speeds of all speculative decoding baselines are close, MJSD is significantly faster. On average, MJSD is 2.21×2.21$\times$2.21 × faster than greedy decoding, 1.49×1.49$\times$1.49 × faster than vanilla speculative decoding, 1.55×1.$5$ faster than BiLD, and 1.64×1.64$\times$1.64 × faster than Spectr. The speed improvement is because MJSD has a larger average acceptance length at each iteration, as MJSD accepts the longest sub-sequence that passes verification. Figure 3 shows the average acceptance length of vanilla speculative decoding and MJSD. We can see that the acceptance length of MJSD is on average 2.15×2.15$\times$2.15 × larger than that of vanilla speculative decoding.
Finally, we focus on energy efficiency. We can see that all speculative decoding algorithms have significantly better energy efficiency over greedy decoding. It supports our analysis that speculative decoding can improve the LLM inference energy efficiency by reducing the number of memory operations. In addition, we can see that the energy efficiency of MJSD is significantly better than that of all baseline algorithms. We believe it is also because of the longer acceptance sequence at each iteration. The number of memory operations is proportional to the number of LLM runs, which is inversely proportional to the average acceptance length. Since by always accepting the longest prefix that passes verification, the energy consumption of MJSD is on average 1.62×1.62$\times$1.62 × smaller than that of vanilla speculative decoding, and 2.84×2.84$\times$2.84 × smaller than that of greedy decoding.
Figure 3: Average number of accepted tokens per iteration of speculative decoding and MJSD. Each iteration generates 4 draft tokens.
5.2 Ablation Study
Table 3: Effect of number of beams to the inference performance on ChatGPT-Prompts dataset.
# beams 2 4 6 8
speed (token/s) $\uparrow$ 55.9 59.9 60.2 61.3
Llama-2 energy (J/token) $\downarrow$ 2.43 2.25 2.22 2.20
perplexity $\downarrow$ 2.44 2.12 2.14 2.10
speed (token/s) $\uparrow$ 51.0 54.1 54.3 55.9
OPT energy (J/token) $\downarrow$ 2.50 2.32 2.36 2.30
perplexity $\downarrow$ 3.63 3.16 3.42 3.19
5.2.1 Effects of Number of Beams
First, we investigate how the number of beams used in the beam decoding of the small model affects the inference performance. Table 3 shows the results. Increasing the number of beams improves the quality of the draft tokens, which not only improves the output perplexity but also increases the average acceptance length and hence leads to better efficiency. But we can see that the increment slows down when the number of beams is large enough. In addition, when the number of beams is too large, the inference cost of the small model will become too high.
5.2.2 Effects of Acceptance Threshold
Figure 4: Effect of acceptance threshold on output perplexity and decoding speed on ChatGPT-Prompts dataset.
Next, we evaluate the effect of acceptance threshold $\tau$. Intuitively, when we increase $\tau$ from 0 to 1, the acceptance criterion becomes more strict, the efficiency drops while the output perplexity increases. Surprisingly, this expectation is only partially correct. As shown in Figure 4, the efficiency indeed drops when $\tau$ increases. However, the perplexity increases when $\tau$ is close to 1. When $\tau=1$, all the draft tokens are rejected, which makes MJSD equivalent to greedy decoding. Similarly, when $\tau$ is close to 1, the advantage of multi-token joint greedy decoding on effectiveness disappears, hence the perplexity becomes close to the perplexity of greedy decoding. Another surprising observation is that the perplexity of MJSD is good when $\tau=0$. When $\tau=0$, MJSD is equivalent to generating $\gamma$ tokens using beam decoding with the small model, then generating an additional token with the large model. The fact that MJSD achieves good perplexity when $\tau=0$ can be explained by the fact that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s predictions deviate from those of the large model” [17, 20]. Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MJSD is relatively stable, suggesting that MJSD is not sensitive to the acceptance threshold.
6 Related Work
Efficient Decoding Inference. There are extensive studies on improving large model inference efficiency. Well-known methods include model quantization [9, 22], model pruning [10, 28], and model distillation [15]. Despite achieving significant speed-ups, a common drawback of these methods is that they have to sacrifice the model’s effectiveness.
A direction closer to our work is non-autoregressive decoding. It is first proposed by [12] to generate multiple tokens in parallel. That is, the model simultaneously predicts $p(x_{t+k}|x_{1:t})$ ($k=1,2,\ldots$). Subsequent studies further improved the performance of parallel decoding by incorporating additional information [35, 31, 21] or using additional iterations to refine predictions [11, 19, 13]. However, these works require continuous training of the model and usually either compromise the model effectiveness or require task-dependent techniques to achieve comparable performance [17].
Speculative Decoding. Speculative decoding was recently proposed in [20, 5] as a way to accelerate LLM inference. Spectr [32] enhances speculative decoding by letting the small model generate multiple i.i.d. draft sequences. While speculative decoding and Spectr use the large model to verify all the tokens drafted by the small model, BiLD [17] only calls the large model when the probability output by the small model is below a pre-defined threshold $\tau_{1}$. The large model rejects a token if its negative log-likelihood is larger than threshold $\tau_{2}$. SpecInfer [24] uses one or multiple small models to generate a draft token tree to increase the average acceptance length for each iteration. All these methods can be perceived as exact or approximate versions of sampling tokens from the conditional distribution $p(x_{t}|x_{<t})$. Therefore, their output perplexity is bounded by greedy decoding.
An orthogonal direction to improve speculative decoding is to improve the effectiveness of the small draft model. It is obvious that if more draft tokens are accepted, the overall inference speed will increase. BiLD [17] uses a model prediction alignment technique to better train the small model. Liu et al. [23] propose online speculative decoding to continually update the draft model based on observed input data. Instead, Rest [14] uses a retrieval model to produce draft tokens. The improvement of small models can benefit all speculative decoding algorithms, including our method.
7 Conclusion
We introduce multi-token joint speculative decoding that significantly enhances output quality with better time and energy efficiency. A distinctive aspect of our work is the exploration of speculative decoding’s impact on inference energy consumption, an often neglected area in existing studies. Experiment results demonstrate our method achieves significant energy reduction. This research contributes not only a novel decoding approach but also valuable insights for optimizing LLM deployment in real-world applications where considerations of both quality and efficiency are crucial.
8 Limitations and Impact Statements
Limitations. This paper mainly considers improving the output perplexity of decoding algorithms. Under the assumption that the model is well-trained for downstream tasks, improving output perplexity usually leads to better downstream effectiveness. But if the model is not well-trained, improving the perplexity may not necessarily improve the downstream effectiveness. However, it is mainly the problem of the model itself. In real-world applications, it is reasonable to assume the model is properly trained for downstream tasks.
Impact Statements. The goal of this work is to advance the field of Large Language Model (LLM) Inference, which has received tremendous attention from both academia and industry. However, LLMs also have received many critiques, including their extremely large carbon footage emission during both training and inference. Our work pays special attention to their energy consumption during inference to provide high-quality and fast-inference LLMs with reduced energy consumption, which has become a serious concern with the rapid increase of LLM-related workloads in the past few years.
Appendix A Pseudocode of MJSD
See Algorithm 1.
Algorithm 1 One Iteration of MJSD Algorithm
1:Input: draft model $M_{q}$, target model $M_{p}$, $input$, threshold $\tau$
[…]
4:$\bm{P}\leftarrow M_{p}(input,\bm{X})$eqparboxCOMMENT# $\bm{P}\in\mathbf{R}^{(\gamma+1)\times|V|}$, $\bm{P}_{i,j}=p(x=j|\bm{x}_{1:i-1},input)$
[…]
6:$p\leftarrow 1$, $\eta\leftarrow-1$
7:for $i=1$ to $\gamma$ do
8:     $j\leftarrow\bm{x}_{i}$
9:     $p\leftarrow p*\bm{P}_{i,j}$, $q\leftarrow\bm{q}_{i}$
10:     if $\tau<\min(1,\frac{p}{q})$ then
11:         $\eta\leftarrow j$eqparboxCOMMENT# longest accepted prefix so far
12:     end if
13:end for
[…]
15:$\bm{p}^{\prime}\leftarrow P_{\eta+1}$
16:$t\sim\bm{p}^{\prime}$
17:return $[\bm{x}_{1},\ldots,\bm{x}_{\eta},t]$
Table 4: Dataset Statistics
Dataset Task # Test Input Avg. Input Len
ChatGPT-Prompt Instruction 360 25.2
ChatAlpaca Chat 43.081 277.7
CNNDM Summarization 11,490 3,967.1
Appendix B Energy Consumption Measurement
We use the command "nvidia-smi --query-gpu=power.draw --format=csv" to get GPU power every second, and sum them up as the total energy consumption. We use average energy consumption per token to measure energy efficiency. There is a recent study pointing out the measurement error using nvidia-smi [37]. We follow the three principles proposed in [37] to minimize the error.
Appendix C Configuration
The experiments are conducted on a machine with 1 Nvidia L40 GPU (48 GB), 4 CPUs, and 50 GB main memory, using a batch size of 1, which is common for online serving [29]. We set the maximum running time to be an hour for each baseline. We use average tokens/second to measure the inference speed and use perplexity (exponentiated average negative log-likelihood) based on the probability of the large model to measure the output quality. Because different methods might finish different numbers of inputs, we only calculate the perplexity of the first $M$ inputs, where $M$ is the number of inputs finished by greedy decoding. We use average energy consumption per token to measure energy efficiency. The details of energy measurement are illustrated in the Appendix.
Appendix D Hyper-parameter Details
In the experiments, we follow the default settings in [2] to warp the sampling distribution $p$ and $q$ with the following steps, which are the default warpping operations in a public speculative decoding implementation.
  1. 1.
    Keep the probabilities of top 20 tokens unchanged, and set the probabilities of other tokens to 0, then normalize the distribution.
  2. 2.
    Sort the tokens based on their distributions descendingly. Keep the first $K$ tokens such that their cumulative probabilities is larger than 0.9, while set the probabilities of other tokens to 0.
For different methods, we choose their hyper-parameters by using a small validation set. We select the set of hyper-parameters that make the corresponding method have best output perplexity. Table 5 shows the hyper-parameters used in the experiments.
Table 5: Hyper-parameters of different methods for different models and datasets. CP: ChatGPT-Prompts, CA: ChatAlpaca, CD: CNNDaily.
Llama,CP OPT,CP Llama,CA OPT,CA Llama,D OPT,CD
speculative step len $\gamma$ 4 4 4 4 4 4
Spectr step len $\gamma$ 4 4 4 4 4 4
num seq $m$ 4 4 2 2 4 2
BiLD step len $\gamma$ 10 10 10 10 10 10
fallback thres $\tau_{1}$ 0.9 0.9 0.9 0.3 0.9 0.3
rollback thres $\tau_{2}$ 2 2 1 2 3 2
MJSD step len $\gamma$ 4 4 4 4 4 4
num beams 8 8 8 8 8 8
acc/rej thres $\tau$ 0.1 0.1 0.1 0.1 0.1 0.1
Appendix E License of Datasets and Models
Datasets:
  • ChatGPT-Prompts: Non (https://huggingface.co/datasets/MohamedRashad/ChatGPT-prompts)
  • ChatAlpaca: Apache-2.0 License
  • CNN Dailymail: Apache-2.0 License
Models
Codes
Appendix F More Discussion on Multi-Token Joint Greedy Decoding
It is not trivial to prove that MJGD is better than vanilla greedy decoding. We might tend to show that given the same $x_{1:t}$ and $prefix$, the $\gamma_{i}$ tokens generated by MJGD has higher likelihood than the $\gamma_{i}$ tokens generated by calling greedy decoding $\gamma_{i}$ times. However, at step $i$ ($\forall i\geq 2$), the previously generated tokens $x_{1:t}$ are different for the two decoding algorithms, making it difficult to compare the overall perplexity.
To explain why MJGD is more effective, let us consider an optimal algorithm that generates next $\gamma_{i}$ tokens at one step based on a score function:
$s^{*}(x_{t+1:t+\gamma_{i}})=\max_{x_{t+\gamma_{i}+1:N}}\log p(x_{t+1:N}|x_{1:t})$ (3)
which is the log-likelihood of the optimal future sequence with prefix $x_{1:t}$ and the generated $\gamma_{i}$ tokens. The optimal greedy decoding chooses $x_{t+1:t+\gamma_{i}}$ that maximizes the score function and is guaranteed to return the sequence with the optimal perplexity.
Both MJGD and vanilla greedy decoding can be viewed as approximations to the optimal greedy decoding by replacing $s^{*}(\cdot)$ with $s_{MJGD{}}$ or $s_{GD}$ that estimates the values of $s^{*}(\cdot)$. Specifically, since MJGD selects the $\gamma_{i}$ tokens with the largest joint likelihood, we have $s_{MJGD}(\cdot)=\log p(x_{1:t+\gamma_{i}}|x_{1:t})+c_{1}$ where $c_{1}$ is an arbitrary constant. Notice that adding a constant to $s(\cdot)$ changes the error of the estimation but does not change the behavior of the algorithm.
On the other hand, greedy decoding generates the $\gamma_{i}$ tokens one by one. At the last token, greedy decoding selects the token with the largest $p(x_{t+\gamma_{i}}|x_{1:t+\gamma_{i}-1})$, which is equivalent to selecting the tokens with the largest $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ provided that $x_{t+1:t+\gamma_{i}-1}$ are the same tokens generated by previous steps. So for greedy decoding, $s_{GD}(x_{t+1:t+\gamma_{i}})=\log p(x_{1:t+\gamma_{i}}|x_{1:t})+c_{2}$ ($c_{2}$ is also an arbitrary constant) if and only if $x_{t+i}=\arg\max_{x}p(x|x_{1:t+i-1})$ for $i=1,\ldots,\gamma_{i}-1$. Otherwise, the score of $x_{t+1:t+\gamma_{i}}$ has to be smaller than any possible value of $\log p(x_{1:t+\gamma_{i}}|x_{1:t})+c_{2}$. Notice that the lower bound of $\log p(x_{1:t+\gamma_{i}}|x_{1:t})$ is $\infty$- ∞, so $s_{GD}(x_{t+1:t+\gamma_{i}})=-\infty$ if and only if $\exists j=1,\ldots,\gamma_{i}-1$ such that $x_{t+j}\neq\arg\max_{x}p(x|x_{1:t+i-1})$.
Let $c^{*}=\arg\min_{c_{2}}\sum_{x_{t+1:t+\gamma_{i}}}|s_{GD}(x_{t+1:t+\gamma_{i}})-s^{*}(x_{t+1:t+\gamma_{i}})|$. By setting $c_{1}=c^{*}$, for $x_{t+1},\ldots,x_{t+\gamma_{i}-1}$ satisfying that $x_{t+i}=\arg\max_{x}p(x|x_{1:t+i-1})$, $\forall i=1,\ldots,\gamma_{i}-1$, we have
$s_{MJGD}(x_{t+1:t+\gamma_{i}})=s_{GD}(x_{t+1:t+\gamma_{i}})=\log p(x_{1:t+\gamma_{i}|x_{1:t}})+c^{*}$
Otherwise, we have
$|s_{MJGD}(x_{t+1:t+\gamma_{i}})-s^{*}(x_{t+1:t+\gamma_{i}})|=|c^{*}-h^{*}(x_{t+1:t+\gamma_{i}})|\leq|-\inf-s^{*}(x_{t+1:t+\gamma_{i}})|$
where $h^{*}(x_{t+1:t+\gamma_{i}})=\max_{x_{t+\gamma_{i}+1:N}}\log p(x_{t+\gamma_{i}+1:N}|x_{1:t+\gamma_{i}},prefix)$. Therefore,
$|s_{MJGD}(x_{t+1:t+\gamma_{i}})-s^{*}(x_{t+1:t+\gamma_{i}})|\leq|s_{GD}(x_{1:K})-s^{*}(x_{t+1:t+\gamma_{i}})|$ (4)
The derivation above provides a heuristic explanation to show that MJGD is a closer approximation to the optimal greedy decoding than vanilla greedy decoding. Hence, its output perplexity is generally better.

### v2

@@
-Multi-Token Joint Speculative Decoding for Accelerating Large Language Model Inference
+Optimized Multi-Token Joint Decoding with Auxiliary Model for LLM Inference
 Abstract
-Transformer-based Large language models (LLMs) have demonstrated their power in various tasks, but their inference incurs significant time and energy costs.
-To accelerate LLM inference, speculative decoding uses a smaller model to propose one sequence of tokens, which are subsequently validated in batch by the target large model.
-Compared with autoregressive decoding, speculative decoding generates the same number of tokens with fewer runs of the large model, hence accelerating the overall inference by $1$-2×2$\times$2 ×.
-However, greedy decoding is not the optimal decoding algorithm in terms of output perplexity, which is a direct measurement of the effectiveness of a decoding algorithm.
-An algorithm that has better output perplexity and even better efficiency than speculative decoding can be more useful in practice.
-To achieve this seemingly contradictory goal, we first introduce multi-token joint greedy decoding (MJGD), which greedily generates multiple tokens at each step based on their joint perplexity.
-We show that it leads to better perplexity for the whole output.
-But the computation cost of MJGD is infeasible in practice.
-So we further propose multi-token joint speculative decoding (MJSD), which approximates and accelerates the MJGD from two aspects: it approximates the joint distribution of the large model with that of a small model, and uses a verification step to guarantee the accuracy of approximation; then it uses beam decoding to accelerate the sequence generation from the joint distribution.
-Compared with vanilla speculative decoding, MJSD has two advantages: (1) it is an approximation of MJGD, thus achieving better output perplexity; (2) verification with joint likelihood allows it to accept the longest prefix sub-sequence of the draft tokens with valid perplexity, leading to better efficiency.
-In addition, we analyze energy consumption during LLM inference and provide theoretical and empirical evidence that, surprisingly, MJSD reduces energy consumption even when the number of operations increases.
-Experiment results show that our approach enhances perplexity by $21.2\%$ than greedy decoding.
-Moreover, MJSD achieves 2.21×2.21$\times$2.21 × speed-up and 2.84×2.84$\times$2.84 × less energy consumption than greedy decoding, and 1.49×1.49$\times$1.49 × speed-up and 1.62×1.62$\times$1.62 × less energy consumption than vanilla speculative decoding.
+Large language models (LLMs) have achieved remarkable success across diverse tasks, yet their inference processes are hindered by substantial time and energy demands due to single-token generation at each decoding step.
+While previous methods such as speculative decoding mitigate these inefficiencies by producing multiple tokens per step, each token is still generated by its single-token distribution, thereby enhancing speed without improving output quality.
+In contrast, our work simultaneously enhances inference speed and improves the output effectiveness.
+We consider multi-token joint decoding (MTJD), which generates multiple tokens from their joint distribution at each iteration, theoretically reducing perplexity and enhancing task performance.
+However, MTJD suffers from the high cost of sampling from the joint distribution of multiple tokens.
+Inspired by speculative decoding, we introduce multi-token assisted decoding (MTAD), a novel framework designed to accelerate MTJD.
+MTAD leverages a smaller auxiliary model to approximate the joint distribution of a larger model, incorporating a verification mechanism that not only ensures the accuracy of this approximation, but also improves the decoding efficiency over conventional speculative decoding.
+Theoretically, we demonstrate that MTAD closely approximates exact MTJD with bounded error.
+Empirical evaluations using Llama-2 and OPT models ranging from 13B to 70B parameters across various tasks reveal that MTAD reduces perplexity by 21.2% and improves downstream performance compared to standard single-token sampling.
+Furthermore, MTAD achieves a 1.42$\times$ speed-up and consumes 1.54$\times$ less energy than conventional speculative decoding methods.
+These results highlight MTAD’s ability to make multi-token joint decoding both effective and efficient, promoting more sustainable and high-performance deployment of LLMs.
 1 Introduction
-Large Language Models (LLMs) have exhibited remarkable performance across real-world tasks spanning text and image domains [4, 7, 33, 34].
-However, the substantial parameter size of these models leads to high computational cost during inference [27].
-For example, ChatGPT is estimated to have an energy demand of 564 MWh per day [8].
-Compounding this issue, the autoregressive generation of $K$ tokens necessitates calling the model $K$ times, involving the repeated loading of weight matrices and intermediate results from GPU global memory to computing units.
-As a result, LLM inference is often hampered by constraints related to memory bandwidth and communication [20], leading to diminished hardware utilization, heightened inference latency, and elevated energy cost.
+Large Language Models (LLMs) such as GPT-4 and Llama-2 (Touvron et al., 2023) have demonstrated extraordinary capabilities across a wide range of tasks (Brown et al., 2020; Chowdhery et al., 2023; Thoppilan et al., 2022; Touvron et al., 2023).
+Despite their impressive performance, the deployment of LLMs is often constrained by substantial inference costs in terms of time and energy.
+This inefficiency primarily stems from the autoregressive nature of these models, where generating a sequence of $K$ tokens requires $K$ separate model calls.
+Each call involves loading large weight matrices and intermediate results from GPU global memory to computing units, leading to repeated memory accesses and limited hardware utilization (Samsi et al., 2023; Leviathan et al., 2023).
 To tackle this challenge, researchers have delved into non-autoregressive decoding approaches.
-Earlier methods [11, 12, 13] aimed at reducing inference latency by concurrently generating multiple tokens.
-But these methods usually require task-dependent techniques and information to achieve comparable performance to that of autoregressive decoding [17, 36].
-More recently, speculative decoding has emerged [20, 5, 17, 32], exploiting the observation that, even when a small model is ten times smaller than a large model, most of the small model’s prediction aligns well with that of a large model.
-So it leverages a smaller auxiliary model to draft a few future tokens autoregressively, which are subsequently validated in parallel by the larger model.
+Early methods (Ghazvininejad et al., 2019; Gu et al., 2017; Guo et al., 2020) aimed at reducing inference latency by concurrently generating multiple tokens.
+But these methods usually require task-dependent techniques and information to match the performance of autoregressive decoding (Kim et al., 2023; Xiao et al., 2023).
+More recently, speculative decoding has emerged (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023), exploiting the observation that most of the small model’s prediction aligns well with that of a large model.
+It leverages a smaller auxiliary model to draft a few future tokens autoregressively, which are subsequently validated in parallel by the larger model.
 As the smaller model operates significantly faster and parallel token verification incurs a similar time cost as generating a single token, speculative decoding achieves an overall speed-up of $1$-2×2$\times$2 ×.
-Meanwhile, it always has an identical sampling distribution as greedy decoding.
-However, there are some limitations of the existing studies on speculative decoding.
-First, they aim to have an identical sampling distribution as greedy sampling.
-But the goal of decoding algorithms is to generate sequences with optimal perplexity, which is defined as the exponentiated average negative log-likelihood of all tokens.
-An algorithm that does not follow the same sampling distribution as greedy decoding, but yields output with better perplexity could be more useful in the real world.
-Second, in these algorithms, if a draft token is rejected, all the draft tokens after it are discarded.
-But the overall draft sequences may have good quality despite there being a few low-quality tokens.
-Rejecting these sequences will shorten the average acceptance length for each iteration, lowering the decoding efficiency.
-Third, existing research does not investigate the impact of speculative decoding on inference energy consumption.
-While speculative decoding increases the number of FLOPs during inference, it simultaneously reduces the overall inference time, communication, and memory operations, which are crucial factors in determining energy consumption [1, 6].
-Thus, it remains unclear if speculative decoding increases or decreases the energy cost of inference.
-So we aim to design a new algorithm that is both more efficient and more effective than greedy decoding and vanilla speculative decoding.
-We first introduce multi-token joint greedy decoding (MJGD), which generates multiple tokens at each step greedily based on their joint likelihood.
-We empirically show that it has better overall perplexity than greedy decoding does.
-However, the computational cost of MJGD is infeasible in practice.
-So we further propose multi-token joint speculative decoding (MJSD), which approximates and accelerates MJGD from two aspects: (1) it approximates the joint distribution of the large model with that of a small model by generating draft tokens with the small model and validating them with the large model; (2) it uses beam decoding to accelerate the sequence generation from the joint distribution of the small model.
-Compared with vanilla speculative decoding, MJSD has two advantages: (1) it is more effective, as it approximates MJGD, which has better output perplexity than greedy decoding does; (2) it is more efficient because it uses joint likelihood as the verification criterion, which allows it to accept the longest prefix sub-sequence of the draft tokens with valid perplexity.
-In addition, we analyze the energy consumption of LLM inference.
-We give theoretical and empirical evidence that, despite that MJSD and other speculative decoding algorithms increase the number of FLOPs needed during LLM inference, they reduce the overall energy consumption by reducing the overhead induced by accessing GPU global memory.
-We evaluate our method on text generation tasks with Llama-2 [34] and OPT models [38].
-Experiment results show that our approach enhances perplexity by $21.2\%$ than the baselines with identical sampling distributions as greedy decoding.
-In addition, MJSD achieves 2.21×2.21$\times$2.21 × speed-up and 2.84×2.84$\times$2.84 × smaller energy consumption than greedy decoding, and 1.49×1.49$\times$1.49 × speed-up and 1.62×1.62$\times$1.62 × smaller energy consumption than vanilla speculative decoding.
-Our code is open-sourced [^1] https://anonymous.4open.science/r/LLMSpeculativeSampling-EE52.
+Despite gains in speed, these methods still generate each token based on its single-token probability.
+Consequently, it does not enhance the effectiveness of the generated sequences.
+In this work, we aim to go beyond the conventional trade-off between efficiency and effectiveness by introducing multi-token joint decoding (MTJD).
+Unlike traditional approaches, MTJD produces multiple tokens from their joint distribution at each decoding step.
+Theoretically, we show this joint generation can lead to lower perplexity and hence improved task performance.
+However, directly sampling from the joint distribution of multiple tokens poses significant computational challenges, rendering MTJD impractical.
+Inspired by speculative decoding, we propose multi-token assisted decoding (MTAD), a novel framework designed to approximate and accelerate MTJD.
+MTAD employs a smaller auxiliary model to estimate the joint distribution of a larger model, significantly reducing computational demands.
+To ensure the accuracy of this approximation, MTAD incorporates a verification mechanism that not only guarantees the accuracy of the draft tokens but also enhances efficiency beyond conventional speculative decoding by maximizing the number of accepted tokens per iteration.
+We provide both theoretical and empirical analyses to demonstrate that MTAD improves perplexity and downstream performance.
+Meanwhile, it achieves significant reductions in energy and time usage compared to existing decoding strategies.
+Our contributions are as follows:
+  1. 1.
+    We introduce multi-token joint decoding (MTJD), a multi-token joint decoding approach that theoretically reduces perplexity by generating tokens from their joint distribution.
+  2. 2.
+    We develop multi-token assisted decoding (MTAD), an efficient approximation of MTJD with bounded error that leverages a smaller model for distribution approximation.
+  3. 3.
+    We analyze the energy consumption of LLM inference.
+To our knowledge, we are the first to give theoretical and empirical evidence that, despite that MTAD and other speculative decoding algorithms increase the number of FLOPs needed during LLM inference, they reduce the overall energy consumption by reducing the overhead induced by accessing GPU global memory.
+  4. 4.
+    We conducted comprehensive evaluations with Llama-2 and OPT models (ranging from 13B to 70B parameters) across various tasks, demonstrating that MTAD enhances perplexity by 21.2% and improves downstream effectiveness compared to standard single-token sampling, while also achieving a 1.42$\times$ speed-up and reducing energy consumption by 1.54$\times$ compared to conventional speculative decoding methods.
+These advancements position MTAD as a robust solution for making multi-token joint decoding both effective and efficient, thereby facilitating more sustainable and high-performance deployment of large-scale language models.
+Our code is publicly available [^1] https://anonymous.4open.science/r/LLMSpeculativeSampling-EE52.
 2 Preliminaries
 2.1 Decodings of LLMs
+Decoding and Perplexity.
 Let $p$ denote the distribution defined by LLM model $M_{p}$.
-Given an input prefix $prefix$, the optimal decoding algorithm is to generate a sequence of $N$ tokens with maximum likelihood $p(x_{1:N}|prefix)$.
+Given an input context $input$, a decoding algorithm generates a sequence of $N$ tokens whose likelihood is denoted as $p(x_{1:N}|input)$.
 The likelihood of the sequence is directly linked to perplexity of the sequence, which is the exponentiated average negative log-likelihood of all tokens.
-Based on autoregressive decomposition $p(x_{1:N}|prefix)=\prod_{t=1}^{N}p(x_{t}|x_{1:t-1},prefix)$, the perplexity is defined as:
+Based on autoregressive decomposition $p(x_{1:N}|input)=\prod_{t=1}^{N}p(x_{t}|x_{1:t-1},input)$ [^2] In the paper, we omit $input$ when there is no ambiguity., the perplexity is defined as:
 $PPL(x_{1:N})=\exp\left\{-\frac{1}{N}\sum_{t=1}^{N}\log p(x_{t}|x_{1:t-1})\right\}$ (1)
+Perplexity serves as a direct metric for assessing the effectiveness of a decoding algorithm.
+In practice, when a model is well-trained, lower perplexity often correlates with improved downstream performance.
+For example, beam sampling aims to return output with lower perplexity and is proven to have better downstream performance in general (Shi et al., 2024).
+To further demonstrate the relationship between perplexity and downstream performance, we evaluate GPT-3.5-turbo on the spider (Yu et al., 2018) dataset.
+Using a temperature of 2, the model generated 10 outputs for each input.
+We measured the average perplexities and execution accuracies for the outputs with the highest, lowest, and median (the 5-th lowest) perplexity.
+As shown in Table 1, lower perplexity correlates with improved downstream performance, even in one of today’s largest models.
+Table 1: Relationship between perplexity and execution accuracy (EA, higher the better) for GPT-3.5-turbo.
+Output Avg. PPL $\downarrow$ EA (%) $\uparrow$
+Highest PPL 4.13 33
+5-th Lowest PPL 1.40 58
+Lowest PPL 1.07 62
 Now we introduce commonly used decoding approaches.
-Greedy Decoding.
-Greedy decoding samples the next single token $x_{t}$ based on $p(\cdot|x_{1:t-1},prefix)$.
-According to the autoregressive decomposition, greedy sampling leads to the desired joint distribution for the sequence.
-In practice, however, the objective of decoding algorithms is to generate sequences with high likelihood.
-Instead of sampling from the joint distribution, a warping operation $\mathcal{T}$ usually is applied to boost the high probability region, which transforms the original distribution $p$ to $\tilde{p}=\mathcal{T}\circ p$.
-A very popular warping operation is $argmax$, which only selects the sample with the highest probability.
-Other possible warping operations include top-k.
-For most of the warping operations, autoregressive decomposition no longer holds after the warping.
-For example, $\arg\max_{x_{1:N}}{p}(x_{1:N}|prefix)\neq\prod_{t=1}^{N}\arg\max_{x_{t}}{p}(x_{t}|x_{1:t-1},prefix)$.
-Beam Decoding.
-Beam decoding aims to do a better job than greedy decoding.
-For each position $t$ ($1\leq t\leq N$), it maintains $M>1$ candidate sequences, which are also called beams.
-Still take $argmax$ warping operator as an example and assume we have already kept the $M$ sequences $\mathcal{I}_{t-1}=\{x_{1:t-1}^{(1)},\ldots,x_{1:t-1}^{(M)}\}$ at position $t-1$ with highest likelihood, $M$ tokens are then sampled from $p(\cdot|x_{1:t-1}^{(m)},prefix)$ for each sequence $m$ and the top $M$ sequences with the highest likelihood $p(x_{1:t}|prefix)$ will be kept.
-It is widely recognized that beam decoding is a closer approximation to $\tilde{p}(x_{1:N}|prefix)$ than greedy decoding [18], with few exceptions under special construction.
-The computation for each beam can be parallelized, thus its run time is about the same as greedy decoding.
-The fact that beam decoding is more effective than greedy decoding also suggests that the output perplexity of greedy decoding still has room for improvement.
+Multinomial Sampling.
+Multinomial sampling, also known as standarized sampling or single-token sampling, samples the next token $x_{t}$ based on $\mathcal{T}\circ p(\cdot|x_{1:t-1},input)$, where $\mathcal{T}$ is a warping operation applied to enhance the high probability region.
+Some common warping operations include top-k warping, which limits the selection to the top k tokens, and top-p warping, where tokens are sampled from the smallest possible subset of the vocabulary whose cumulative probability mass exceeds a specified threshold.
+The deterministic version of multinomial sampling (i.e., greedy decoding) is a special case when $k=1$.
+Beam Sampling.
+Beam sampling aims to improve output perplexity over multinomial sampling.
+For each position $t$ ($1\leq t\leq N$), it maintains $W>1$ candidate sequences, which are also called beams.
+Assume we have already kept the $W$ sequences $\mathcal{I}_{t-1}=\{x_{1:t-1}^{(1)},\ldots,x_{1:t-1}^{(W)}\}$ at position $t-1$, $W$ sequences with length $t$ are then sampled from $\mathcal{T}\circ p_{beam}$, where $p_{beam}:\mathcal{I}_{t-1}\times V\rightarrow[0,1]$ is the beam sampling probability:
+$p_{beam}(x_{1:t-1}^{(i)},x_{t})=\frac{p(x_{1:t-1}^{(i)},x_{t}|input)}{\sum_{x_{1:t-1}^{(j)},x^{\prime}_{t}\in\mathcal{I}_{t-1}\times V}p(x_{1:t-1}^{(j)},x_{t}^{\prime}|input)}$ (2)
+Notice that $p(x_{1:t-1}^{(i)},x_{t}|input)=p(x_{t}|x_{1:t-1}^{(i)},input)\cdot p(x_{1:t-1}^{(i)}|input)$.
+In practice, beam sampling stores the likelihood $p(x_{1:t-1}^{(i)}|input)$ for each beam, and the computation complexity of $p_{beam}$ is $O(W\cdot|V|)$.
+In deterministic beam sampling, the top $W$ sequences with the highest likelihood $p_{beam}(x_{1:t})$ will be kept.
 2.2 Vanilla Speculative Decoding
-Besides effectiveness, speculative decoding is proposed by [20, 5] to accelerate the inference of LLMs.
+Besides effectiveness, speculative decoding is proposed by (Leviathan et al., 2023; Chen et al., 2023) to accelerate the inference of LLMs.
 It utilizes a small model to generate the next $\gamma$ tokens and then uses the large model to verify the drafted tokens in parallel, which is summarized below:
   1. 1.
-    Let $prefix$ be the input context, the small model samples $\gamma$ draft tokens $x_{1},\ldots,x_{\gamma}$ using greedy decoding based on the warped predicted conditional probability $\tilde{q}(x_{t}|x_{1:t-1},prefix))$ for $t=1,\ldots,\gamma$, where $\tilde{q}=\mathcal{T}\circ q$ and $q$ is the small model’s output distribution.
+    Let $input$ be the input context, the small model samples $\gamma$ draft tokens $x_{1},\ldots,x_{\gamma}$ using multinomial sampling based on $\tilde{q}(x_{t}|x_{1:t-1},input))$ for $t=1,\ldots,\gamma$, where $\tilde{q}=\mathcal{T}\circ q$ and $q$ is the small model’s output distribution.
   2. 2.
-    The large model verifies the draft tokens in parallel by computing the conditional probability $\tilde{p}(x_{t}|x_{1:t-1},prefix)$ for $t=1,\ldots,\gamma$.
+    The large model verifies the draft tokens in parallel by computing the conditional probability $\tilde{p}(x_{t}|x_{1:t-1},input)$ for $t=1,\ldots,\gamma$.
   3. 3.
@@
 An additional token is sampled from a residual distribution as a correction to the first rejected token.
-Then the accepted tokens and the resampled token are appended to the context $prefix$ as the input to the next iteration.
+Then the accepted tokens and the resampled token are appended to the context $input$ as the input to the next iteration.
   4. 4.
@@
 Meanwhile, although the small model still runs in an autoregressive way, its inference speed is much faster than the large model.
-So speculative decoding can accelerate the inference process of LLMs.
-Additionally, it is proven that each token $x_{i}$ generated by speculative sampling follows the identical sampling distribution $\tilde{p}(x_{t}|x_{1:t-1},prefix)$ as greedy decoding [20].
-However, since $\tilde{p}(x_{1:N}|prefix))\neq\prod_{t=1}^{N}\tilde{p}(x_{t}|x_{1:t-1},prefix)$, the effectiveness of speculative decoding also has room for improvement.
-Therefore, instead of having an identical sampling distributon as greedy decoding, our algorithm aims to achieve a better approximation to $\tilde{p}(x_{1:N}|prefix)$.
+As a result, speculative decoding achieves a speedup of 1–2× compared to multinomial sampling while maintaining an identical sampling distribution.
 3 Methodology
-The perplexity is a direct measurement of the effectiveness of a decoding algorithm.
-In practice, when the model is well-trained, a better perplexity usually leads to better downstream effectiveness.
-Although it is possible that improving perplexity does not improve the performance on downstream tasks, the problem is due to the model itself instead of the decoding algorithm.
-Therefore, the goal of this work is to design an algorithm that yields better output perplexity and better efficiency than greedy decoding and vanilla speculative decoding.
-In this section, we first introduce multi-token joint greedy decoding (MJGD) to give a motivating example for why generating multiple tokens based on their joint likelihood can lead to better perplexity.
-Then we introduce multi-token joint speculative decoding (MJSD), which approximates and accelerates MJGD by exploiting a small model.
-3.1 Multi-Token Joint Greedy Decoding
-Let $M_{p}$ be the large target model, the goal is to generate a sequence of $N$ tokens with maximum likelihood $p(x_{1:N}|prefix)$.
-We now introduce a new decoding algorithm to improve greedy decoding in terms of perplexity.
+As discussed in Section 2, the goal of this work is to design an algorithm that yields lower perplexity and better efficiency than multinomial sampling and vanilla speculative decoding.
+In this section, we first introduce multi-token joint decoding (MTJD), which generates multiple tokens based on their joint likelihood.
+We prove it can yield lower perplexity.
+Then we introduce multi-token assisted decoding (MTAD), which approximates and accelerates MTJD by exploiting an auxiliary model.
+3.1 Multi-Token Joint Decoding
+We first introduce a new decoding algorithm to improve multinomial sampling in terms of perplexity.
 Definition 3.1.
-Multi-Token Joint Greedy Decoding.
+Multi-Token Joint Decoding.
 Let $M_{p}$ be the large target model with distribution $p$.
-Different from single-token greedy decoding, multi-token joint greedy decoding (MJGD) generates the next $\gamma_{i}$ tokens at step $i$ based on their joint conditional probability $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^2] In the paper, we omit $prefix$ when there is no ambiguity., where $\gamma_{i}$ is an integer no less than 1 and $t=\sum_{i^{\prime}=1}^{i-1}{\gamma_{i^{\prime}}}$, i.e., the total tokens generated in the previous $i-1$ steps.
-Figure 1: Perplexity and rougeL score of the output when $\gamma_{i}=K$ for MJGD with OPT-125M and Llama-2-68M fined-tuned on ChatGPT-Prompts [26] dataset.
-Greedy decoding is a special case of MJGD where $\gamma_{i}=1$, $\forall i$.
-When $\gamma_{1}=N$, MJGD generates the optimal sequence by returning the sequence with the highest likelihood.
+Different from single-token multinomial sampling, multi-token joint decoding (MTJD) generates the next $\gamma_{i}$ tokens at step $i$ based on their joint conditional probability $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$, where $\gamma_{i}$ is an integer no less than 1 and $t=\sum_{i^{\prime}=1}^{i-1}{\gamma_{i^{\prime}}}$, i.e., the total tokens generated in the previous $i-1$ steps.
+Figure 1: Perplexity and Rouge-L score of the output when $\gamma_{i}=K$ for MTJD with OPT-125M and Llama-2-68M fine-tuned on ChatGPT-Prompts (Rashad, 2023) dataset.
+Multinomial sampling is a special case of MTJD where $\gamma_{i}=1$, $\forall i$.
+When $\gamma_{1}=N$, MTJD generates the sequence directly based on their joint likelihood.
 So intuitively, output perplexity should improve as $\gamma_{i}$ increases.
-Additionally, generating $\gamma_{i}$ tokens simultaneously allows MJGD to return the optimal $\gamma_{i}$ tokens for step $i$.
-In contrast, vanilla greedy decoding selects each token without considering any future tokens.
-So MJGD is less prone to choosing local optima than greedy decoding.
+Besides, generating $\gamma_{i}$ tokens simultaneously allows MTJD to consider their interactions.
+In contrast, multinomial sampling selects each token without considering any future tokens.
+So MTJD is less prone to choosing local optima.
+Theorem 3.2 shows the limit of perplexity of MTJD when $N$ approaches infinity. The proofs are included in the Appendix A.
+Theorem 3.2.
+Assume at the $i$-th ($i=1,\ldots,N$) iteration, MTJD generates $\gamma_{i}$ tokens.
+Let $\Gamma_{i}$ denote the total number of tokens generated at the first $i$ iterations.
+Let $x_{1:\Gamma_{N}}$ denote the generated tokens.
+When $N\rightarrow\infty$
+$PPL_{p}(x_{1:\Gamma_{N}})\rightarrow\exp\left(-\frac{1}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{p}(\gamma,\tilde{p})\right)$ (3)
+where $\bar{\gamma}$ is the expected number of $\gamma_{i}$, $\tilde{p}=\mathcal{T}\circ p$ represents how we sample the next $\gamma_{i}$ tokens from $p$ (e.g., in deterministic sampling, $\tilde{p}=\arg\max\circ p$ always returns the tokens with the highest joint likelihood), and $L_{p}(\gamma,\tilde{p})$ is the expected log-likelihood of the $\gamma$ tokens sampled from $\tilde{p}$:
+$L_{p}(\gamma,\tilde{p})=\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{x_{t+1:t+\gamma}}\tilde{p}(x_{t+1:t+\gamma}|x_{1:t})\log p(x_{t+1:t+\gamma}|x_{1:t})$ (4)
+Here $\mathcal{X}$ is the space of all possible inputs.
+Corollary 3.3.
+Based on Theorem 3.2, we can show that when $N\rightarrow\infty$, greedy MTJD (i.e., top-1 MTJD sampling) has lower perplexity than greedy decoding (top-1 single-token sampling).
 Empirical evidence supports our claim.
-Figure 1 shows the output perplexity and downstream ROUGE-L scores of MJGD with $\gamma_{i}$ set to a constant $K$, where $K=1,\ldots,5$.
-Notice that setting $K=1$ is equivalent to vanilla greedy decoding.
-We use beam decoding to approximate the $\arg\max$ sampling from the joint distribution $p(x_{t+1:t+K}|x+{1:t},prefix)$.
-We can see that when $K$ increases from 1 to 2, there is a significant improvement in perplexity.
-The perplexity continues to decrease when $K$ increases further.
+We fine-tune both a Llama and an OPT model on the ChatGPT-Prompts dataset and evaluate the output perplexity and Rouge-L scores with example outputs.
+Figure 1 shows the output perplexity and Rouge-L scores of MTJD with $\gamma_{i}$ set to a constant $K$, where $K=1,\ldots,5$.
+Notice that setting $K=1$ is equivalent to multinomial sampling.
+We use beam sampling to approximate the $\arg\max$ sampling from the joint distribution $p(x_{t+1:t+K}|x+{1:t},input)$.
+We can see that the perplexity keeps dropping when $K$ increases.
 It confirms our claim that increasing $\gamma_{i}$ will increase the output perplexity.
-So $\prod_{i=1}^{n}\tilde{p}(x_{\Gamma_{i-1}+1:\Gamma_{i}}|prefix,x_{1:\Gamma_{i-1}})$ is a better approximation to $\tilde{p}(x_{1:N}|prefix)$ than $\prod_{t=1}^{N}\tilde{p}(x_{t}|prefix,x_{1:t-1})$.
-Here $\Gamma_{0}=0$, and $\Gamma_{i}=\sum_{j=1}^{i}\gamma_{i}$ for $i\geq 1$, and $n$ is the total number of steps of MJGD.
 Moreover, the Rouge-L score also improves with $K$, supporting our claim that better perplexity reflects enhanced performance in downstream tasks.
-This demonstrates that MJGD not only achieves lower perplexity but also yields higher quality text generation, making it a superior decoding strategy compared to vanilla greedy decoding [^3] Additional discussions comparing MJGD and greedy decoding are provided in the Appendix F..
-3.2 Multi-Token Joint Speculative Decoding
-Unfortunately, the computation cost of MJGD is infeasible in practice, since the time and space complexity to compute the joint distribution of $\gamma_{i}$ tokens is $|V|^{\gamma_{i}}$.
-Inspired by speculative decoding and the facts that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s prediction deviate from those of the large model” [20, 17], we propose multi-token joint speculative decoding (MJSD), which exploits a small model $M_{q}$ to accelerate MJGD approximately.
+3.2 Multi-Token Assisted Decoding
+Unfortunately, the computation cost of MTJD is infeasible in practice, since the time and space complexity to compute the joint distribution of $\gamma_{i}$ tokens is $|V|^{\gamma_{i}}$.
+Inspired by speculative decoding and the facts that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s prediction deviate from those of the large model” (Leviathan et al., 2023; Kim et al., 2023), we propose multi-token assisted decoding (MTAD), which exploits a small auxiliary model $M_{q}$ to accelerate MTJD approximately.
 The core idea is to (1) use the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$ output by $M_{q}$ to approximate $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^3] It is also valid to approximate $\tilde{p}$ with $\tilde{q}$.
-Without loss of generality, we consider non-warped distribution in the illustration of MJSD.
-and generate $\gamma$ draft tokens from $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$, then (2) use the large model to validate draft tokens in parallel and accept the longest draft sub-sequence that passes verification, and (3) sample an additional token from the distribution of the large model without extra overhead to ensure at least one token is generated at each iteration.
+Without loss of generality, we consider non-warped distribution in the illustration of MTAD.
+and generate $\gamma$ draft tokens from $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$, then (2) use the large model to validate draft tokens in parallel and accept the longest draft prefix sub-sequence that passes verification, and (3) sample an additional token from the distribution of the large model without extra overhead to ensure at least one token is generated at each iteration.
 However, it is still infeasible to directly generate draft tokens from the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$.
-So we propose to further approximate this process with beam decoding, which is an effective and efficient algorithm to generate sequences with high likelihood.
-In this way, MJSD reduces the number of runs of the large model to generate $N$ tokens, thus accelerating the inference in the same way as vanilla speculative decoding does.
-Algorithm 1 in the Appendix illustrates the pseudocode of MJSD algorithm.
-Figure 2: An example of MJSD’s verification process. MJSD accepts the longest draft sub-sequence that passes verification based on joint likelihood.
+So we propose to further approximate this process with beam sampling, which is an effective and efficient algorithm to generate sequences with high likelihood.
+In this way, MTAD reduces the number of runs of the large model to generate $N$ tokens, thus accelerating the inference in the same way as vanilla speculative decoding does.
+Algorithm 1 in the Appendix illustrates the pseudocode of MTAD algorithm.
+Figure 2: An example of MTAD’s verification process. MTAD accepts the longest draft sub-sequence that passes verification based on joint likelihood.
 Draft Tokens Verification
-Figure 2 illustrates the verification process of MJSD.
-Let $x_{t+1},\ldots,x_{t+\gamma}$ be the draft tokens generated by beam decoding with the small model.
-Since beam decoding is a widely recognized algorithm to generate sequences with high overall likelihood [18], it is reasonable to assume $q(x_{t+1:t+\gamma}|x_{1:t})$ is large.
-Also, since beam decoding works in an autoregressive way, we can also assume that $\forall j\in\{1,\ldots,\gamma\}$, $q(x_{t+1:t+j}|x_{1:t})$ is large.
-To approximate MJGD, for each step $i$, MJSD needs to ensure the accepted tokens $x_{t+1:t+\gamma_{i}}$ ($0\leq\gamma_{i}\leq\gamma$) also have high joint likelihood with the large model $M_{p}$.
-So MJSD first computes the joint likelihood $p(x_{t+1:t+j}|x_{1:t})$ for $j=1,\ldots,\gamma$.
+Figure 2 illustrates the verification process of MTAD.
+Let $x_{t+1},\ldots,x_{t+\gamma}$ be the draft tokens generated by beam sampling with the auxiliary model.
+Since beam sampling is a widely recognized algorithm to generate sequences with high overall likelihood (Leblond et al., 2021), it is reasonable to assume $q(x_{t+1:t+\gamma}|x_{1:t})$ is large.
+Also, since beam sampling works in an autoregressive way, we can also assume that $\forall j\in\{1,\ldots,\gamma\}$, $q(x_{t+1:t+j}|x_{1:t})$ is large.
+To approximate MTJD, for each step $i$, MTAD needs to ensure the accepted tokens $x_{t+1:t+\gamma_{i}}$ ($0\leq\gamma_{i}\leq\gamma$) also have high joint likelihood with the large model $M_{p}$.
+So MTAD first computes the joint likelihood $p(x_{t+1:t+j}|x_{1:t})$ for $j=1,\ldots,\gamma$.
 Then for each prefix sub-sequence $x_{t+1:t+j}$, it passes verification if and only if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, where $\tau\in[0,1)$ is a pre-defined threshold.
@@
 And if $q(x_{t+1:t+j}|x_{1:t})\leq p(x_{t+1:t+j}|x_{1:t})$, it means the sub-sequence has higher likelihood in the large model, then it is reasonable to accept it.
-After verifying all the sub-sequences, MJSD accepts the longest draft sub-sequence that passes verification.
-The verification step of MJSD ensures that the accepted tokens have a high joint likelihood with the large model.
+After verifying all the sub-sequences, MTAD accepts the longest prefix sub-sequence that passes verification.
+The verification step of MTAD ensures that the accepted tokens have a high joint likelihood with the large model.
 We have shown that selecting multiple tokens based on their joint likelihood lead to better output perplexity.
-Thus, MJSD is more effective than greedy decoding and vanilla speculative decoding.
-Furthermore, since MJSD accepts the longest draft sub-sequence with high likelihood, it can tolerate low-quality tokens as long as the joint likelihood is high.
-So at each iteration, MJSD can accept more draft tokens than vanilla speculative decoding, which results in better efficiency.
+Thus, MTAD is more effective than multinomial sampling and vanilla speculative decoding.
+Furthermore, since MTAD accepts the longest draft sub-sequence with high likelihood, it can tolerate low-quality tokens as long as the joint likelihood is high.
+So at each iteration, MTAD can accept more draft tokens than vanilla speculative decoding, which results in better efficiency.
+Next, we theoretically analyze the approximation error of MTAD.
+Lemma 3.4 shows the upper bound of MTAD’s perplexity.
+And Theorem 3.5 shows the upper bound of the ratio between the perplexity of approximate MTAD and exact MTJD.
+The proofs are given in Appendix A.
+Lemma 3.4.
+Let us assume that when the small auxiliary model generates draft tokens with beam sampling, the beam width is large enough such that the returned log-likelihood is close to the maximum log-likelihood, i.e.,
+$\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))\geq(1-\epsilon)\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\max_{x_{\Gamma_{i-1}+1:\Gamma_{i}-1}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))$ (5)
+where $\epsilon$ is an error term and $\epsilon\leq 0$ because $\log q\leq 0$.
+Furthermore, let $H(p,q)$ the single-token cross entropy between $p$ and $q$, i.e., $H(p,q)=-\mathbb{E}_{x_{1:t\in\mathcal{X}}}\sum_{x_{t+1}}p(x_{t+1}|x_{1:t})\log q(x_{t+1}|x_{1:t})$.
+With the two assumption above, when $N\rightarrow\infty$ we have
+$PPL_{q}({x_{1:\Gamma_{N}}})\leq\exp(-\frac{1-\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{q}(\gamma-1,\arg\max\circ q)+\frac{H(p,q)}{\bar{\gamma}})$ (6)
+where
+$L_{q}(\gamma,\arg\max\circ q)=\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log q(x_{t+1:t+\gamma}|x_{1:t}))$ (7)
+Theorem 3.5.
+Let $x_{1:\Gamma_{N}}$ be the tokens generated by approximate MTAD, and $x^{*}_{1:\Gamma_{N}}$ be the tokens generated by deterministic exact MTJD.
+Assume $\forall x_{1:t}\in\mathcal{X}$, $\|\log p(x|x_{1:t})-\log q(x|x_{1:t})\|_{\infty}\leq U$, where $U$ is a constant.
+We have
+$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{p}(x^{*}_{1:\Gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{(1-\epsilon\bar{\gamma})H(p)+(1-\epsilon+\bar{\gamma})U}{\bar{\gamma}}\right)$ (8)
+where $H(p)$ is the entropy of $p$ and $\epsilon<0$ is the error term of beam sampling (see Lemma 3.4).
+Theorem 3.5 suggests the approximation error of MTAD is bounded by a factor related to the verification threshold $\tau$, average number of accepted tokens $\bar{\gamma}$, the difference between the large and small models (measured by $U$), the error of beam sampling $\epsilon$, and the entropy of the large model itself.
+In addition, the following theorem analyzes $\bar{\gamma}$.
+The proof is illustrated in Appendix A.
+Theorem 3.6.
+Following the assumption in Theorem 3.5, we have $\bar{\gamma}\geq\frac{|\log\tau|}{U}$.
+With Theorem 3.6, we observe that when $q\rightarrow p$, we have $U\rightarrow 0$ and $\bar{\gamma}\rightarrow\infty$.
+Meanwhile, when $\epsilon\rightarrow 0$, meaning the beam width for the auxiliary model is large enough, the ratio bound in Theorem 3.5 converges to 1, It implies that MTAD converges to MTJD under these limiting conditions.
+Similar to Spectr (Sun et al., 2023) and SpecInfer (Miao et al., 2023), it is possible to enhance the number of accepted tokens in MTAD by allowing the draft model to generate multiple draft sequences and applying tree-based attention (Miao et al., 2023) for simultaneous verification.
+However, our preliminary experiment results suggest that since MTAD already selects the longest accepted prefix sub-sequence, the advantage of generating multiple draft tokens is less significant.
+Moreover, this approach increases the memory cost during inference and may affect the error bounds derived above.
+Therefore, we leave a more detailed exploration of this extension as future work.
 4 Energy Efficiency Analysis
-Previous studies [20, 5, 17, 32] only focus on the speed of speculative decoding.
+Previous studies (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023) only focus on the speed of speculative decoding.
 However, an equally important consideration is energy consumption.
 To our knowledge, there is no existing work evaluating the impact of speculative decoding on inference energy consumption.
-Although MJSD and speculative decoding increase the number of FLOPs due to the involvement of an additional small model and the rollback operation, they concurrently reduce the inference time and memory operations, which are key factors of GPU energy consumption [1, 6].
+Although MTAD and speculative decoding increase the number of FLOPs due to the involvement of a small auxiliary model and the rollback operation, they concurrently reduce the inference time and memory operations, which are key factors of GPU energy consumption (Allen & Ge, 2016; Chen et al., 2011).
 Consequently, it poses an open question regarding whether speculative decoding increases or decreases overall energy consumption.
-To understand the net effect of speculative decoding, we decompose the total energy consumption into two parts following [1]:
+To understand the net effect of speculative decoding, we decompose the total energy consumption into two parts following (Allen & Ge, 2016):
 $E_{total}=PW_{flop}T_{flop}+PW_{mem}T_{mem}$ (9)
@@
 Meanwhile, $PW_{mem}$ is irrelevant to the input batch size, as it only depends on the memory hardware.
-To determine the relative magnitude relationship between $PW_{flop}$ and $PW_{mem}$, we first point out the fact that GPU memory operations in LLM inference are dominated by accessing off-chip global memory, which consumes about 100×100$\times$100 × of energy compared to accessing on-chip shared memory [16].
-It is because each multiprocessor on GPU usually has 64KB of on-chip memory shared by multiple threads, while storing a single layer of LLM, say T5-11b [25], requires about 1GB memory.
-Moreover, Allen and Ge showed that doing sequential read from off-chip memory consumes $20$-$30\%$ more power than running maximum FLOPs [1].
-So we have $PW_{mem}>PW^{*}_{flop}\geq PW_{flop}$.
-Notice that $PW^{*}_{flop}=PW_{flop}$ only if the batch size reaches the maximum parallelization capacity of GPUs.
-During greedy decoding and speculative decoding, the batch size is usually small [20].
-So most of the computing power is not utilized [20], which means $PW_{mem}\gg PW_{flop}$.
-In addition, previous studies have shown that during LLM inference $T_{mem}\gg T_{flop}$ [20].
-Therefore, the energy induced by memory operations, i.e., $PW_{mem}T_{mem}$ dominates $E_{total}$.
-Since speculative decoding reduces $T_{mem}$ by reducing the number of runs of the large model, it should reduce the inference energy consumption to a similar extent as it reduces time consumption.
-To validate our hypothesis, we conducted an experiment to evaluate how batch size influences energy consumption during inference.
-We ran OPT-13b models on 2 Nvidia A10 GPUs with 24GB memory.
-Fixing the total number of runs of the large model while varying the input batch size $b\in\{1,2,4,8,16\}$ for each run, we measured time and energy cost.
-The details of energy measurement are illustrated in the Appendix B.
-Table 1 shows the results.
-As batch size doubles, although the number of FLOPs doubles, the energy consumption per run increases slightly.
-This observation demonstrates that $PW_{mem}T_{mem}$ dominates $E_{total}$.
-In the next section, we will show more direct evidence for the energy efficiency of MJSD and other speculative decoding algorithms.
 Table 2: The effect of batch size to inference speed and energy consumption. The number of inputs is the product of the number of LLM runs and input batch size.
@@
 16 74,058 24.7 1.54 1,255 0.418 0.026
+Table 3: Speed and energy cost of multinomial sampling (ms) and speculative decoding (spec).
+OPT Llama-2
+ms spec ms Spec
+tokens/s 23.8 35.6 22.0 31.6
+J/token 11.3 5.74 11.2 6.97
+To determine the relative magnitude relationship between $PW_{flop}$ and $PW_{mem}$, we first point out the fact that GPU memory operations in LLM inference are dominated by accessing off-chip global memory, which consumes about 100×100$\times$100 × of energy compared to accessing on-chip shared memory (Jouppi et al., 2021).
+It is because each multiprocessor on GPU usually has 64KB of on-chip memory shared by multiple threads, while storing a single layer of LLM, say T5-11b (Raffel et al., 2020), requires about 1GB memory.
+Moreover, Allen and Ge showed that doing sequential read from off-chip memory consumes $20$-$30\%$ more power than running maximum FLOPs (Allen & Ge, 2016).
+So we have $PW_{mem}>PW^{*}_{flop}\geq PW_{flop}$.
+Notice that $PW^{*}_{flop}=PW_{flop}$ only if the batch size reaches the maximum parallelization capacity of GPUs.
+During multinomial sampling and speculative decoding, the batch size is usually small (Leviathan et al., 2023).
+So most of the computing power is not utilized (Leviathan et al., 2023), which means $PW_{mem}\gg PW_{flop}$.
+In addition, previous studies have shown that during LLM inference $T_{mem}\gg T_{flop}$ (Leviathan et al., 2023).
+Therefore, the energy induced by memory operations, i.e., $PW_{mem}T_{mem}$ dominates $E_{total}$.
+Since speculative decoding reduces $T_{mem}$ by reducing the number of runs of the large model, it should reduce the inference energy consumption to a similar extent as it reduces time consumption.
+To validate our hypothesis, we conducted an experiment to evaluate how batch size influences energy consumption during inference.
+We ran OPT-13b models on a Nvidia L40 GPUs with 48GB memory.
+Fixing the total number of runs of the large model while varying the input batch size $b\in\{1,2,4,8,16\}$ for each run, we measured time and energy cost.
+The details of energy measurement are illustrated in the Appendix D.
+Table 2 shows the results.
+As batch size doubles, although the number of FLOPs doubles, the energy consumption per run increases slightly.
+This observation demonstrates that $PW_{mem}T_{mem}$ dominates $E_{total}$.
+Moreover, we measured the speed and energy consumption of running multinomial sampling with the large model and speculative decoding using OPT (125M, 13B) and Llama-2 (68M, 13B) models.
+The results, shown in Table 3, indicate that speculative decoding reduces the energy consumption and the time cost.
+This observation corroborates our claim to the energy efficiency of speculative decoding.
 5 Experiments
-Datasets and Models.We use three public datasets for evaluation: (1) ChatGPT-Prompt [26], (2) ChatAlpaca [3], and (3) CNN Dailymail [30].
-The three datasets represent three common tasks for LLMs: instruction, multi-turn chat, and summarization.
+Datasets and Models.
+We use five public datasets for evaluation: (1) ChatGPT-Prompt (Rashad, 2023), (2) ChatAlpaca (Bian et al., 2023), (3) CNN Dailymail (See et al., 2017), (4) Spider (Yu et al., 2018), and (5) MT-bench (Zheng et al., 2023).
 Table 8 in the Appendix shows more details of the datasets.
-We use two pairs of small and large models in our experiments: OPT-125m and OPT-13B [38], and Llama-68M [24] and Llama-2-13B [34].
+Following previous studies (Kim et al., 2023), we use two public LLM families in our experiments: OPT (Zhang et al., 2022) and Llama-2 (Touvron et al., 2023).
+In this section, we set the large model to be OPT-13B and Llama-2-13B as they are the largest models that can run on a single 40GB GPU, and use Llama-68M (Miao et al., 2023) and OPT-125M as the small models.
+Appendix C reports additional experiment results with OPT-30B and Llama-2-chat-70B.
 Baselines.
-For each pair of small and large models, we compare our method with greedy decoding (greedy), and three speculative decoding methods: vanilla speculative decoding (speculative[19, 5], Spectr [32], and BiLD [17].
-Our implementation of MJSD and all the baselines are based on a public implementation of speculative decoding [2].
-We also tried to compare with SpecInfer [24] using their released implementation.
-But it runs out of memory on our machine.
+For each pair of small and large models, we compare our method with four speculative decoding methods: vanilla speculative decoding (speculative(Lee et al., 2018; Chen et al., 2023), Spectr (Sun et al., 2023), SpecInfer (Miao et al., 2023), and BiLD (Kim et al., 2023).
+Our implementation of MTAD and all the baselines are based on a public implementation of speculative decoding (Bear, 2024).
 For each method, we let it generate at most 128 tokens for each input and run it for $1,000$ seconds.
 We open-sourced our code for reproduction.
-More details of the hyper-parameters, warping operations, and machine configurations of the experiments can be found in the Appendix B, C, and D .
-Table 2: Inference efficiency and output perplexity of different methods on ChatGPT-Prompt (CP), ChatAlpaca (CA), and CNNDailyMail (CD) datasets.
+All the methods are stochastic with top-$k$ and top-$p$ sampling.
+The details of the hyper-parameters (e.g., $k$ and $p$) and machine configurations of the experiments can be found in the Appendix D, E, and F.
+Appendix C reports additional experiments and ablation studies.
+Table 4: Inference efficiency and output perplexity of different methods on ChatGPT-Prompt (CP), ChatAlpaca (CA), CNNDailyMail (CD), Spider (SP), and MT-Bench (MT) datasets.
 Bold numbers mark the best result, underlined numbers mark the second best.
@@
 […]
-speed up $\uparrow$ 1.00$\pm$0.00 1.63$\pm$0.02 1.52$\pm$0.04 1.99$\pm$0.06 2.78$\pm$0.01
-energy (J/token) $\downarrow$ 11.7$\pm$1.56 6.62$\pm$0.91 7.45$\pm$0.90 5.17$\pm$0.88 3.38$\pm$0.02
-perplexity $\downarrow$ 3.74$\pm$0.14 3.64$\pm$0.11 3.15$\pm$0.06 3.64$\pm$0.08 2.06$\pm$0.06
-OPT speed (token/s) $\uparrow$ 22.4$\pm$0.48 33.8$\pm$2.47 31.5$\pm$1.87 38.0$\pm$2.20 55.8$\pm$0.30
-speed $\uparrow$ 1.00$\pm$0.00 1.51$\pm$0.08 1.41$\pm$0.06 1.70$\pm$0.06 2.50$\pm$0.05
-energy (J/token) $\downarrow$ 13.2$\pm$0.28 7.48$\pm$0.07 8.75$\pm$0.13 6.08$\pm$0.11 3.61$\pm$0.03
-perplexity $\downarrow$ 5.49$\pm$0.15 5.47$\pm$0.11 4.51$\pm$0.09 5.27$\pm$0.09 3.00$\pm$0.09
-CA Llama-2 speed (token/s) $\uparrow$ 22.0$\pm$0.15 31.6$\pm$0.35 28.8$\pm$0.20 27.7$\pm$0.29 44.1$\pm$0.25
-speed up $\uparrow$ 1.00$\pm$0.00 1.43$\pm$0.02 1.31$\pm$0.01 1.26$\pm$0.00 2.00$\pm$0.02
-energy (J/token) $\downarrow$ 11.2$\pm$0.16 6.98$\pm$0.15 7.99$\pm$0.15 7.20$\pm$0.08 4.72$\pm$0.03
-perplexity $\downarrow$ 2.11$\pm$0.01 2.13$\pm$0.03 1.95$\pm$0.03 2.15$\pm$0.01 1.88$\pm$0.05
-OPT speed (token/s) $\uparrow$ 23.8$\pm$0.10 35.6$\pm$0.45 38.5$\pm$0.93 28.4$\pm$0.34 49.6$\pm$0.42
-speed $\uparrow$ 1.00$\pm$0.00 1.49$\pm$0.01 1.62$\pm$0.03 1.19$\pm$0.01 2.08$\pm$0.03
-energy (J/token) $\downarrow$ 11.3$\pm$0.22 5.74$\pm$0.11 5.12$\pm$0.06 6.24$\pm$0.11 4.03$\pm$0.02
-perplexity $\downarrow$ 3.28$\pm$0.06 3.32$\pm$0.10 2.60$\pm$0.06 3.16$\pm$0.06 2.07$\pm$0.03
-CD Llama-2 speed (token/s) $\uparrow$ 21.7$\pm$0.08 30.7$\pm$0.18 30.5$\pm$0.21 25.0$\pm$0.31 44.2$\pm$0.99
-speed up $\uparrow$ 1.00$\pm$0.00 1.41$\pm$0.00 1.41$\pm$0.01 1.15$\pm$0.01 2.04$\pm$0.05
-energy (J/token) $\downarrow$ 11.3$\pm$0.21 7.07$\pm$0.19 7.41$\pm$0.16 8.22$\pm$0.19 4.80$\pm$0.12
-perplexity $\downarrow$ 2.88$\pm$0.04 2.87$\pm$0.08 2.93$\pm$0.03 3.06$\pm$0.11 2.63$\pm$0.10
-OPT speed (token/s) $\uparrow$ 23.3$\pm$0.81 31.7$\pm$0.91 30.9$\pm$0.80 23.7$\pm$0.40 43.6$\pm$0.33
-speed $\uparrow$ 1.00$\pm$0.00 1.36$\pm$0.04 1.32$\pm$0.04 1.01$\pm$0.02 1.87$\pm$0.02
-energy (J/token) $\downarrow$ 11.5$\pm$0.14 6.37$\pm$0.11 6.71$\pm$0.17 7.31$\pm$0.17 4.86$\pm$0.03
-perplexity $\downarrow$ 3.93$\pm$0.14 3.97$\pm$0.06 3.74$\pm$0.09 4.04$\pm$0.07 3.17$\pm$0.06
+energy (J/token) $\downarrow$ 6.62$\pm$0.91 7.45$\pm$0.90 5.17$\pm$0.88 9.52$\pm$0.10 3.38$\pm$0.02
+perplexity $\downarrow$ 3.64$\pm$0.11 3.15$\pm$0.06 3.64$\pm$0.08 3.64$\pm$0.11 2.06$\pm$0.06
+OPT speed (token/s) $\uparrow$ 33.8$\pm$2.47 31.5$\pm$1.87 38.0$\pm$2.20 32.8$\pm$ 0.58 55.8$\pm$0.30
+energy (J/token) $\downarrow$ 7.48$\pm$0.07 8.75$\pm$0.13 6.08$\pm$0.11 10.3$\pm$1.49 3.61$\pm$0.03
+perplexity $\downarrow$ 5.47$\pm$0.11 4.51$\pm$0.09 5.27$\pm$0.09 5.12$\pm$0.01 3.00$\pm$0.09
+CA Llama-2 speed (token/s) $\uparrow$ 31.6$\pm$0.35 28.8$\pm$0.20 27.7$\pm$0.29 26.5$\pm$0.49 44.1$\pm$0.25
+energy (J/token) $\downarrow$ 6.98$\pm$0.15 7.99$\pm$0.15 7.20$\pm$0.08 7.52$\pm$0.32 4.72$\pm$0.03
+perplexity $\downarrow$ 2.13$\pm$0.03 1.95$\pm$0.03 2.15$\pm$0.01 2.15$\pm$0.01 1.88$\pm$0.05
+OPT speed (token/s) $\uparrow$ 35.6$\pm$0.45 38.5$\pm$0.93 28.4$\pm$0.34 31.4$\pm$0.39 49.6$\pm$0.42
+energy (J/token) $\downarrow$ 5.74$\pm$0.11 5.12$\pm$0.06 6.24$\pm$0.11 8.68$\pm$1.83 4.03$\pm$0.02
+perplexity $\downarrow$ 3.32$\pm$0.10 2.60$\pm$0.06 3.16$\pm$0.06 3.42$\pm$0.03 2.07$\pm$0.03
+CD Llama-2 speed (token/s) $\uparrow$ 30.7$\pm$0.18 30.5$\pm$0.21 25.0$\pm$0.31 24.6$\pm$0.06 44.2$\pm$0.99
+energy (J/token) $\downarrow$ 7.07$\pm$0.19 7.41$\pm$0.16 8.22$\pm$0.19 7.59$\pm$0.85 4.80$\pm$0.12
+perplexity $\downarrow$ 2.87$\pm$0.08 2.93$\pm$0.03 3.06$\pm$0.11 2.92$\pm$0.09 2.63$\pm$0.10
+OPT speed (token/s) $\uparrow$ 31.7$\pm$0.91 30.9$\pm$0.80 23.7$\pm$0.40 25.7$\pm$0.36 43.6$\pm$0.33
+energy (J/token) $\downarrow$ 6.37$\pm$0.11 6.71$\pm$0.17 7.31$\pm$0.17 8.03$\pm$0.63 4.86$\pm$0.03
+perplexity $\downarrow$ 3.97$\pm$0.06 3.74$\pm$0.09 4.04$\pm$0.07 3.92$\pm$ 0.34 3.17$\pm$0.06
+SP Llama-2 speed (token/s) $\uparrow$ 24.0$\pm$0.28 26.2$\pm$0.08 24.2$\pm$0.29 23.8$\pm$0.20 26.4$\pm$0.28
+energy (J/token) $\downarrow$ 10.75$\pm$0.02 9.84$\pm$0.07 11.0$\pm$0.08 11.0$\pm$0.76 9.01$\pm$0.07
+perplexity $\downarrow$ 2.26$\pm$0.01 2.13$\pm$0.03 2.29$\pm$0.04 2.29$\pm$0.03 1.87$\pm$0.03
+OPT speed (token/s) $\uparrow$ 24.6$\pm$0.30 29.9$\pm$0.55 19.8$\pm$0.13 24.1$\pm$0.10 34.4$\pm$0.46
+energy (J/token) $\downarrow$ 15.6$\pm$3.55 13.6$\pm$3.07 20.1$\pm$2.52 16.9$\pm$2.75 11.7$\pm$2.36
+perplexity $\downarrow$ 2.30$\pm$0.07 1.90$\pm$0.01 2.20$\pm$0.09 2.21$\pm$ 0.01 1.63$\pm$0.03
+MT Llama-2 speed (token/s) $\uparrow$ 23.0$\pm$1.10 23.7$\pm$1.43 19.1$\pm$2.71 23.7$\pm$2.03 29.4$\pm$2.71
+energy (J/token) $\downarrow$ 7.99$\pm$0.26 7.40$\pm$0.19 9.27$\pm$0.54 9.20$\pm$0.73 6.71$\pm$1.19
+perplexity $\downarrow$ 3.64$\pm$0.51 3.44$\pm$0.76 3.64$\pm$0.51 3.63$\pm$0.50 2.21$\pm$0.18
+OPT speed (token/s) $\uparrow$ 34.0$\pm$3.00 44.7$\pm$2.92 28.7$\pm$2.46 28.5$\pm$2.74 48.0$\pm$1.80
+energy (J/token) $\downarrow$ 12.1$\pm$0.36 6.23$\pm$0.67 12.9$\pm$1.73 13.2$\pm$1.88 6.11$\pm$0.82
+perplexity $\downarrow$ 2.02$\pm$0.40 1.50$\pm$0.27 1.97$\pm$0.38 1.99$\pm$ 0.33 1.10$\pm$0.03
 […]
 5.1 Comparison with Baselines
-Table 2 shows the main results of our experiments.
-The standard deviations in the table are computed by repeating the experiment four times.
-First, we compare the output perplexity of different algorithms.
-We can see that the perplexity of greedy decoding, vanilla speculative decoding, and Spectr decoding are close because they have identical sampling distributions for each token.
-Meanwhile, BiLD approximates the sampling distribution of greedy decoding but yields better perplexity.
-It is because we set a strict acceptance threshold for BiLD decoding, which lowers the acceptance rate but ensures every token has a high probability in the large model.
-It can be viewed as having an additional warping operation towards the sampling distribution.
-Nevertheless, there is a significant gap between the perplexity of MJSD and other baselines.
-On average, the output perplexity of MJSD is $21.2\%$ lower than that of BiLD.
-It provides strong evidence that MJSD has better effectiveness than existing speculative decoding methods that aim to have identical sampling distribution as greedy decoding.
-Next, we compare the speeds of different algorithms.
-While the speeds of all speculative decoding baselines are close, MJSD is significantly faster.
-On average, MJSD is 2.21×2.21$\times$2.21 × faster than greedy decoding, 1.49×1.49$\times$1.49 × faster than vanilla speculative decoding, 1.55×1.$5$ faster than BiLD, and 1.64×1.64$\times$1.64 × faster than Spectr.
-The speed improvement is because MJSD has a larger average acceptance length at each iteration, as MJSD accepts the longest sub-sequence that passes verification.
-Figure 3 shows the average acceptance length of vanilla speculative decoding and MJSD.
-We can see that the acceptance length of MJSD is on average 2.15×2.15$\times$2.15 × larger than that of vanilla speculative decoding.
-Finally, we focus on energy efficiency.
-We can see that all speculative decoding algorithms have significantly better energy efficiency over greedy decoding.
-It supports our analysis that speculative decoding can improve the LLM inference energy efficiency by reducing the number of memory operations.
-In addition, we can see that the energy efficiency of MJSD is significantly better than that of all baseline algorithms.
-We believe it is also because of the longer acceptance sequence at each iteration.
-The number of memory operations is proportional to the number of LLM runs, which is inversely proportional to the average acceptance length.
-Since by always accepting the longest prefix that passes verification, the energy consumption of MJSD is on average 1.62×1.62$\times$1.62 × smaller than that of vanilla speculative decoding, and 2.84×2.84$\times$2.84 × smaller than that of greedy decoding.
-Figure 3: Average number of accepted tokens per iteration of speculative decoding and MJSD. Each iteration generates 4 draft tokens.
+Table 4 shows the primary results of our experiments while Table 5 details the block efficiency, defined as the average number of tokens generated per iteration, of different methods.
+The standard deviations in the tables are computed by repeating each experiment four times [^4] For MT-Bench, the standard deviation also accounts for variations across different tasks..
+First, we observe that MTAD is significantly more efficient than all baselines in terms of both energy and time.
+On one hand, the energy consumption of MTAD is on average 1.54×1.54$\times$1.54 × smaller than that of vanilla speculative decoding.
+On the other hand, MTAD is 1.101.71×1.10-1.71$\times$1.10 - 1.71 × faster than vanilla speculative decoding, 1.38×1.38$\times$1.38 × faster than BiLD, 1.59×1.59$\times$1.59 × faster than Spectr, and 1.60×1.60$\times$1.60 × faster than SpecInfer.
+SpecInfer and Spectr has better block efficiency than vanilla speculative decoding, but are slower.
+This may be due to the fact that they have to verify multiple draft sequences, which introduces extra overhead and may not be perfectly parallelized, especially when the memory overhead exceeds the GPU memory capacity.
+Meanwhile, MTAD has the best block efficiency without causing any extra overhead, hence it is significantly more efficient.
+Table 5: Average number of tokens generated at each iteration across all datasets.
+Llama-2 OPT
+spec 2.02$\pm$0.05 2.60$\pm$0.06
+BiLD 1.83$\pm$0.10 2.68$\pm$0.36
+Spectr 2.73$\pm$0.43 3.45$\pm$0.42
+SpecInfer 2.74$\pm$0.46 3.45$\pm$0.40
+MTAD 3.17$\pm$0.43 4.30$\pm$0.03
+Next, we compare the output perplexity of different algorithms.
+The perplexity scores of vanilla speculative decoding, SpecInfer, and Spectr are close since their sampling distributions are equivalent.
+Meanwhile, BiLD approximates the sampling distribution of single-token multinomial sampling but yields better perplexity.
+It is because we set a strict acceptance threshold for BiLD, which lowers the acceptance rate but ensures every token has a high probability in the large model, thus improving the overall likelihood.
+More importantly, there is a significant gap between MTAD and other baselines.
+On average, the perplexity of MTAD is $21.2\%$ lower than that of speculative decoding.
+Table 6: Downstream task scores of speculative decoding and MTAD. All the scores are higher the better.
+spec MTAD
+CD Rouge-L 0.114 0.118
+SP EA 11.5 13.0
+MT Humanities 2.95 3.15
+Extraction 1.80 2.50
+Roleplay 3.10 3.80
+Math 1.10 1.00
+Coding 1.25 1.10
+Reasoning 3.80 3.15
+STEM 2.85 3.10
+Writing 3.80 3.65
+Average 2.58 2.68
+In addition, to show MTAD indeed improves the downstream effectiveness, we compare the performance metrics of speculative decoding and MTAD on CNNDM, Spider, and MT-Bench datasets.
+[^5] All speculative baselines are equivalent to multinomial sampling We exclude the other two datasets as they lack explicit downstream metrics.
+And we exclude the results of OPT models due to their consistently poor performance across all evaluated datasets.
+As illustrated in Table 6, MTAD outperforms speculative decoding across all three datasets, thereby validating our claim that MTAD achieves superior effectiveness compared to conventional decoding methods that rely on single-token distributions.
 5.2 Ablation Study
+5.2.1 Number of Beams
+First, we investigate how the number of beams used in the beam decoding of the small model affects the inference performance.
+Table 7 shows the results.
+Increasing the number of beams improves the quality of the draft tokens, which not only improves the output perplexity but also increases the average acceptance length and hence leads to better efficiency.
+But we can see that the increment slows down when the number of beams is large enough.
+In addition, when the number of beams is too large, the inference cost of the small model will become too high.
 Table 7: Effect of number of beams to the inference performance on ChatGPT-Prompts dataset.
@@
 perplexity $\downarrow$ 3.63 3.16 3.42 3.19
-5.2.1 Effects of Number of Beams
-First, we investigate how the number of beams used in the beam decoding of the small model affects the inference performance.
-Table 3 shows the results.
-Increasing the number of beams improves the quality of the draft tokens, which not only improves the output perplexity but also increases the average acceptance length and hence leads to better efficiency.
-But we can see that the increment slows down when the number of beams is large enough.
-In addition, when the number of beams is too large, the inference cost of the small model will become too high.
-5.2.2 Effects of Acceptance Threshold
-Figure 4: Effect of acceptance threshold on output perplexity and decoding speed on ChatGPT-Prompts dataset.
+5.2.2 Acceptance Threshold
+Figure 3: Effect of acceptance threshold on output perplexity and decoding speed.
 Next, we evaluate the effect of acceptance threshold $\tau$.
@@
 However, the perplexity increases when $\tau$ is close to 1.
-When $\tau=1$, all the draft tokens are rejected, which makes MJSD equivalent to greedy decoding.
-Similarly, when $\tau$ is close to 1, the advantage of multi-token joint greedy decoding on effectiveness disappears, hence the perplexity becomes close to the perplexity of greedy decoding.
-Another surprising observation is that the perplexity of MJSD is good when $\tau=0$.
-When $\tau=0$, MJSD is equivalent to generating $\gamma$ tokens using beam decoding with the small model, then generating an additional token with the large model.
-The fact that MJSD achieves good perplexity when $\tau=0$ can be explained by the fact that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s predictions deviate from those of the large model” [17, 20].
-Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MJSD is relatively stable, suggesting that MJSD is not sensitive to the acceptance threshold.
+When $\tau=1$, all the draft tokens are rejected, which makes MTAD equivalent to multinomial sampling.
+Similarly, when $\tau$ is close to 1, the advantage of multi-token joint decoding on effectiveness disappears, hence the perplexity becomes close to the perplexity of multinomial sampling.
+Another surprising observation is that the perplexity of MTAD is good when $\tau=0$.
+When $\tau=0$, MTAD is equivalent to generating $\gamma$ tokens using beam decoding with the small model, then generating an additional token with the large model.
+The fact that MTAD achieves good perplexity when $\tau=0$ can be explained by the fact that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s predictions deviate from those of the large model” (Kim et al., 2023; Leviathan et al., 2023).
+Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MTAD is relatively stable, suggesting that MTAD is not sensitive to the acceptance threshold.
 6 Related Work
@@
 There are extensive studies on improving large model inference efficiency.
-Well-known methods include model quantization [9, 22], model pruning [10, 28], and model distillation [15].
+Well-known methods include model quantization (Frantar et al., 2022; Lin et al., 2023), model pruning (Gale et al., 2019; Sanh et al., 2020), and model distillation (Hinton et al., 2015).
 Despite achieving significant speed-ups, a common drawback of these methods is that they have to sacrifice the model’s effectiveness.
 A direction closer to our work is non-autoregressive decoding.
-It is first proposed by [12] to generate multiple tokens in parallel.
+It is first proposed by (Gu et al., 2017) to generate multiple tokens in parallel.
 That is, the model simultaneously predicts $p(x_{t+k}|x_{1:t})$ ($k=1,2,\ldots$).
-Subsequent studies further improved the performance of parallel decoding by incorporating additional information [35, 31, 21] or using additional iterations to refine predictions [11, 19, 13].
-However, these works require continuous training of the model and usually either compromise the model effectiveness or require task-dependent techniques to achieve comparable performance [17].
+Subsequent studies further improved the performance of parallel decoding by incorporating additional information (Wang et al., 2019; Sun et al., 2019; Li et al., 2019) or using additional iterations to refine predictions (Ghazvininejad et al., 2019; Lee et al., 2018; Guo et al., 2020).
+However, these works require continuous training of the model and usually either compromise the model effectiveness or require task-dependent techniques to achieve comparable performance (Kim et al., 2023).
 Speculative Decoding.
-Speculative decoding was recently proposed in [20, 5] as a way to accelerate LLM inference.
-Spectr [32] enhances speculative decoding by letting the small model generate multiple i.i.d.
+Speculative decoding was recently proposed in (Leviathan et al., 2023; Chen et al., 2023) as a way to accelerate LLM inference.
+Spectr (Sun et al., 2023) enhances speculative decoding by letting the small model generate multiple i.i.d.
 draft sequences.
-While speculative decoding and Spectr use the large model to verify all the tokens drafted by the small model, BiLD [17] only calls the large model when the probability output by the small model is below a pre-defined threshold $\tau_{1}$.
+While speculative decoding and Spectr use the large model to verify all the tokens drafted by the small model, BiLD (Kim et al., 2023) only calls the large model when the probability output by the small model is below a pre-defined threshold $\tau_{1}$.
 The large model rejects a token if its negative log-likelihood is larger than threshold $\tau_{2}$.
-SpecInfer [24] uses one or multiple small models to generate a draft token tree to increase the average acceptance length for each iteration.
+SpecInfer (Miao et al., 2023) uses one or multiple small models to generate a draft token tree to increase the average acceptance length for each iteration.
 All these methods can be perceived as exact or approximate versions of sampling tokens from the conditional distribution $p(x_{t}|x_{<t})$.
@@
 It is obvious that if more draft tokens are accepted, the overall inference speed will increase.
-BiLD [17] uses a model prediction alignment technique to better train the small model.
+BiLD (Kim et al., 2023) uses a model prediction alignment technique to better train the small model.
 Liu et al.
-[23] propose online speculative decoding to continually update the draft model based on observed input data.
-Instead, Rest [14] uses a retrieval model to produce draft tokens.
-The improvement of small models can benefit all speculative decoding algorithms, including our method.
+(Liu et al., 2023) propose online speculative decoding to continually update the draft model based on observed input data.
+Instead, Rest (He et al., 2023) uses a retrieval model to produce draft tokens.
+An alternative way is to train additional heads in the large model to predict future tokens.
+Representative works include EAGLE (Li et al., 2024) and MEDUSA (Cai et al., 2024).
+Importantly, these works are orthogonal to speculative decoding techniques, including our proposed method.
+This orthogonality means that the improvements offered by more accurate draft tokens could be combined with our method for better effectiveness.
 7 Conclusion
-We introduce multi-token joint speculative decoding that significantly enhances output quality with better time and energy efficiency.
+We introduce multi-token assisted decoding that significantly enhances output quality along with better time and energy efficiency.
 A distinctive aspect of our work is the exploration of speculative decoding’s impact on inference energy consumption, an often neglected area in existing studies.
-Experiment results demonstrate our method achieves significant energy reduction.
 This research contributes not only a novel decoding approach but also valuable insights for optimizing LLM deployment in real-world applications where considerations of both quality and efficiency are crucial.
-8 Limitations and Impact Statements
-Limitations.
-This paper mainly considers improving the output perplexity of decoding algorithms.
-Under the assumption that the model is well-trained for downstream tasks, improving output perplexity usually leads to better downstream effectiveness.
-But if the model is not well-trained, improving the perplexity may not necessarily improve the downstream effectiveness.
-However, it is mainly the problem of the model itself.
-In real-world applications, it is reasonable to assume the model is properly trained for downstream tasks.
-Impact Statements.
-The goal of this work is to advance the field of Large Language Model (LLM) Inference, which has received tremendous attention from both academia and industry.
-However, LLMs also have received many critiques, including their extremely large carbon footage emission during both training and inference.
-Our work pays special attention to their energy consumption during inference to provide high-quality and fast-inference LLMs with reduced energy consumption, which has become a serious concern with the rapid increase of LLM-related workloads in the past few years.
-Appendix A Pseudocode of MJSD
+Appendix A Proof
+A.1 Proof of Theorem 3.2
+Proof.
+$PPL(x_{1:\Gamma_{N}})$ $=\exp\left(-\frac{1}{\Gamma_{N}}\sum_{i=1}^{\Gamma_{N}}\log p(x_{i}|x_{1:i-1})\right)$ (10)
+$=\exp\left(-\frac{N}{\Gamma_{N}}\frac{1}{N}\sum_{i=1}^{N}\log p(x_{\Gamma_{i-1}:\Gamma_{i}}|x_{1:\Gamma_{i-1}})\right)$
+When $N\rightarrow\infty$, $\frac{\Gamma_{N}}{N}\rightarrow\bar{\gamma}$, and $\frac{1}{N}\sum_{i=1}^{N}\log p(x_{\Gamma_{i-1}:\Gamma_{i}}|x_{1:\Gamma_{i-1}})\rightarrow\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{\gamma}\sum_{x_{t+1:t+\gamma}}P(\gamma)\tilde{p}(x_{t+1:t+\gamma}|x_{1:t})\log p(x_{t+1:t+\gamma}|x_{1:t})=\mathbb{E}_{\gamma}L_{p}(\gamma,\tilde{p})$
+A.2 Proof of Corollary 3.3
+Proof.
+For deterministic multi token sampling, $\tilde{p}_{multi}=\arg\max\circ p$, so we have
+$L_{p}(\gamma,\tilde{p}_{multi})=E_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})$ (11)
+Notice that deterministic greedy sampling can be seen as a special case of MJGD where $\tilde{p}_{single}(x_{t+1:t+\gamma}|x_{1:t})=1$ if and only if $x_{t+i}=\arg\max_{x}p(x|x_{1:t+i-1})$ for $i=1,\ldots,\gamma$.
+Let $x^{*}_{t+1:t+\gamma}$ be the tokens generated by deterministic MJGD and let $x^{\prime}_{t+1:t+\gamma}$ be the tokens generated by deterministic greedy decoding.
+For any fixed $\gamma$ and $x_{1:t}$, we have $\log p(x^{\prime}_{t+1:t+\gamma}|x_{1:t})\leq\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})=\log p(x^{*}_{t+1:t+\gamma}|x_{1:t})$.
+Therefore, $L_{p}(\gamma,\tilde{p}_{single})\leq L_{p}(\gamma,\tilde{p}_{multi})$.
+Then with Theorem 3.2, we know that the perplexity of greedy decoding will be higher.
+∎
+A.3 Proof of Lemma 3.4
+We first prove the following Lemma.
+Lemma A.1.
+Let $PPL_{p}$ and $PPL_{q}$ denote the perplexity of tokens under distribution $p$ and $q$. When $N\rightarrow\infty$, we have
+$\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{q}(x_{1:\Gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}$ (12)
+where $\tau$ is the verification threshold.
+Proof.
+In the $i$-th iteration, the first $\gamma_{i}-1$ tokens are the accepted draft tokens and the last token is sampled from $p$.
+Based on our verification criteria, we know that for the accepted draft tokens, we have
+$\frac{p(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}{q(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}\geq\tau.$ (13)
+So,
+$\frac{p({x_{1:\Gamma_{N}}})}{q({x_{1:\Gamma_{N}}})}\geq\tau^{N}\prod_{i=1}^{N}\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}$ (14)
+Notice that
+$\left(\prod_{i=1}^{N}\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)^{\frac{1}{N}}=\exp\left(\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)\right)$ (15)
+When $N\rightarrow\infty$, since the last token at each iteration is sampled from $p$, we have
+$\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)\rightarrow\mathbb{E}_{p}\log\left(\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)=KL(p,q)\geq 0$ (16)
+So
+$\left(\prod_{i=1}^{N}\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)^{\frac{1}{N}}\geq 1$ (17)
+Therefore,
+$\frac{p({x_{1:\Gamma_{N}}})}{q({x_{1:\Gamma_{N}}})}\geq\tau^{N}$ (18)
+Thus,
+$\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{q}(x_{1:\Gamma_{N}})}=\left(\frac{p({x_{1:\Gamma_{N}}})}{q({x_{1:\Gamma_{N}}})}\right)^{-\frac{1}{\Gamma_{N}}}\leq\tau^{-\frac{N}{\Gamma_{N}}}\rightarrow\tau^{-\frac{1}{\bar{\gamma}}}$ (19)
+Now, we prove Lemma 3.4.
+Proof.
+$-\log PPL_{q}(x_{1:\Gamma_{N}})=\frac{1}{\Gamma_{N}}\sum_{i=1}^{N}(\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}})+\log q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1}))\$ (20)
+When $N\rightarrow\infty$, since the first $\gamma_{i}-1$ tokens are sampled with beam decoding, we have
+$\frac{1}{N}\sum_{i=1}^{N}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))$ $\rightarrow\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t\in\mathcal{X}}}\log q(x_{t+1:t+\gamma-1|x_{1:t}})$ (21)
+$\geq(1-\epsilon)\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\max_{x_{\Gamma_{i-1}+1:\Gamma_{i}-1}}q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))$
+$=(1-\epsilon)\mathbb{E}_{\gamma}L_{q}(\gamma-1,\arg\max\circ q)$
+Since the last token at each iteration is sampled from $p$, we have
+$\frac{1}{N}\sum_{i=1}^{N}\log q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1}))\rightarrow\mathbb{E}_{x_{1:t}\in\mathcal{X}}\mathbb{E}_{p}\log q(x_{t+1}|x_{1:t})=-H(p,q)$ (22)
+So
+$-\log PPL_{q}(x_{1:\Gamma_{N}})\geq\frac{1-\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma,x_{1:\Gamma_{i-1}}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}q(x_{t+1:t+\gamma}|x_{1:t}))-\frac{H(p,q)}{\bar{\gamma}}$ (23)
+$PPL_{q}(x_{1:\Gamma_{N}})\leq\exp\left(\frac{H(p,q)}{\bar{\gamma}}-\frac{1-\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{q}(\gamma-1,\arg\max\circ q)\right)$ (24)
+A.4 Proof of Theorem 3.5
+Proof.
+We have
+[…]
+[…]
+[…]
+$=\tau^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{H(p,q)}{\bar{\gamma}}-\frac{1-\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{q}(\gamma-1,\arg\max\circ q)+\frac{1}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)\right)$
+Notice that $L_{p}(\gamma,\arg\max\circ p)\geq L_{p}(\gamma+1,\arg\max\circ p)$ for any $\gamma$.
+This is because for any $x_{1:t}$, $\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})\geq\max_{x_{t+1:t+\gamma+1}}\left(\log p(x_{t+1:t+\gamma}|x_{1:t})+\log p(x_{t+\gamma+1}|x_{1:t+\gamma})\right)=\max_{x_{t+1:t+\gamma+1}}\log p(x_{t+1:t+\gamma+1}|x_{1:t})$.
+So
+$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{p}(x^{*}_{1:\Gamma_{N}})}$ (26)
+$\leq$ $\tau^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{H(p,q)}{\bar{\gamma}}+\frac{\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)+\frac{1-\epsilon}{\bar{\gamma}}(\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)-\mathbb{E}_{\gamma}L_{q}(\gamma,\arg\max\circ q))\right)$
+Since $\epsilon\leq 0$, and $L_{p}(\gamma,\arg\max\circ p)$ is the maximum log-likelihood, which is larger than the expected log-likelihood (i.e., negative entropy), we have
+$\frac{\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)$ (27)
+$=$ $\frac{\epsilon}{\bar{\gamma}}{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})$
+$\leq$ $\frac{\epsilon}{\bar{\gamma}}{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{x_{t+1:t+\gamma}}p(x_{t+1:t+\gamma}|x_{1:t})\log p(x_{t+1:t+\gamma}|x_{1:t})$
+$=$ $-\epsilon H(p)$
+In addition
+$\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)-\mathbb{E}_{\gamma}L_{q}(\gamma,\arg\max\circ q)$ (28)
+$=$ $\mathbb{E}_{\gamma}(L_{p}(\gamma,\arg\max\circ p)-L_{q}(\gamma,\arg\max\circ q))$
+$=$ $\mathbb{E}_{\gamma}\left(\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})-\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log q(x_{t+1:t+\gamma}|x_{1:t})\right)$
+$=$ $\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\left(\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})-\max_{x_{t+1:t+\gamma}}\log q(x_{t+1:t+\gamma}|x_{1:t})\right)$
+$\leq\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\left(\log p(x_{t+1:t+\gamma}|x_{1:t})-\log q(x_{t+1:t+\gamma}|x_{1:t})\right)$
+$=\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\left(\sum_{i=1}^{\gamma}\log p(x_{t+i}|x_{1:t+i-1})-\log q(x_{t+i}|x_{1:t+i-1})\right)$
+$\leq\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}U\gamma\quad(because\ \|\log p(x|x_{1:t})-\log q(x|x_{1:t})\|_{\infty}\leq U)$
+$=U\bar{\gamma}$
+And $H(p,q)=H(p)+KL(p\|q)$.
+$KL(p\|q)$ $=\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{x}p(x|x_{1:t})(\log p(x|x_{1:t})-\log p(x|x_{1:t}))$ (29)
+$\leq\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{x}p(x|x_{1:t})U\leq U$
+So $H(p,q)\leq H(p)+U$. Therefore,
+$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{p}(x^{*}_{1:\Gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{(1-\epsilon\bar{\gamma})H(p)+(1-\epsilon+\bar{\gamma})U}{\bar{\gamma}}\right)$ (30)
+A.5 Proof of Theorem 3.6
+Proof.
+Recall that we accept $x_{t+1:t+j}$ if and only if $\log p(x_{t+1:t+j}|x_{1:t})-\log q(x_{t+1:t+j}|x_{1:t})\geq\log\tau$. Since $\|\log p(x|x_{1:t})-\log q(x|x_{1:t})\|_{\infty}\leq U$, we have
+$\log p(x_{t+1:t+j}|x_{1:t})-\log q(x_{t+1:t+j}|x_{1:t})\geq-jU$ (31)
+Therefore $x_{t+1:t+j}$ is always accepted if $j\leq\frac{|\log\tau|}{U}$. So $\bar{\gamma}\geq\frac{|\log\tau|}{U}$
+Appendix B Pseudocode of MJSD
 See Algorithm 1.
-Algorithm 1 One Iteration of MJSD Algorithm
+Algorithm 1 One Iteration of MTAD Algorithm
 1:Input: draft model $M_{q}$, target model $M_{p}$, $input$, threshold $\tau$
@@
 Table 8: Dataset Statistics
-Dataset Task # Test Input Avg. Input Len
-ChatGPT-Prompt Instruction 360 25.2
-ChatAlpaca Chat 43.081 277.7
-CNNDM Summarization 11,490 3,967.1
-Appendix B Energy Consumption Measurement
+Dataset Task Avg. Input Len
+ChatGPT-Prompt Instruction 25.2
+ChatAlpaca Chat 277.7
+CNNDM Summarization 3,967.1
+Spider Text-to-SQL 347.68
+MT-Bench Various [^1] The tasks of MT-Bench cover humanities, extraction, roleplay, math, coding, reasoning, stem, writing, and STEM.
+N/A [^2] MT-Bench contains multi-turn tasks where the input includes the responses of LLMs, so the input length is not fixed.
+Appendix C Additional Experiments
+C.1 Resutlts with OPT-30B and Llama-2-70B
+Here we report the performances of different methods for OPT (350M and 30B) and Llama-2-Chat (7B and 70B).
+Table 9 shows the average performances across all datasets.
+MTAD always achieves the lowest perplexity and the best efficiency.
+Table 9: Inference efficiency and output perplexity of different methods with OPT (350M,30B) and Llama-2-Chat (7B,70B).
+The mean and standard deviation are computed across all datasets.
+Bold numbers mark the best result, underlined numbers mark the second best.
+[…]
+[…]
+energy (J/token) $\downarrow$ 138$\pm$87.7 142$\pm$99.7 122$\pm$66.4 125$\pm$65.4 119$\pm$67.7
+perplexity $\downarrow$ 1.77$\pm$0.22 1.69$\pm$0.25 1.73$\pm$0.24 1.73$\pm$0.24 1.52$\pm$0.19
+OPT speed (token/s) $\uparrow$ 15.3$\pm$1.64 14.5$\pm$1.96 17.0$\pm$4.14 17.4$\pm$4.00 19.5$\pm$4.11
+energy (J/token) $\downarrow$ 72.4$\pm$11.5 79.6$\pm$3.03 68.2$\pm$16.7 62.4$\pm$10.3 60.0$\pm$12.8
+perplexity $\downarrow$ 4.74$\pm$1.96 3.50$\pm$1.42 4.55$\pm$1.93 4.49$\pm$ 1.95 2.74$\pm$0.87
+[…]
+C.2 Ablation study of top-k and top-p sampling
+Here we conduct an ablation study to show how the value of $k$ and $p$ in top-k and top-p warping affects our method.
+Table 10 shows the results.
+We can see that when changing the value of $k$ and $p$, MTAD consistently achieves significantly better performances.
+Table 10: Ablation study of $k$ and $p$ in top-k and top-p sampling
+K P Greedy Speculative MJSD
+PPL Tokens/sec PPL Tokens/sec PPL Tokens/sec
+20 0.9 3.74 22.6 3.64 36.8 2.06 63.0
+20 0.8 3.06 22.7 3.10 38.5 1.93 58.8
+10 0.9 3.03 22.7 3.22 38.5 1.95 62.5
+10 0.8 2.56 22.7 2.53 40.0 1.80 62.5
+C.3 Additional Experiments on CNNDM and Spider
+We also report the downstream effectiveness of our method on CNNDM and Spider when the small model and large model are fine-tuned on the dataset.
+Table 11 and Table 12 show the results.
+We can see that MTAD consistenly achieve better effectiveness as well as faster decoding speed.
+Table 11: Comparison of ROUGE-L Scores and Tokens per Second under Different Fine-Tuning Conditions on CNNDM
+No Fine-Tune Fine-Tune 68M Fine-Tune Both
+Method ROUGE-L Tokens/sec ROUGE-L Tokens/sec ROUGE-L Tokens/sec
+Speculative 0.114 37.7 0.114 20.4 0.164 24.3
+MJSD 0.118 44.2 0.121 25.0 0.168 27.1
+Table 12: Comparison of Execution Accuracy (EA) and Tokens per Second under Different Fine-Tuning Conditions on Spider
+No Fine-Tune Fine-Tune 68M Fine-Tune Both
+Method EA Tokens/sec EA Tokens/sec EA Tokens/sec
+Speculative 11.5 28.5 11.5 27.8 16.3 25.6
+MJSD 13.0 30.3 14.8 32.3 18.3 29.4
+Appendix D Energy Consumption Measurement
 We use the command "nvidia-smi --query-gpu=power.draw --format=csv" to get GPU power every second, and sum them up as the total energy consumption.
 We use average energy consumption per token to measure energy efficiency.
-There is a recent study pointing out the measurement error using nvidia-smi [37].
-We follow the three principles proposed in [37] to minimize the error.
-Appendix C Configuration
-The experiments are conducted on a machine with 1 Nvidia L40 GPU (48 GB), 4 CPUs, and 50 GB main memory, using a batch size of 1, which is common for online serving [29].
+There is a recent study pointing out the measurement error using nvidia-smi (Yang et al., 2023).
+We follow the three principles proposed in (Yang et al., 2023) to minimize the error.
+Appendix E Configuration
+The experiments are conducted on a machine with 1 Nvidia L40 GPU (48 GB), 4 CPUs, and 50 GB main memory, using a batch size of 1, which is common for online serving (Schuster et al., 2022).
 We set the maximum running time to be an hour for each baseline.
@@
 The details of energy measurement are illustrated in the Appendix.
-Appendix D Hyper-parameter Details
-In the experiments, we follow the default settings in [2] to warp the sampling distribution $p$ and $q$ with the following steps, which are the default warpping operations in a public speculative decoding implementation.
+Appendix F Hyper-parameter Details
+In the experiments, we follow the default settings in (Bear, 2024) to warp the sampling distribution $p$ and $q$ with the following steps, which are the default warpping operations in a public speculative decoding implementation.
   1. 1.
@@
 Table 13 shows the hyper-parameters used in the experiments.
-Table 5: Hyper-parameters of different methods for different models and datasets. CP: ChatGPT-Prompts, CA: ChatAlpaca, CD: CNNDaily.
-Llama,CP OPT,CP Llama,CA OPT,CA Llama,D OPT,CD
-speculative step len $\gamma$ 4 4 4 4 4 4
-Spectr step len $\gamma$ 4 4 4 4 4 4
-num seq $m$ 4 4 2 2 4 2
-BiLD step len $\gamma$ 10 10 10 10 10 10
-fallback thres $\tau_{1}$ 0.9 0.9 0.9 0.3 0.9 0.3
-rollback thres $\tau_{2}$ 2 2 1 2 3 2
-MJSD step len $\gamma$ 4 4 4 4 4 4
-num beams 8 8 8 8 8 8
-acc/rej thres $\tau$ 0.1 0.1 0.1 0.1 0.1 0.1
-Appendix E License of Datasets and Models
+Table 13: Hyper-parameters of different methods for different models and datasets. L: Llama, O:OPT, CP: ChatGPT-Prompts, CA: ChatAlpaca, CD: CNNDaily.
+L,CP O,CP L,CA O,CA L,CD O,CD L,SP O,SP L,MT O,MT
+speculative step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
+Spectr&SpecInfer step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
+num seq $m$ 4 4 2 2 4 2 2 2 2 2
+BiLD step len $\gamma$ 10 10 10 10 10 10 10 10 10 10
+fallback thres $\tau_{1}$ 0.9 0.9 0.9 0.3 0.9 0.3 0.9 0.9 0.9 0.9
+rollback thres $\tau_{2}$ 2 2 1 2 3 2 1 1 1 1
+MTAD step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
+num beams 8 8 8 8 8 8 8 8 8 8
+acc/rej thres $\tau$ 0.1 0.1 0.1 0.1 0.1 0.1 0.9 0.1 0.9 0.1
+Appendix G License of Datasets and Models
 Datasets:
@@
 Codes
-Appendix F More Discussion on Multi-Token Joint Greedy Decoding
-It is not trivial to prove that MJGD is better than vanilla greedy decoding.
-We might tend to show that given the same $x_{1:t}$ and $prefix$, the $\gamma_{i}$ tokens generated by MJGD has higher likelihood than the $\gamma_{i}$ tokens generated by calling greedy decoding $\gamma_{i}$ times.
-However, at step $i$ ($\forall i\geq 2$), the previously generated tokens $x_{1:t}$ are different for the two decoding algorithms, making it difficult to compare the overall perplexity.
-To explain why MJGD is more effective, let us consider an optimal algorithm that generates next $\gamma_{i}$ tokens at one step based on a score function:
-$s^{*}(x_{t+1:t+\gamma_{i}})=\max_{x_{t+\gamma_{i}+1:N}}\log p(x_{t+1:N}|x_{1:t})$ (3)
-which is the log-likelihood of the optimal future sequence with prefix $x_{1:t}$ and the generated $\gamma_{i}$ tokens.
-The optimal greedy decoding chooses $x_{t+1:t+\gamma_{i}}$ that maximizes the score function and is guaranteed to return the sequence with the optimal perplexity.
-Both MJGD and vanilla greedy decoding can be viewed as approximations to the optimal greedy decoding by replacing $s^{*}(\cdot)$ with $s_{MJGD{}}$ or $s_{GD}$ that estimates the values of $s^{*}(\cdot)$.
-Specifically, since MJGD selects the $\gamma_{i}$ tokens with the largest joint likelihood, we have $s_{MJGD}(\cdot)=\log p(x_{1:t+\gamma_{i}}|x_{1:t})+c_{1}$ where $c_{1}$ is an arbitrary constant.
-Notice that adding a constant to $s(\cdot)$ changes the error of the estimation but does not change the behavior of the algorithm.
-On the other hand, greedy decoding generates the $\gamma_{i}$ tokens one by one.
-At the last token, greedy decoding selects the token with the largest $p(x_{t+\gamma_{i}}|x_{1:t+\gamma_{i}-1})$, which is equivalent to selecting the tokens with the largest $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ provided that $x_{t+1:t+\gamma_{i}-1}$ are the same tokens generated by previous steps.
-So for greedy decoding, $s_{GD}(x_{t+1:t+\gamma_{i}})=\log p(x_{1:t+\gamma_{i}}|x_{1:t})+c_{2}$ ($c_{2}$ is also an arbitrary constant) if and only if $x_{t+i}=\arg\max_{x}p(x|x_{1:t+i-1})$ for $i=1,\ldots,\gamma_{i}-1$.
-Otherwise, the score of $x_{t+1:t+\gamma_{i}}$ has to be smaller than any possible value of $\log p(x_{1:t+\gamma_{i}}|x_{1:t})+c_{2}$.
-Notice that the lower bound of $\log p(x_{1:t+\gamma_{i}}|x_{1:t})$ is $\infty$- ∞, so $s_{GD}(x_{t+1:t+\gamma_{i}})=-\infty$ if and only if $\exists j=1,\ldots,\gamma_{i}-1$ such that $x_{t+j}\neq\arg\max_{x}p(x|x_{1:t+i-1})$.
-Let $c^{*}=\arg\min_{c_{2}}\sum_{x_{t+1:t+\gamma_{i}}}|s_{GD}(x_{t+1:t+\gamma_{i}})-s^{*}(x_{t+1:t+\gamma_{i}})|$.
-By setting $c_{1}=c^{*}$, for $x_{t+1},\ldots,x_{t+\gamma_{i}-1}$ satisfying that $x_{t+i}=\arg\max_{x}p(x|x_{1:t+i-1})$, $\forall i=1,\ldots,\gamma_{i}-1$, we have
-$s_{MJGD}(x_{t+1:t+\gamma_{i}})=s_{GD}(x_{t+1:t+\gamma_{i}})=\log p(x_{1:t+\gamma_{i}|x_{1:t}})+c^{*}$
-Otherwise, we have
-$|s_{MJGD}(x_{t+1:t+\gamma_{i}})-s^{*}(x_{t+1:t+\gamma_{i}})|=|c^{*}-h^{*}(x_{t+1:t+\gamma_{i}})|\leq|-\inf-s^{*}(x_{t+1:t+\gamma_{i}})|$
-where $h^{*}(x_{t+1:t+\gamma_{i}})=\max_{x_{t+\gamma_{i}+1:N}}\log p(x_{t+\gamma_{i}+1:N}|x_{1:t+\gamma_{i}},prefix)$. Therefore,
-$|s_{MJGD}(x_{t+1:t+\gamma_{i}})-s^{*}(x_{t+1:t+\gamma_{i}})|\leq|s_{GD}(x_{1:K})-s^{*}(x_{t+1:t+\gamma_{i}})|$ (4)
-The derivation above provides a heuristic explanation to show that MJGD is a closer approximation to the optimal greedy decoding than vanilla greedy decoding.
-Hence, its output perplexity is generally better.

### v3

@@
 Abstract
-Large language models (LLMs) have achieved remarkable success across diverse tasks, yet their inference processes are hindered by substantial time and energy demands due to single-token generation at each decoding step.
-While previous methods such as speculative decoding mitigate these inefficiencies by producing multiple tokens per step, each token is still generated by its single-token distribution, thereby enhancing speed without improving output quality.
-In contrast, our work simultaneously enhances inference speed and improves the output effectiveness.
-We consider multi-token joint decoding (MTJD), which generates multiple tokens from their joint distribution at each iteration, theoretically reducing perplexity and enhancing task performance.
+Large language models (LLMs) have achieved remarkable success across diverse tasks, but, due to single-token generation at each decoding step, their inference processes are hindered by substantial time and energy demands.
+While previous methods such as speculative decoding mitigate these inefficiencies by producing multiple tokens per step, each token is still generated by its single-token distribution.
+Although this enhances the speed, it does not improve the output quality.
+In contrast, our work simultaneously boosts inference speed and improves the output effectiveness.
+We consider multi-token joint decoding (MTJD), which generates multiple tokens from their joint distribution at each iteration, theoretically reducing perplexity and raising task performance.
 However, MTJD suffers from the high cost of sampling from the joint distribution of multiple tokens.
 Inspired by speculative decoding, we introduce multi-token assisted decoding (MTAD), a novel framework designed to accelerate MTJD.
-MTAD leverages a smaller auxiliary model to approximate the joint distribution of a larger model, incorporating a verification mechanism that not only ensures the accuracy of this approximation, but also improves the decoding efficiency over conventional speculative decoding.
-Theoretically, we demonstrate that MTAD closely approximates exact MTJD with bounded error.
-Empirical evaluations using Llama-2 and OPT models ranging from 13B to 70B parameters across various tasks reveal that MTAD reduces perplexity by 21.2% and improves downstream performance compared to standard single-token sampling.
-Furthermore, MTAD achieves a 1.42$\times$ speed-up and consumes 1.54$\times$ less energy than conventional speculative decoding methods.
-These results highlight MTAD’s ability to make multi-token joint decoding both effective and efficient, promoting more sustainable and high-performance deployment of LLMs.
+MTAD leverages a smaller auxiliary model to approximate the joint distribution of a larger model, incorporating a verification mechanism that not only ensures the accuracy of this approximation, but also increases the decoding efficiency over conventional speculative decoding.
+Theoretically, we demonstrate that MTAD closely approximates exact MTJD with a bounded error.
+Empirical evaluations across various tasks reveal that MTAD improves downstream performance by 25% compared to standard single-token sampling.
+Furthermore, MTAD achieves a 1.42$\times$ speed-up and consumes 1.54$\times$ less energy than vanilla speculative decoding methods.
+These results highlight MTAD’s ability to make multi-token joint decoding both effective and efficient, promoting more productive and high-performance deployment of LLMs.
+[^1] We release our code at https://github.com/ZongyueQin/MTAD
 1 Introduction
@@
 But these methods usually require task-dependent techniques and information to match the performance of autoregressive decoding (Kim et al., 2023; Xiao et al., 2023).
-More recently, speculative decoding has emerged (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023), exploiting the observation that most of the small model’s prediction aligns well with that of a large model.
+More recently, speculative decoding has emerged (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023).
+This method exploits the observation that most of the small model’s prediction aligns well with that of a large model.
 It leverages a smaller auxiliary model to draft a few future tokens autoregressively, which are subsequently validated in parallel by the larger model.
-As the smaller model operates significantly faster and parallel token verification incurs a similar time cost as generating a single token, speculative decoding achieves an overall speed-up of $1$-2×2$\times$2 ×.
+As the smaller model operates significantly faster and parallel token verification incurs a similar time cost as generating a single token, speculative decoding attains an overall speed-up of $1$-2×2$\times$2 ×.
 Despite gains in speed, these methods still generate each token based on its single-token probability.
 Consequently, it does not enhance the effectiveness of the generated sequences.
-In this work, we aim to go beyond the conventional trade-off between efficiency and effectiveness by introducing multi-token joint decoding (MTJD).
+In this work, we first go beyond the conventional trade-off between efficiency and effectiveness and explore multi-token joint decoding (MTJD).
 Unlike traditional approaches, MTJD produces multiple tokens from their joint distribution at each decoding step.
@@
 To ensure the accuracy of this approximation, MTAD incorporates a verification mechanism that not only guarantees the accuracy of the draft tokens but also enhances efficiency beyond conventional speculative decoding by maximizing the number of accepted tokens per iteration.
-We provide both theoretical and empirical analyses to demonstrate that MTAD improves perplexity and downstream performance.
-Meanwhile, it achieves significant reductions in energy and time usage compared to existing decoding strategies.
+We provide both theoretical and empirical analyses to demonstrate that MTAD boosts perplexity and downstream performance.
+Meanwhile, it significantly reduces the energy and time usage compared to existing decoding strategies.
 Our contributions are as follows:
@@
     We analyze the energy consumption of LLM inference.
-To our knowledge, we are the first to give theoretical and empirical evidence that, despite that MTAD and other speculative decoding algorithms increase the number of FLOPs needed during LLM inference, they reduce the overall energy consumption by reducing the overhead induced by accessing GPU global memory.
+To our knowledge, we are the first to give quantified and empirical evidence that, despite the fact that MTAD and other speculative decoding algorithms increase the number of FLOPs needed during LLM inference, they actually use less energy with fewer accesses to the GPU global memory.
   4. 4.
-    We conducted comprehensive evaluations with Llama-2 and OPT models (ranging from 13B to 70B parameters) across various tasks, demonstrating that MTAD enhances perplexity by 21.2% and improves downstream effectiveness compared to standard single-token sampling, while also achieving a 1.42$\times$ speed-up and reducing energy consumption by 1.54$\times$ compared to conventional speculative decoding methods.
+    We conducted comprehensive evaluations across various tasks, demonstrating that MTAD improves downstream performance by 25% compared to standard single-token sampling, while, at the same time, realizing a 1.42$\times$ speed-up and paring energy consumption by 1.54$\times$ compared to vanilla speculative decoding methods.
 These advancements position MTAD as a robust solution for making multi-token joint decoding both effective and efficient, thereby facilitating more sustainable and high-performance deployment of large-scale language models.
-Our code is publicly available [^1] https://anonymous.4open.science/r/LLMSpeculativeSampling-EE52.
 2 Preliminaries
+In this section, we discuss preliminaries relevant to contextualizing our paper.
 2.1 Decodings of LLMs
@@
 Let $p$ denote the distribution defined by LLM model $M_{p}$.
-Given an input context $input$, a decoding algorithm generates a sequence of $N$ tokens whose likelihood is denoted as $p(x_{1:N}|input)$.
-The likelihood of the sequence is directly linked to perplexity of the sequence, which is the exponentiated average negative log-likelihood of all tokens.
+Given an input context $input$, a decoding algorithm generates a sequence of $N$ tokens whose likelihood is designated as $p(x_{1:N}|input)$.
+The likelihood of the sequence is directly linked to the perplexity of the sequence, which is the exponentiated average negative log-likelihood of all tokens.
 Based on autoregressive decomposition $p(x_{1:N}|input)=\prod_{t=1}^{N}p(x_{t}|x_{1:t-1},input)$ [^2] In the paper, we omit $input$ when there is no ambiguity., the perplexity is defined as:
@@
 In practice, when a model is well-trained, lower perplexity often correlates with improved downstream performance.
-For example, beam sampling aims to return output with lower perplexity and is proven to have better downstream performance in general (Shi et al., 2024).
+For example, beam sampling (explained below) aims to return output with lower perplexity and is empirically proven to have better downstream performance in general (Shi et al., 2024).
 To further demonstrate the relationship between perplexity and downstream performance, we evaluate GPT-3.5-turbo on the spider (Yu et al., 2018) dataset.
-Using a temperature of 2, the model generated 10 outputs for each input.
-We measured the average perplexities and execution accuracies for the outputs with the highest, lowest, and median (the 5-th lowest) perplexity.
-As shown in Table 1, lower perplexity correlates with improved downstream performance, even in one of today’s largest models.
+Employing a temperature of 2, the model generated 10 outputs for each input.
+We measured the average perplexities and execution accuracies for the outputs with the highest, lowest, and median (the 5th lowest) perplexity.
+As seen in Table 1, lower perplexity correlates with improved downstream performance, even in one of today’s largest models.
 Table 1: Relationship between perplexity and execution accuracy (EA, higher the better) for GPT-3.5-turbo.
@@
 Multinomial sampling, also known as standarized sampling or single-token sampling, samples the next token $x_{t}$ based on $\mathcal{T}\circ p(\cdot|x_{1:t-1},input)$, where $\mathcal{T}$ is a warping operation applied to enhance the high probability region.
-Some common warping operations include top-k warping, which limits the selection to the top k tokens, and top-p warping, where tokens are sampled from the smallest possible subset of the vocabulary whose cumulative probability mass exceeds a specified threshold.
-The deterministic version of multinomial sampling (i.e., greedy decoding) is a special case when $k=1$.
+Some common warping operations include top-k warping.
+This limits the selection to the top k tokens, and top-p warping, where tokens are sampled from the smallest possible subset of the vocabulary whose cumulative probability mass exceeds a specified threshold.
+The deterministic version of multinomial sampling is a special case with $k=1$, also called greedy decoding.
 Beam Sampling.
-Beam sampling aims to improve output perplexity over multinomial sampling.
+Beam sampling is intended to decrease output perplexity over multinomial sampling.
 For each position $t$ ($1\leq t\leq N$), it maintains $W>1$ candidate sequences, which are also called beams.
-Assume we have already kept the $W$ sequences $\mathcal{I}_{t-1}=\{x_{1:t-1}^{(1)},\ldots,x_{1:t-1}^{(W)}\}$ at position $t-1$, $W$ sequences with length $t$ are then sampled from $\mathcal{T}\circ p_{beam}$, where $p_{beam}:\mathcal{I}_{t-1}\times V\rightarrow[0,1]$ is the beam sampling probability:
-$p_{beam}(x_{1:t-1}^{(i)},x_{t})=\frac{p(x_{1:t-1}^{(i)},x_{t}|input)}{\sum_{x_{1:t-1}^{(j)},x^{\prime}_{t}\in\mathcal{I}_{t-1}\times V}p(x_{1:t-1}^{(j)},x_{t}^{\prime}|input)}$ (2)
+Assume we have already kept the $W$ sequences $\mathcal{I}_{t-1}=\{x_{1:t-1}^{(1)},\ldots,x_{1:t-1}^{(W)}\}$ at position $t-1$.
+$W$ sequences with length $t$ are then sampled from $\mathcal{T}\circ p_{beam}$, where $p_{beam}:\mathcal{I}_{t-1}\times V\rightarrow[0,1]$ is the beam sampling probability:
+$p_{beam}(x_{1:t-1}^{(i)},x_{t})=\frac{p(x_{1:t-1}^{(i)},x_{t}|input)}{\sum_{1\leq j\leq W,x^{\prime}_{t}\in V}p(x_{1:t-1}^{(j)},x_{t}^{\prime}|input)}$ (2)
 Notice that $p(x_{1:t-1}^{(i)},x_{t}|input)=p(x_{t}|x_{1:t-1}^{(i)},input)\cdot p(x_{1:t-1}^{(i)}|input)$.
@@
   1. 1.
-    Let $input$ be the input context, the small model samples $\gamma$ draft tokens $x_{1},\ldots,x_{\gamma}$ using multinomial sampling based on $\tilde{q}(x_{t}|x_{1:t-1},input))$ for $t=1,\ldots,\gamma$, where $\tilde{q}=\mathcal{T}\circ q$ and $q$ is the small model’s output distribution.
+    Let $input$ be the input context, the small model samples $\gamma$ draft tokens $x_{1},\ldots,x_{\gamma}$ with multinomial sampling based on $\tilde{q}(x_{t}|x_{1:t-1},input))$ for $t=1,\ldots,\gamma$, where $\tilde{q}=\mathcal{T}\circ q$ and $q$ is the small model’s output distribution.
   2. 2.
@@
   4. 4.
-    Repeat step 1-3 until reaching the stopping criteria, e.g., reaching the length limit..
+    Repeat step 1-3 until reaching the stopping criteria, e.g., reaching the length limit.
 Because the large model verifies $\gamma$ tokens in parallel with one run, the time cost is smaller than calling it $\gamma$ times.
-Meanwhile, although the small model still runs in an autoregressive way, its inference speed is much faster than the large model.
-As a result, speculative decoding achieves a speedup of 1–2× compared to multinomial sampling while maintaining an identical sampling distribution.
+Moreover, the global memory access is also pared, which saves energy consumption, as we shall illustrate in Section 4.
+Meanwhile, although the small model still runs in an autoregressive way, its inference speed is more efficient than the large model.
+As a result, speculative decoding maintains an identical sampling distribution while realizing a speedup of 1–2$\times$ compared to multinomial sampling and using less energy.
 3 Methodology
 As discussed in Section 2, the goal of this work is to design an algorithm that yields lower perplexity and better efficiency than multinomial sampling and vanilla speculative decoding.
-In this section, we first introduce multi-token joint decoding (MTJD), which generates multiple tokens based on their joint likelihood.
+In this section, we first introduce multi-token joint decoding (MTJD).
+This generates multiple tokens based on their joint likelihood.
 We prove it can yield lower perplexity.
-Then we introduce multi-token assisted decoding (MTAD), which approximates and accelerates MTJD by exploiting an auxiliary model.
+Then we present multi-token assisted decoding (MTAD), which approximates and accelerates MTJD by exploiting an auxiliary model.
 3.1 Multi-Token Joint Decoding
-We first introduce a new decoding algorithm to improve multinomial sampling in terms of perplexity.
+We first explain a new decoding algorithm to improve multinomial sampling in terms of perplexity.
 Definition 3.1.
@@
 Let $M_{p}$ be the large target model with distribution $p$.
-Different from single-token multinomial sampling, multi-token joint decoding (MTJD) generates the next $\gamma_{i}$ tokens at step $i$ based on their joint conditional probability $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$, where $\gamma_{i}$ is an integer no less than 1 and $t=\sum_{i^{\prime}=1}^{i-1}{\gamma_{i^{\prime}}}$, i.e., the total tokens generated in the previous $i-1$ steps.
+Different from single-token multinomial sampling, multi-token joint decoding (MTJD) produces the next $\gamma_{i}$ tokens at step $i$ based on their joint conditional probability $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$, where $\gamma_{i}$ is an integer no less than 1 and $t=\sum_{i^{\prime}=1}^{i-1}{\gamma_{i^{\prime}}}$, i.e., the total tokens generated in the previous $i-1$ steps.
 Figure 1: Perplexity and Rouge-L score of the output when $\gamma_{i}=K$ for MTJD with OPT-125M and Llama-2-68M fine-tuned on ChatGPT-Prompts (Rashad, 2023) dataset.
@@
 So MTJD is less prone to choosing local optima.
-Theorem 3.2 shows the limit of perplexity of MTJD when $N$ approaches infinity. The proofs are included in the Appendix A.
+Theorem 3.2 demonstrates the limit of perplexity of MTJD when $N$ approaches infinity. The proofs are included in the Appendix A.
 Theorem 3.2.
@@
 We fine-tune both a Llama and an OPT model on the ChatGPT-Prompts dataset and evaluate the output perplexity and Rouge-L scores with example outputs.
-Figure 1 shows the output perplexity and Rouge-L scores of MTJD with $\gamma_{i}$ set to a constant $K$, where $K=1,\ldots,5$.
+Figure 1 depicts the output perplexity and Rouge-L scores of MTJD with $\gamma_{i}$ set to a constant $K$, where $K=1,\ldots,5$.
 Notice that setting $K=1$ is equivalent to multinomial sampling.
@@
 Unfortunately, the computation cost of MTJD is infeasible in practice, since the time and space complexity to compute the joint distribution of $\gamma_{i}$ tokens is $|V|^{\gamma_{i}}$.
-Inspired by speculative decoding and the facts that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s prediction deviate from those of the large model” (Leviathan et al., 2023; Kim et al., 2023), we propose multi-token assisted decoding (MTAD), which exploits a small auxiliary model $M_{q}$ to accelerate MTJD approximately.
+Inspired by speculative decoding and the fact that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s prediction deviate from those of the large model” (Leviathan et al., 2023; Kim et al., 2023), we propose multi-token assisted decoding (MTAD), which exploits a small auxiliary model $M_{q}$ to accelerate MTJD approximately.
 The core idea is to (1) use the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$ output by $M_{q}$ to approximate $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^3] It is also valid to approximate $\tilde{p}$ with $\tilde{q}$.
 Without loss of generality, we consider non-warped distribution in the illustration of MTAD.
-and generate $\gamma$ draft tokens from $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$, then (2) use the large model to validate draft tokens in parallel and accept the longest draft prefix sub-sequence that passes verification, and (3) sample an additional token from the distribution of the large model without extra overhead to ensure at least one token is generated at each iteration.
+and produce $\gamma$ draft tokens from $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$, then (2) utilize the large model to validate draft tokens in parallel and accept the longest draft prefix sub-sequence that passes verification, and (3) sample an additional token from the distribution of the large model without extra overhead to ensure at least one token is generated at each iteration.
 However, it is still infeasible to directly generate draft tokens from the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$.
 So we propose to further approximate this process with beam sampling, which is an effective and efficient algorithm to generate sequences with high likelihood.
-In this way, MTAD reduces the number of runs of the large model to generate $N$ tokens, thus accelerating the inference in the same way as vanilla speculative decoding does.
+In this way, MTAD decreases the number of runs of the large model to generate $N$ tokens, thus accelerating the inference in the same way as vanilla speculative decoding does.
 Algorithm 1 in the Appendix illustrates the pseudocode of MTAD algorithm.
@@
 Draft Tokens Verification
-Figure 2 illustrates the verification process of MTAD.
+Figure 2 displays the verification process of MTAD.
 Let $x_{t+1},\ldots,x_{t+\gamma}$ be the draft tokens generated by beam sampling with the auxiliary model.
-Since beam sampling is a widely recognized algorithm to generate sequences with high overall likelihood (Leblond et al., 2021), it is reasonable to assume $q(x_{t+1:t+\gamma}|x_{1:t})$ is large.
-Also, since beam sampling works in an autoregressive way, we can also assume that $\forall j\in\{1,\ldots,\gamma\}$, $q(x_{t+1:t+j}|x_{1:t})$ is large.
+Since beam sampling is a widely recognized algorithm to produce sequences with high overall likelihood (Leblond et al., 2021), it is reasonable to assume $q(x_{t+1:t+\gamma}|x_{1:t})$ is large.
+Also, since beam sampling works in an autoregressive way, we can also infer that $\forall j\in\{1,\ldots,\gamma\}$, $q(x_{t+1:t+j}|x_{1:t})$ is large.
 To approximate MTJD, for each step $i$, MTAD needs to ensure the accepted tokens $x_{t+1:t+\gamma_{i}}$ ($0\leq\gamma_{i}\leq\gamma$) also have high joint likelihood with the large model $M_{p}$.
@@
 The verification step of MTAD ensures that the accepted tokens have a high joint likelihood with the large model.
-We have shown that selecting multiple tokens based on their joint likelihood lead to better output perplexity.
+We have shown that selecting multiple tokens based on their joint likelihood leads to better output perplexity.
 Thus, MTAD is more effective than multinomial sampling and vanilla speculative decoding.
 Furthermore, since MTAD accepts the longest draft sub-sequence with high likelihood, it can tolerate low-quality tokens as long as the joint likelihood is high.
-So at each iteration, MTAD can accept more draft tokens than vanilla speculative decoding, which results in better efficiency.
+So at each iteration, MTAD can admit more draft tokens than vanilla speculative decoding, which results in better efficiency.
 Next, we theoretically analyze the approximation error of MTAD.
 Lemma 3.4 shows the upper bound of MTAD’s perplexity.
-And Theorem 3.5 shows the upper bound of the ratio between the perplexity of approximate MTAD and exact MTJD.
+Theorem 3.5 reveals the upper bound of the ratio between the perplexity of approximate MTAD and exact MTJD.
 The proofs are given in Appendix A.
@@
 Let us assume that when the small auxiliary model generates draft tokens with beam sampling, the beam width is large enough such that the returned log-likelihood is close to the maximum log-likelihood, i.e.,
-$\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))\geq(1-\epsilon)\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\max_{x_{\Gamma_{i-1}+1:\Gamma_{i}-1}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))$ (5)
+$\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}})\geq$ (5)
+$(1-\epsilon)\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\left(\max_{x_{\Gamma_{i-1}+1:\Gamma_{i}-1}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}})\right)$
 where $\epsilon$ is an error term and $\epsilon\leq 0$ because $\log q\leq 0$.
 Furthermore, let $H(p,q)$ the single-token cross entropy between $p$ and $q$, i.e., $H(p,q)=-\mathbb{E}_{x_{1:t\in\mathcal{X}}}\sum_{x_{t+1}}p(x_{t+1}|x_{1:t})\log q(x_{t+1}|x_{1:t})$.
-With the two assumption above, when $N\rightarrow\infty$ we have
+With the two assumptions above, when $N\rightarrow\infty$ we have
 $PPL_{q}({x_{1:\Gamma_{N}}})\leq\exp(-\frac{1-\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{q}(\gamma-1,\arg\max\circ q)+\frac{H(p,q)}{\bar{\gamma}})$ (6)
@@
 With Theorem 3.6, we observe that when $q\rightarrow p$, we have $U\rightarrow 0$ and $\bar{\gamma}\rightarrow\infty$.
-Meanwhile, when $\epsilon\rightarrow 0$, meaning the beam width for the auxiliary model is large enough, the ratio bound in Theorem 3.5 converges to 1, It implies that MTAD converges to MTJD under these limiting conditions.
-Similar to Spectr (Sun et al., 2023) and SpecInfer (Miao et al., 2023), it is possible to enhance the number of accepted tokens in MTAD by allowing the draft model to generate multiple draft sequences and applying tree-based attention (Miao et al., 2023) for simultaneous verification.
+Meanwhile, when the beam width for the auxiliary model is large enough, $\epsilon\rightarrow 0$, and the ratio bound in Theorem 3.5 converges to 1, This implies that MTAD converges to MTJD under these limiting conditions.
+Similar to Spectr (Sun et al., 2023) and SpecInfer (Miao et al., 2023), it is possible to enhance the number of accepted tokens in MTAD by allowing the draft model to produce multiple draft sequences and by applying tree-based attention (Miao et al., 2023) for simultaneous verification.
 However, our preliminary experiment results suggest that since MTAD already selects the longest accepted prefix sub-sequence, the advantage of generating multiple draft tokens is less significant.
@@
 To our knowledge, there is no existing work evaluating the impact of speculative decoding on inference energy consumption.
-Although MTAD and speculative decoding increase the number of FLOPs due to the involvement of a small auxiliary model and the rollback operation, they concurrently reduce the inference time and memory operations, which are key factors of GPU energy consumption (Allen & Ge, 2016; Chen et al., 2011).
+Although MTAD and speculative decoding raise the number of FLOPs due to the involvement of a small auxiliary model and the rollback operation, they concurrently reduce the inference time and memory operations, which are key factors of GPU (or TPU) energy consumption (Allen & Ge, 2016; Chen et al., 2011).
 Consequently, it poses an open question regarding whether speculative decoding increases or decreases overall energy consumption.
@@
 $E_{total}=PW_{flop}T_{flop}+PW_{mem}T_{mem}$ (9)
-where $PW_{flop},PW_{mem}$ denote the power (energy/second) of FLOPs and memory operations, and $T_{flop},T_{mem}$ denote the time spent on these operations.
-When input batch size increases, $PW_{flop}$ increases until it reaches the power of maximum FLOPs, denoted as $PW^{*}_{flop}$.
-Meanwhile, $PW_{mem}$ is irrelevant to the input batch size, as it only depends on the memory hardware.
+where $PW_{flop},PW_{mem}$ denote the power (energy/second) of a unit FLOP and memory operation, respectively, and $T_{flop},T_{mem}$ are the total time spent on these operations.
+When input batch size increases, $PW_{flop}$ rises until it reaches the power of maximum FLOPs, designated as $PW^{*}_{flop}$.
+$PW_{mem}$ is irrelevant to the input batch size because it only depends on the memory hardware.
 Table 2: The effect of batch size to inference speed and energy consumption. The number of inputs is the product of the number of LLM runs and input batch size.
@@
 J/token 11.3 5.74 11.2 6.97
-To determine the relative magnitude relationship between $PW_{flop}$ and $PW_{mem}$, we first point out the fact that GPU memory operations in LLM inference are dominated by accessing off-chip global memory, which consumes about 100×100$\times$100 × of energy compared to accessing on-chip shared memory (Jouppi et al., 2021).
-It is because each multiprocessor on GPU usually has 64KB of on-chip memory shared by multiple threads, while storing a single layer of LLM, say T5-11b (Raffel et al., 2020), requires about 1GB memory.
-Moreover, Allen and Ge showed that doing sequential read from off-chip memory consumes $20$-$30\%$ more power than running maximum FLOPs (Allen & Ge, 2016).
+To determine the relative magnitude relationship between $PW_{flop}$ and $PW_{mem}$, we first point out the fact that GPU memory operations in LLM inference are dominated by accessing off-chip global memory.
+This consumes about 100×100$\times$100 × of energy compared to accessing on-chip shared memory (Jouppi et al., 2021).
+Because each multiprocessor on a GPU usually has 64KB of on-chip memory shared by multiple threads, but to store a single layer of LLM, say T5-11b (Raffel et al., 2020), requires about 1GB of memory.
+Moreover, Allen and Ge showed that doing a sequential read from off-chip memory consumes $20$-$30\%$ more power than running maximum FLOPs (Allen & Ge, 2016).
 So we have $PW_{mem}>PW^{*}_{flop}\geq PW_{flop}$.
 Notice that $PW^{*}_{flop}=PW_{flop}$ only if the batch size reaches the maximum parallelization capacity of GPUs.
-During multinomial sampling and speculative decoding, the batch size is usually small (Leviathan et al., 2023).
+During multinomial sampling and speculative decoding, the batch size is generally small (Leviathan et al., 2023).
 So most of the computing power is not utilized (Leviathan et al., 2023), which means $PW_{mem}\gg PW_{flop}$.
-In addition, previous studies have shown that during LLM inference $T_{mem}\gg T_{flop}$ (Leviathan et al., 2023).
+In addition, previous studies have revealed that during LLM inference $T_{mem}\gg T_{flop}$ (Leviathan et al., 2023).
 Therefore, the energy induced by memory operations, i.e., $PW_{mem}T_{mem}$ dominates $E_{total}$.
-Since speculative decoding reduces $T_{mem}$ by reducing the number of runs of the large model, it should reduce the inference energy consumption to a similar extent as it reduces time consumption.
+Since speculative decoding lowers $T_{mem}$ by reducing the number of runs of the large model, it should cut the inference energy consumption to a similar extent as it reduces time consumption.
 To validate our hypothesis, we conducted an experiment to evaluate how batch size influences energy consumption during inference.
@@
 Table 2 shows the results.
-As batch size doubles, although the number of FLOPs doubles, the energy consumption per run increases slightly.
+As batch size doubles, although the number of FLOPs doubles, the energy consumption per run goes up slightly.
 This observation demonstrates that $PW_{mem}T_{mem}$ dominates $E_{total}$.
-Moreover, we measured the speed and energy consumption of running multinomial sampling with the large model and speculative decoding using OPT (125M, 13B) and Llama-2 (68M, 13B) models.
-The results, shown in Table 3, indicate that speculative decoding reduces the energy consumption and the time cost.
+Moreover, we measured the speed and energy consumption of running multinomial sampling with the large model and speculative decoding with OPT (125M, 13B) and Llama-2 (68M, 13B) models.
+The results, seen in Table 3, indicate that speculative decoding lowers the energy consumption and the time cost.
 This observation corroborates our claim to the energy efficiency of speculative decoding.
@@
 Datasets and Models.
-We use five public datasets for evaluation: (1) ChatGPT-Prompt (Rashad, 2023), (2) ChatAlpaca (Bian et al., 2023), (3) CNN Dailymail (See et al., 2017), (4) Spider (Yu et al., 2018), and (5) MT-bench (Zheng et al., 2023).
-Table 8 in the Appendix shows more details of the datasets.
-Following previous studies (Kim et al., 2023), we use two public LLM families in our experiments: OPT (Zhang et al., 2022) and Llama-2 (Touvron et al., 2023).
-In this section, we set the large model to be OPT-13B and Llama-2-13B as they are the largest models that can run on a single 40GB GPU, and use Llama-68M (Miao et al., 2023) and OPT-125M as the small models.
-Appendix C reports additional experiment results with OPT-30B and Llama-2-chat-70B.
+In the main paper, we report results with three public datasets for evaluation: (1) Spider (Yu et al., 2018), MTBench (Zheng et al., 2023), and HumanEval (Chen et al., 2021).
+We use Llama-3-8B and Llama-3-8B-Instruct (Dubey et al., 2024) as target models, and Llama-3-1B and Llama-3-1B-Instruct as their draft models, respectively.
+We provide additional experiments with other datasets and model families in Appendix C.
 Baselines.
-For each pair of small and large models, we compare our method with four speculative decoding methods: vanilla speculative decoding (speculative(Lee et al., 2018; Chen et al., 2023), Spectr (Sun et al., 2023), SpecInfer (Miao et al., 2023), and BiLD (Kim et al., 2023).
-Our implementation of MTAD and all the baselines are based on a public implementation of speculative decoding (Bear, 2024).
+We compare our method with six speculative decoding methods, including four lossless decoding methods: vanilla speculative decoding (SpD(Lee et al., 2018; Chen et al., 2023), Spectr (Sun et al., 2023), SpecInfer (Miao et al., 2023), MCSS (Yang et al., 2024), and two lossy speculative decoding methods: BiLD (Kim et al., 2023) and typical decoding (Cai et al., 2024).
+All the baselines and our method utilize the same pair of draft and target models without any fine-tuning.
 For each method, we let it generate at most 128 tokens for each input and run it for $1,000$ seconds.
-We open-sourced our code for reproduction.
-All the methods are stochastic with top-$k$ and top-$p$ sampling.
-The details of the hyper-parameters (e.g., $k$ and $p$) and machine configurations of the experiments can be found in the Appendix D, E, and F.
+All the methods are stochastic with top-$k$ and top-$p$ sampling with the temperature = 1.
+The details of the hyper-parameters (e.g., $k$ and $p$) and machine configurations of the experiments can listed in the Appendix D, E, and F.
 Appendix C reports additional experiments and ablation studies.
-Table 4: Inference efficiency and output perplexity of different methods on ChatGPT-Prompt (CP), ChatAlpaca (CA), CNNDailyMail (CD), Spider (SP), and MT-Bench (MT) datasets.
-Bold numbers mark the best result, underlined numbers mark the second best.
-[…]
-[…]
-energy (J/token) $\downarrow$ 6.62$\pm$0.91 7.45$\pm$0.90 5.17$\pm$0.88 9.52$\pm$0.10 3.38$\pm$0.02
-perplexity $\downarrow$ 3.64$\pm$0.11 3.15$\pm$0.06 3.64$\pm$0.08 3.64$\pm$0.11 2.06$\pm$0.06
-OPT speed (token/s) $\uparrow$ 33.8$\pm$2.47 31.5$\pm$1.87 38.0$\pm$2.20 32.8$\pm$ 0.58 55.8$\pm$0.30
-energy (J/token) $\downarrow$ 7.48$\pm$0.07 8.75$\pm$0.13 6.08$\pm$0.11 10.3$\pm$1.49 3.61$\pm$0.03
-perplexity $\downarrow$ 5.47$\pm$0.11 4.51$\pm$0.09 5.27$\pm$0.09 5.12$\pm$0.01 3.00$\pm$0.09
-CA Llama-2 speed (token/s) $\uparrow$ 31.6$\pm$0.35 28.8$\pm$0.20 27.7$\pm$0.29 26.5$\pm$0.49 44.1$\pm$0.25
-energy (J/token) $\downarrow$ 6.98$\pm$0.15 7.99$\pm$0.15 7.20$\pm$0.08 7.52$\pm$0.32 4.72$\pm$0.03
-perplexity $\downarrow$ 2.13$\pm$0.03 1.95$\pm$0.03 2.15$\pm$0.01 2.15$\pm$0.01 1.88$\pm$0.05
-OPT speed (token/s) $\uparrow$ 35.6$\pm$0.45 38.5$\pm$0.93 28.4$\pm$0.34 31.4$\pm$0.39 49.6$\pm$0.42
-energy (J/token) $\downarrow$ 5.74$\pm$0.11 5.12$\pm$0.06 6.24$\pm$0.11 8.68$\pm$1.83 4.03$\pm$0.02
-perplexity $\downarrow$ 3.32$\pm$0.10 2.60$\pm$0.06 3.16$\pm$0.06 3.42$\pm$0.03 2.07$\pm$0.03
-CD Llama-2 speed (token/s) $\uparrow$ 30.7$\pm$0.18 30.5$\pm$0.21 25.0$\pm$0.31 24.6$\pm$0.06 44.2$\pm$0.99
-energy (J/token) $\downarrow$ 7.07$\pm$0.19 7.41$\pm$0.16 8.22$\pm$0.19 7.59$\pm$0.85 4.80$\pm$0.12
-perplexity $\downarrow$ 2.87$\pm$0.08 2.93$\pm$0.03 3.06$\pm$0.11 2.92$\pm$0.09 2.63$\pm$0.10
-OPT speed (token/s) $\uparrow$ 31.7$\pm$0.91 30.9$\pm$0.80 23.7$\pm$0.40 25.7$\pm$0.36 43.6$\pm$0.33
-energy (J/token) $\downarrow$ 6.37$\pm$0.11 6.71$\pm$0.17 7.31$\pm$0.17 8.03$\pm$0.63 4.86$\pm$0.03
-perplexity $\downarrow$ 3.97$\pm$0.06 3.74$\pm$0.09 4.04$\pm$0.07 3.92$\pm$ 0.34 3.17$\pm$0.06
-SP Llama-2 speed (token/s) $\uparrow$ 24.0$\pm$0.28 26.2$\pm$0.08 24.2$\pm$0.29 23.8$\pm$0.20 26.4$\pm$0.28
-energy (J/token) $\downarrow$ 10.75$\pm$0.02 9.84$\pm$0.07 11.0$\pm$0.08 11.0$\pm$0.76 9.01$\pm$0.07
-perplexity $\downarrow$ 2.26$\pm$0.01 2.13$\pm$0.03 2.29$\pm$0.04 2.29$\pm$0.03 1.87$\pm$0.03
-OPT speed (token/s) $\uparrow$ 24.6$\pm$0.30 29.9$\pm$0.55 19.8$\pm$0.13 24.1$\pm$0.10 34.4$\pm$0.46
-energy (J/token) $\downarrow$ 15.6$\pm$3.55 13.6$\pm$3.07 20.1$\pm$2.52 16.9$\pm$2.75 11.7$\pm$2.36
-perplexity $\downarrow$ 2.30$\pm$0.07 1.90$\pm$0.01 2.20$\pm$0.09 2.21$\pm$ 0.01 1.63$\pm$0.03
-MT Llama-2 speed (token/s) $\uparrow$ 23.0$\pm$1.10 23.7$\pm$1.43 19.1$\pm$2.71 23.7$\pm$2.03 29.4$\pm$2.71
-energy (J/token) $\downarrow$ 7.99$\pm$0.26 7.40$\pm$0.19 9.27$\pm$0.54 9.20$\pm$0.73 6.71$\pm$1.19
-perplexity $\downarrow$ 3.64$\pm$0.51 3.44$\pm$0.76 3.64$\pm$0.51 3.63$\pm$0.50 2.21$\pm$0.18
-OPT speed (token/s) $\uparrow$ 34.0$\pm$3.00 44.7$\pm$2.92 28.7$\pm$2.46 28.5$\pm$2.74 48.0$\pm$1.80
-energy (J/token) $\downarrow$ 12.1$\pm$0.36 6.23$\pm$0.67 12.9$\pm$1.73 13.2$\pm$1.88 6.11$\pm$0.82
-perplexity $\downarrow$ 2.02$\pm$0.40 1.50$\pm$0.27 1.97$\pm$0.38 1.99$\pm$ 0.33 1.10$\pm$0.03
-[…]
-5.1 Comparison with Baselines
-Table 4 shows the primary results of our experiments while Table 5 details the block efficiency, defined as the average number of tokens generated per iteration, of different methods.
-The standard deviations in the tables are computed by repeating each experiment four times [^4] For MT-Bench, the standard deviation also accounts for variations across different tasks..
-First, we observe that MTAD is significantly more efficient than all baselines in terms of both energy and time.
-On one hand, the energy consumption of MTAD is on average 1.54×1.54$\times$1.54 × smaller than that of vanilla speculative decoding.
-On the other hand, MTAD is 1.101.71×1.10-1.71$\times$1.10 - 1.71 × faster than vanilla speculative decoding, 1.38×1.38$\times$1.38 × faster than BiLD, 1.59×1.59$\times$1.59 × faster than Spectr, and 1.60×1.60$\times$1.60 × faster than SpecInfer.
-SpecInfer and Spectr has better block efficiency than vanilla speculative decoding, but are slower.
-This may be due to the fact that they have to verify multiple draft sequences, which introduces extra overhead and may not be perfectly parallelized, especially when the memory overhead exceeds the GPU memory capacity.
-Meanwhile, MTAD has the best block efficiency without causing any extra overhead, hence it is significantly more efficient.
-Table 5: Average number of tokens generated at each iteration across all datasets.
-Llama-2 OPT
-spec 2.02$\pm$0.05 2.60$\pm$0.06
-BiLD 1.83$\pm$0.10 2.68$\pm$0.36
-Spectr 2.73$\pm$0.43 3.45$\pm$0.42
-SpecInfer 2.74$\pm$0.46 3.45$\pm$0.40
-MTAD 3.17$\pm$0.43 4.30$\pm$0.03
-Next, we compare the output perplexity of different algorithms.
-The perplexity scores of vanilla speculative decoding, SpecInfer, and Spectr are close since their sampling distributions are equivalent.
-Meanwhile, BiLD approximates the sampling distribution of single-token multinomial sampling but yields better perplexity.
-It is because we set a strict acceptance threshold for BiLD, which lowers the acceptance rate but ensures every token has a high probability in the large model, thus improving the overall likelihood.
-More importantly, there is a significant gap between MTAD and other baselines.
-On average, the perplexity of MTAD is $21.2\%$ lower than that of speculative decoding.
-Table 6: Downstream task scores of speculative decoding and MTAD. All the scores are higher the better.
-spec MTAD
-CD Rouge-L 0.114 0.118
-SP EA 11.5 13.0
-MT Humanities 2.95 3.15
-Extraction 1.80 2.50
-Roleplay 3.10 3.80
-Math 1.10 1.00
-Coding 1.25 1.10
-Reasoning 3.80 3.15
-STEM 2.85 3.10
-Writing 3.80 3.65
-Average 2.58 2.68
-In addition, to show MTAD indeed improves the downstream effectiveness, we compare the performance metrics of speculative decoding and MTAD on CNNDM, Spider, and MT-Bench datasets.
-[^5] All speculative baselines are equivalent to multinomial sampling We exclude the other two datasets as they lack explicit downstream metrics.
-And we exclude the results of OPT models due to their consistently poor performance across all evaluated datasets.
-As illustrated in Table 6, MTAD outperforms speculative decoding across all three datasets, thereby validating our claim that MTAD achieves superior effectiveness compared to conventional decoding methods that rely on single-token distributions.
-5.2 Ablation Study
-5.2.1 Number of Beams
-First, we investigate how the number of beams used in the beam decoding of the small model affects the inference performance.
-Table 7 shows the results.
-Increasing the number of beams improves the quality of the draft tokens, which not only improves the output perplexity but also increases the average acceptance length and hence leads to better efficiency.
-But we can see that the increment slows down when the number of beams is large enough.
-In addition, when the number of beams is too large, the inference cost of the small model will become too high.
-Table 7: Effect of number of beams to the inference performance on ChatGPT-Prompts dataset.
-# beams 2 4 6 8
-speed (token/s) $\uparrow$ 55.9 59.9 60.2 61.3
-Llama-2 energy (J/token) $\downarrow$ 2.43 2.25 2.22 2.20
-perplexity $\downarrow$ 2.44 2.12 2.14 2.10
-speed (token/s) $\uparrow$ 51.0 54.1 54.3 55.9
-OPT energy (J/token) $\downarrow$ 2.50 2.32 2.36 2.30
-perplexity $\downarrow$ 3.63 3.16 3.42 3.19
-5.2.2 Acceptance Threshold
-Figure 3: Effect of acceptance threshold on output perplexity and decoding speed.
-Next, we evaluate the effect of acceptance threshold $\tau$.
-Intuitively, when we increase $\tau$ from 0 to 1, the acceptance criterion becomes more strict, the efficiency drops while the output perplexity increases.
-Surprisingly, this expectation is only partially correct.
-As shown in Figure 3, the efficiency indeed drops when $\tau$ increases.
-However, the perplexity increases when $\tau$ is close to 1.
-When $\tau=1$, all the draft tokens are rejected, which makes MTAD equivalent to multinomial sampling.
-Similarly, when $\tau$ is close to 1, the advantage of multi-token joint decoding on effectiveness disappears, hence the perplexity becomes close to the perplexity of multinomial sampling.
-Another surprising observation is that the perplexity of MTAD is good when $\tau=0$.
-When $\tau=0$, MTAD is equivalent to generating $\gamma$ tokens using beam decoding with the small model, then generating an additional token with the large model.
-The fact that MTAD achieves good perplexity when $\tau=0$ can be explained by the fact that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s predictions deviate from those of the large model” (Kim et al., 2023; Leviathan et al., 2023).
-Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MTAD is relatively stable, suggesting that MTAD is not sensitive to the acceptance threshold.
+5.1 Performance of Multi-Token Joint Decoding
+While most speculative decoding approaches focus on inference speed up, we want to design approaches that can also improve inference quality.
+We propose multi-token joint decoding (MTJD, Section 3.1) to accomplish the goal, due to its capability to achieve a lower perplexity and higher likelihood than single-token multinomial sampling.
+To validate that MTJD indeed betters output quality, we test MTJD (k=4) and standard multinomial sampling on Spider, MTBench, and HumanEval using the Llama-3 series models.
+We follow the same way introduced in Section 3.1 to implement MTJD.
+For this process, the higher the scores, the better the downstream performance.
+Under all settings, MTJD realizes the highest scores and lower perplexity.
+These results show a clear advantage for MTJD in terms of output quality.
+Table 4: Performance comparison of single-token sampling and multi-token joint sampling.
+We use Llama-3.1-8B and Llama-3.1-8B-Instruct as target models, and Llama-3.2-1B and Llama-3.2-1B-Instruct as the draft models.
+Llama-3 (8B,1B) Llama-3-Instruct (8B,1B)
+  Spider  MTBench HumanEval   Spider  MTBench HumanEval
+Single-token multinomial sampling
+Score 22.0 3.40 15.9 36.0 4.11 28.0
+PPL 2.58 2.40 2.09 2.23 1.91 1.85
+Multi-token joint sampling
+Score 52.5 3.77 36.6 60.5 4.40 49.4
+PPL 1.16 1.32 1.26 1.18 1.27 1.15
+5.2 Performance of Multi-Token Assisted Decoding
+Next, we evaluate the efficiency and effectiveness of MTAD, an approximate algorithm that accelerates MTJD while preserving its downstream performance advantages.
+Table 5 presents the decoding speed, energy consumption, and downstream performance of various decoding algorithms across different datasets.
+Efficiency Analysis.
+We first observe that MTAD is the most efficient among all baselines in terms of both energy and time.
+On average, MTAD is 17.2% faster than the most efficient lossless baseline, MCSS, while consuming 6.1% less energy.
+Compared to lossy decoding algorithms, it attains 27.2% (15.1%) higher speed and 17.9% (14.1%) lower energy consumption than BiLD (typical decoding), respectively.
+Interestingly, despite utilizing only a single draft sequence, MTAD outperforms baselines that employ multiple draft sequences, such as Spectr, MCSS, and SpecInfer.
+In these methods, verification terminates immediately upon rejecting a token.
+In contrast, MTAD continues verification even after a rejection, searching for future tokens that may still pass.
+This mechanism results in a greater acceptance length per iteration than the baselines.
+Downstream Performance Comparison.
+Next, we compare the downstream performance of different decoding algorithms.
+Notice that while lossless decoding algorithms theoretically sample from the target distribution, they exhibit slight variations in downstream performance.
+This discrepancy arises because, despite preserving the original distribution, differences in the sampling process prevent them from generating identical sequences even when the random seed is fixed.
+Furthermore, we observe that lossy decoding algorithms can reach higher downstream performance at the expense of efficiency.
+This suggests that all lossy decoding methods can trade off efficiency for performance by adjusting verification strictness.
+Most notably, MTAD consistently achieves the highest downstream performance.
+On average, it surpasses lossless decoding algorithms by 25.0%, BiLD by 11.8%, and typical decoding by 9.9%.
+These results confirm our claim that MTAD offers superior effectiveness compared to conventional decoding methods that rely solely on single-token distributions.
+Table 5: Comparison of different speculative decoding methods across various models and metrics. Bold indicates best values, underline indicates second-best.
+Lossy Decoding Lossless Decoding Ours
+BiLD Typical SpD Spectr SpecInfer MCSS MTAD
+HumanEval
+Llama-3-Instruct
+tokens/s $\uparrow$ 17.4 21.7 22.2 23.8 22.8 23.7 24.8
+J/token $\downarrow$ 10.0 8.1 7.8 7.8 7.9 7.8 7.6
+pass@1 $\uparrow$ 37.8 35.9 32.9 32.9 31.0 32.0 38.4
+Llama-3
+tokens/s $\uparrow$ 19.6 22.5 22.2 24.4 22.5 23.8 25.6
+J/token $\downarrow$ 9.7 8.9 8.9 8.9 8.1 7.9 7.6
+pass@1 $\uparrow$ 19.5 20.0 15.9 16.0 17.7 17.0 22.0
+Spider
+Llama-3-Instruct
+tokens/s $\uparrow$ 20.1 22.3 19.6 22.4 21.1 21.7 23.5
+J/token $\downarrow$ 10.2 9.5 10.5 9.6 10.2 10.0 9.2
+Acc $\uparrow$ 35.0 42.0 36.0 35.5 37.0 35.0 44.0
+Llama-3
+tokens/s $\uparrow$ 23.3 32.3 31.1 32.1 32.6 32.7 $\downarrow$ 33.3
+J/token 8.2 7.9 7.5 7.1 8.1 8.0 7.8
+Acc $\uparrow$ 30.5 29.5 21.5 23.0 21.5 24.0 35.0
+MT-Bench
+Llama-3-Instruct
+tokens/s $\uparrow$ 25.9 23.4 26.0 26.2 26.3 26.8 $\downarrow$ 29.8
+J/token $\downarrow$ 10.8 12.2 10.0 9.9 10.0 9.9 9.2
+score $\uparrow$ 4.15 4.26 4.10 4.11 4.01 4.02 4.40
+Llama-3
+tokens/s $\uparrow$ 24.5 22.3 24.1 24.5 24.5 25.7 28.2
+J/token $\downarrow$ 11.5 12.4 11.0 11.6 11.7 11.1 10.0
+score $\uparrow$ 3.41 3.24 3.39 3.41 3.35 3.36 3.75
 6 Related Work
@@
 Despite achieving significant speed-ups, a common drawback of these methods is that they have to sacrifice the model’s effectiveness.
-A direction closer to our work is non-autoregressive decoding.
+Non-autoregressive decoding more closely resembles our work.
 It is first proposed by (Gu et al., 2017) to generate multiple tokens in parallel.
 That is, the model simultaneously predicts $p(x_{t+k}|x_{1:t})$ ($k=1,2,\ldots$).
-Subsequent studies further improved the performance of parallel decoding by incorporating additional information (Wang et al., 2019; Sun et al., 2019; Li et al., 2019) or using additional iterations to refine predictions (Ghazvininejad et al., 2019; Lee et al., 2018; Guo et al., 2020).
-However, these works require continuous training of the model and usually either compromise the model effectiveness or require task-dependent techniques to achieve comparable performance (Kim et al., 2023).
+Subsequent studies further improved the performance of parallel decoding by incorporating additional information (Wang et al., 2019; Sun et al., 2019; Li et al., 2019) or employing additional iterations to refine predictions (Ghazvininejad et al., 2019; Lee et al., 2018; Guo et al., 2020).
+However, these works require continuous training of the model and generally either compromise the model effectiveness or require task-dependent techniques to attain a comparable performance (Kim et al., 2023).
 Speculative Decoding.
@@
 The large model rejects a token if its negative log-likelihood is larger than threshold $\tau_{2}$.
-SpecInfer (Miao et al., 2023) uses one or multiple small models to generate a draft token tree to increase the average acceptance length for each iteration.
+SpecInfer (Miao et al., 2023) utilizes one or multiple small models to generate a draft token tree to increase the average acceptance length for each iteration.
+MCSS (Yang et al., 2024) further strengthens SpecInfer via sampling without replacement.
 All these methods can be perceived as exact or approximate versions of sampling tokens from the conditional distribution $p(x_{t}|x_{<t})$.
 Therefore, their output perplexity is bounded by greedy decoding.
-An orthogonal direction to improve speculative decoding is to improve the effectiveness of the small draft model.
+An orthogonal direction to boost speculative decoding is to improve the effectiveness of the small draft model.
 It is obvious that if more draft tokens are accepted, the overall inference speed will increase.
-BiLD (Kim et al., 2023) uses a model prediction alignment technique to better train the small model.
+BiLD (Kim et al., 2023) employs a model prediction alignment technique to better train the small model.
 Liu et al.
@@
 7 Conclusion
-We introduce multi-token assisted decoding that significantly enhances output quality along with better time and energy efficiency.
-A distinctive aspect of our work is the exploration of speculative decoding’s impact on inference energy consumption, an often neglected area in existing studies.
+We introduce multi-token assisted decoding, a process that enhances output quality while improving time and energy efficiency A distinctive aspect of our work is the exploration of speculative decoding’s impact on inference energy consumption, an often neglected area in existing studies.
 This research contributes not only a novel decoding approach but also valuable insights for optimizing LLM deployment in real-world applications where considerations of both quality and efficiency are crucial.
+8 Acknowlegement
+This work was partially supported by NSF grants 2211557, 1937599, 2119643, 2303037, NSF 2312501, SRC JUMP 2.0 PRISM Center, NASA, Okawa Foundation, Amazon Research, Snapchat, and the CDSC industrial partners (https://cdsc.ucla.edu/partners/).
+The authors would also like to thank Marci Baun for editing the paper.
 Appendix A Proof
@@
 Therefore $x_{t+1:t+j}$ is always accepted if $j\leq\frac{|\log\tau|}{U}$. So $\bar{\gamma}\geq\frac{|\log\tau|}{U}$
-Appendix B Pseudocode of MJSD
+Appendix B Pseudocode of MTAD
 See Algorithm 1.
@@
 N/A [^2] MT-Bench contains multi-turn tasks where the input includes the responses of LLMs, so the input length is not fixed.
+HumanEval Coding 67
 Appendix C Additional Experiments
-C.1 Resutlts with OPT-30B and Llama-2-70B
+Table 7: Inference efficiency and output perplexity of different methods on ChatGPT-Prompt (CP), ChatAlpaca (CA), CNNDailyMail (CD), Spider (SP), and MT-Bench (MT) datasets.
+Bold numbers mark the best result, underlined numbers mark the second best.
+[…]
+[…]
+energy (J/token) $\downarrow$ 6.62$\pm$0.91 7.45$\pm$0.90 5.17$\pm$0.88 9.52$\pm$0.10 3.38$\pm$0.02
+perplexity $\downarrow$ 3.64$\pm$0.11 3.15$\pm$0.06 3.64$\pm$0.08 3.64$\pm$0.11 2.06$\pm$0.06
+OPT speed (token/s) $\uparrow$ 33.8$\pm$2.47 31.5$\pm$1.87 38.0$\pm$2.20 32.8$\pm$ 0.58 55.8$\pm$0.30
+energy (J/token) $\downarrow$ 7.48$\pm$0.07 8.75$\pm$0.13 6.08$\pm$0.11 10.3$\pm$1.49 3.61$\pm$0.03
+perplexity $\downarrow$ 5.47$\pm$0.11 4.51$\pm$0.09 5.27$\pm$0.09 5.12$\pm$0.01 3.00$\pm$0.09
+CA Llama-2 speed (token/s) $\uparrow$ 31.6$\pm$0.35 28.8$\pm$0.20 27.7$\pm$0.29 26.5$\pm$0.49 44.1$\pm$0.25
+energy (J/token) $\downarrow$ 6.98$\pm$0.15 7.99$\pm$0.15 7.20$\pm$0.08 7.52$\pm$0.32 4.72$\pm$0.03
+perplexity $\downarrow$ 2.13$\pm$0.03 1.95$\pm$0.03 2.15$\pm$0.01 2.15$\pm$0.01 1.88$\pm$0.05
+OPT speed (token/s) $\uparrow$ 35.6$\pm$0.45 38.5$\pm$0.93 28.4$\pm$0.34 31.4$\pm$0.39 49.6$\pm$0.42
+energy (J/token) $\downarrow$ 5.74$\pm$0.11 5.12$\pm$0.06 6.24$\pm$0.11 8.68$\pm$1.83 4.03$\pm$0.02
+perplexity $\downarrow$ 3.32$\pm$0.10 2.60$\pm$0.06 3.16$\pm$0.06 3.42$\pm$0.03 2.07$\pm$0.03
+CD Llama-2 speed (token/s) $\uparrow$ 30.7$\pm$0.18 30.5$\pm$0.21 25.0$\pm$0.31 24.6$\pm$0.06 44.2$\pm$0.99
+energy (J/token) $\downarrow$ 7.07$\pm$0.19 7.41$\pm$0.16 8.22$\pm$0.19 7.59$\pm$0.85 4.80$\pm$0.12
+perplexity $\downarrow$ 2.87$\pm$0.08 2.93$\pm$0.03 3.06$\pm$0.11 2.92$\pm$0.09 2.63$\pm$0.10
+OPT speed (token/s) $\uparrow$ 31.7$\pm$0.91 30.9$\pm$0.80 23.7$\pm$0.40 25.7$\pm$0.36 43.6$\pm$0.33
+energy (J/token) $\downarrow$ 6.37$\pm$0.11 6.71$\pm$0.17 7.31$\pm$0.17 8.03$\pm$0.63 4.86$\pm$0.03
+perplexity $\downarrow$ 3.97$\pm$0.06 3.74$\pm$0.09 4.04$\pm$0.07 3.92$\pm$ 0.34 3.17$\pm$0.06
+SP Llama-2 speed (token/s) $\uparrow$ 24.0$\pm$0.28 26.2$\pm$0.08 24.2$\pm$0.29 23.8$\pm$0.20 26.4$\pm$0.28
+energy (J/token) $\downarrow$ 10.75$\pm$0.02 9.84$\pm$0.07 11.0$\pm$0.08 11.0$\pm$0.76 9.01$\pm$0.07
+perplexity $\downarrow$ 2.26$\pm$0.01 2.13$\pm$0.03 2.29$\pm$0.04 2.29$\pm$0.03 1.87$\pm$0.03
+OPT speed (token/s) $\uparrow$ 24.6$\pm$0.30 29.9$\pm$0.55 19.8$\pm$0.13 24.1$\pm$0.10 34.4$\pm$0.46
+energy (J/token) $\downarrow$ 15.6$\pm$3.55 13.6$\pm$3.07 20.1$\pm$2.52 16.9$\pm$2.75 11.7$\pm$2.36
+perplexity $\downarrow$ 2.30$\pm$0.07 1.90$\pm$0.01 2.20$\pm$0.09 2.21$\pm$ 0.01 1.63$\pm$0.03
+MT Llama-2 speed (token/s) $\uparrow$ 23.0$\pm$1.10 23.7$\pm$1.43 19.1$\pm$2.71 23.7$\pm$2.03 29.4$\pm$2.71
+energy (J/token) $\downarrow$ 7.99$\pm$0.26 7.40$\pm$0.19 9.27$\pm$0.54 9.20$\pm$0.73 6.71$\pm$1.19
+perplexity $\downarrow$ 3.64$\pm$0.51 3.44$\pm$0.76 3.64$\pm$0.51 3.63$\pm$0.50 2.21$\pm$0.18
+OPT speed (token/s) $\uparrow$ 34.0$\pm$3.00 44.7$\pm$2.92 28.7$\pm$2.46 28.5$\pm$2.74 48.0$\pm$1.80
+energy (J/token) $\downarrow$ 12.1$\pm$0.36 6.23$\pm$0.67 12.9$\pm$1.73 13.2$\pm$1.88 6.11$\pm$0.82
+perplexity $\downarrow$ 2.02$\pm$0.40 1.50$\pm$0.27 1.97$\pm$0.38 1.99$\pm$ 0.33 1.10$\pm$0.03
+[…]
+Table 8: Average number of tokens generated at each iteration across all datasets.
+Llama-2 OPT
+SpD 2.02$\pm$0.05 2.60$\pm$0.06
+BiLD 1.83$\pm$0.10 2.68$\pm$0.36
+Spectr 2.73$\pm$0.43 3.45$\pm$0.42
+SpecInfer 2.74$\pm$0.46 3.45$\pm$0.40
+MTAD 3.17$\pm$0.43 4.30$\pm$0.03
+Table 9: Downstream task scores of speculative decoding and MTAD. All the scores are higher the better.
+SpD MTAD
+CD Rouge-L 0.114 0.118
+SP EA 11.5 13.0
+MT Humanities 2.95 3.15
+Extraction 1.80 2.50
+Roleplay 3.10 3.80
+Math 1.10 1.00
+Coding 1.25 1.10
+Reasoning 3.80 3.15
+STEM 2.85 3.10
+Writing 3.80 3.65
+Average 2.58 2.68
+C.1 Additional Datasets and Model Family
+Here we report the additional experiment results with three more datasets: (1) ChatGPT-Prompt (Rashad, 2023), (2) ChatAlpaca (Bian et al., 2023), (3) CNN Dailymail (See et al., 2017).
+We use two public LLM families in our experiments: OPT (Zhang et al., 2022) and Llama-2 (Touvron et al., 2023).
+We set the large model to be OPT-13B and Llama-2-13B as they are the largest models that can run on a single 40GB GPU, and utilize Llama-68M (Miao et al., 2023) and OPT-125M as the small models.
+Table 7 shows the full evaluation results, and Table 9 displays the downstream performance.
+Table 8 depicts the average number of generated token per iteration for different algorithms.
+The experiment results demonstrate that MTAD achieves better efficiency, better perplexity, as well as better downstream performance.
+Table 10: Number of Generated Tokens Per Iteration Across Thresholds
+Threshold
+0.1 0.3 0.5 0.7 0.9
+Llama-3 4.72 4.43 4.12 3.93 3.64
+Llama-3-Instruct 4.74 4.51 4.28 4.06 3.79
+C.2 Ablation Studies on Acceptance Threshold
+Figure 3: Effect of acceptance threshold on output perplexity and decoding speed.
+Next, we evaluate the effect of acceptance threshold $\tau$.
+Intuitively, when we increase $\tau$ from 0 to 1, the acceptance criterion becomes stricter, the efficiency drops while the output perplexity increases.
+Surprisingly, this expectation is only partially correct.
+As shown in Figure 3, the efficiency indeed drops when $\tau$ increases.
+However, the perplexity rises when $\tau$ is close to 1.
+When $\tau=1$, all the draft tokens are rejected, which makes MTAD equivalent to multinomial sampling.
+Similarly, when $\tau$ is close to 1, the advantage of multi-token joint decoding on effectiveness disappears, hence MTAD’s perplexity is similar to the perplexity of multinomial sampling.
+Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MTAD is relatively stable, suggesting that MTAD is not sensitive to the acceptance threshold.
+In addition, Table 10 exhibits how the number of accepted tokens change with thresholds.
+We can see that the number of accepted tokens grows smaller when the threshold increases, which explains the speed degradation.
+C.3 Ablation Studies on Number of Beams
+In Table 11, we investigate how the number of beams used in the beam decoding of the small model affects the inference performance.
+Increasing the number of beams boosts the quality of the draft tokens, which not only improves the output perplexity but also increases the average acceptance length and hence leads to better efficiency.
+But we can see that the increment slows down when the number of beams is too large.
+In addition, when the number of beams is too large, the inference cost of the small model will become too high.
+Table 11: Effect of beam count on inference performance (ChatGPT-Prompts dataset). Arrows indicate optimal direction ($\uparrow$=higher better, $\downarrow$=lower better).
+Number of Beams
+2 4 6 8
+Llama-2 Speed (tok/s) $\uparrow$ 55.9 59.9 60.2 61.3
+Energy (J/tok) $\downarrow$ 2.43 2.25 2.22 2.20
+Perplexity $\downarrow$ 2.44 2.12 2.14 2.10
+OPT Speed (tok/s) $\uparrow$ 51.0 54.1 54.3 55.9
+Energy (J/tok) $\downarrow$ 2.50 2.32 2.36 2.30
+Perplexity $\downarrow$ 3.63 3.16 3.42 3.19
+C.4 Ablation study of top-k and top-p sampling
+Table 12 demonstrates how the value of k and p in top-k and top-p warping affects our method. We can see that by changing the value of $k$ and $p$, MTAD consistently performs significantly better.
+Table 12: Ablation study of $k$ and $p$ in top-k and top-p sampling
+K P Multinomial SpD MTAD
+PPL Tokens/sec PPL Tokens/sec PPL Tokens/sec
+20 0.9 3.74 22.6 3.64 36.8 2.06 63.0
+20 0.8 3.06 22.7 3.10 38.5 1.93 58.8
+10 0.9 3.03 22.7 3.22 38.5 1.95 62.5
+10 0.8 2.56 22.7 2.53 40.0 1.80 62.5
+C.5 Resutlts with OPT-30B and Llama-2-70B
 Here we report the performances of different methods for OPT (350M and 30B) and Llama-2-Chat (7B and 70B).
 Table 13 shows the average performances across all datasets.
-MTAD always achieves the lowest perplexity and the best efficiency.
+MTAD always realizes the lowest perplexity and the best efficiency.
 Table 13: Inference efficiency and output perplexity of different methods with OPT (350M,30B) and Llama-2-Chat (7B,70B).
@@
 […]
-C.2 Ablation study of top-k and top-p sampling
-Here we conduct an ablation study to show how the value of $k$ and $p$ in top-k and top-p warping affects our method.
-Table 10 shows the results.
-We can see that when changing the value of $k$ and $p$, MTAD consistently achieves significantly better performances.
-Table 10: Ablation study of $k$ and $p$ in top-k and top-p sampling
-K P Greedy Speculative MJSD
-PPL Tokens/sec PPL Tokens/sec PPL Tokens/sec
-20 0.9 3.74 22.6 3.64 36.8 2.06 63.0
-20 0.8 3.06 22.7 3.10 38.5 1.93 58.8
-10 0.9 3.03 22.7 3.22 38.5 1.95 62.5
-10 0.8 2.56 22.7 2.53 40.0 1.80 62.5
-C.3 Additional Experiments on CNNDM and Spider
-We also report the downstream effectiveness of our method on CNNDM and Spider when the small model and large model are fine-tuned on the dataset.
-Table 11 and Table 12 show the results.
-We can see that MTAD consistenly achieve better effectiveness as well as faster decoding speed.
+C.6 Additional Experiments on CNNDM and Spider
+Table 14 and Table 15 reveal the downstream effectiveness of our method on CNNDM and Spider when the small model and large model are fine-tuned on the dataset.
+We can see that MTAD is consistently more effective and has a faster decoding speed.
 Table 14: Comparison of ROUGE-L Scores and Tokens per Second under Different Fine-Tuning Conditions on CNNDM
@@
 Method ROUGE-L Tokens/sec ROUGE-L Tokens/sec ROUGE-L Tokens/sec
-Speculative 0.114 37.7 0.114 20.4 0.164 24.3
-MJSD 0.118 44.2 0.121 25.0 0.168 27.1
+SpeD 0.114 37.7 0.114 20.4 0.164 24.3
+MTAD 0.118 44.2 0.121 25.0 0.168 27.1
 Table 15: Comparison of Execution Accuracy (EA) and Tokens per Second under Different Fine-Tuning Conditions on Spider
@@
 Method EA Tokens/sec EA Tokens/sec EA Tokens/sec
-Speculative 11.5 28.5 11.5 27.8 16.3 25.6
-MJSD 13.0 30.3 14.8 32.3 18.3 29.4
+SpD 11.5 28.5 11.5 27.8 16.3 25.6
+MTAD 13.0 30.3 14.8 32.3 18.3 29.4
+C.7 Visualization of perplexity and output quality
+To further illustrate the relationship between perplexity and downstream performance, we present a scatter plot in Figure 4.
+The plot depicts the correlation between relative downstream scores (normalized by the score of multinomial sampling) and relative perplexity (normalized by the perplexity of multinomial sampling) across 7 decoding algorithms, 3 datasets, and 2 model configurations.
+The results confirm that lower perplexity generally correlates with higher output quality.
+Figure 4: Relationship between relative perplexity (normalized by multinomial sampling’s perplexity) and relative performance score (normalized by multinomial sampling’s score).
+C.8 Correlation between Energy and Speed
+Figure 5 shows there is a correlation between speed and energy, whether we consider the entire table or focus on a specific dataset and model.
+For fairness, all methods for a given dataset and model were run on the same machine nodes.
+However, for a fixed method (e.g., Spectr), experiments on different datasets and models might be conducted on different nodes (all equipped with L40 GPUs).
+We did notice that the same configuration run on different machines may have varied energy consumption.
+This variation introduces some randomness, which could make the correlation appear less consistent across datasets and models.
+(a) group by dataset and model
+(b) group by methods
+Figure 5: Correlation between speed and energy
 Appendix D Energy Consumption Measurement
-We use the command "nvidia-smi --query-gpu=power.draw --format=csv" to get GPU power every second, and sum them up as the total energy consumption.
+To get GPU power every second, we run the command “nvidia-smi -query-gpu=power.draw -format=csv”.
+We add the results up to determine the total energy consumption.
 We use average energy consumption per token to measure energy efficiency.
 There is a recent study pointing out the measurement error using nvidia-smi (Yang et al., 2023).
-We follow the three principles proposed in (Yang et al., 2023) to minimize the error.
+We follow the three principles proposed in (Yang et al., 2023) to minimize this error.
 Appendix E Configuration
@@
 We set the maximum running time to be an hour for each baseline.
-We use average tokens/second to measure the inference speed and use perplexity (exponentiated average negative log-likelihood) based on the probability of the large model to measure the output quality.
-Because different methods might finish different numbers of inputs, we only calculate the perplexity of the first $M$ inputs, where $M$ is the number of inputs finished by greedy decoding.
-We use average energy consumption per token to measure energy efficiency.
-The details of energy measurement are illustrated in the Appendix.
+We use average tokens/second to measure the inference speed and use average energy consumption per token to measure energy efficiency.
 Appendix F Hyper-parameter Details
-In the experiments, we follow the default settings in (Bear, 2024) to warp the sampling distribution $p$ and $q$ with the following steps, which are the default warpping operations in a public speculative decoding implementation.
-  1. 1.
-    Keep the probabilities of top 20 tokens unchanged, and set the probabilities of other tokens to 0, then normalize the distribution.
-  2. 2.
-    Sort the tokens based on their distributions descendingly. Keep the first $K$ tokens such that their cumulative probabilities is larger than 0.9, while set the probabilities of other tokens to 0.
-For different methods, we choose their hyper-parameters by using a small validation set.
+In the experiments, we follow the settings in (Bear, 2024) to warp the sampling distribution $p$ and $q$ with the following steps, which are the default warping operations in a public speculative decoding implementation.
+Specifically, we first keep the probabilities of top 10 tokens unchanged, and set the probabilities of other tokens to 0, then normalize the distribution.
+Then we sort the tokens based on their distributions in descending order and keep the first $K$ tokens such that their cumulative probabilities is larger than 0.9, while set the probabilities of other tokens to 0.
+For different methods, we choose their hyper-parameters based on a small validation set.
 We select the set of hyper-parameters that make the corresponding method have best output perplexity.
-Table 13 shows the hyper-parameters used in the experiments.
-Table 13: Hyper-parameters of different methods for different models and datasets. L: Llama, O:OPT, CP: ChatGPT-Prompts, CA: ChatAlpaca, CD: CNNDaily.
-L,CP O,CP L,CA O,CA L,CD O,CD L,SP O,SP L,MT O,MT
-speculative step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
-Spectr&SpecInfer step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
-num seq $m$ 4 4 2 2 4 2 2 2 2 2
-BiLD step len $\gamma$ 10 10 10 10 10 10 10 10 10 10
-fallback thres $\tau_{1}$ 0.9 0.9 0.9 0.3 0.9 0.3 0.9 0.9 0.9 0.9
-rollback thres $\tau_{2}$ 2 2 1 2 3 2 1 1 1 1
-MTAD step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
-num beams 8 8 8 8 8 8 8 8 8 8
-acc/rej thres $\tau$ 0.1 0.1 0.1 0.1 0.1 0.1 0.9 0.1 0.9 0.1
+For MTAD, we choose the beam width from $\{4,8\}$, the number of draft tokens from $\{3,4\}$, and the acceptance threshold from $\{0.1,0.3,0.5,0.7,0.9\}$.
 Appendix G License of Datasets and Models
@@
   • CNN Dailymail: Apache-2.0 License
+  • MTBench: Apache-2.0 License
+  • Spider: CC BY-SA 4.0 License
+  • HumanEval: MIT License
 Models

### v4

@@
 Abstract
-Large language models (LLMs) have achieved remarkable success across diverse tasks, but, due to single-token generation at each decoding step, their inference processes are hindered by substantial time and energy demands.
-While previous methods such as speculative decoding mitigate these inefficiencies by producing multiple tokens per step, each token is still generated by its single-token distribution.
-Although this enhances the speed, it does not improve the output quality.
-In contrast, our work simultaneously boosts inference speed and improves the output effectiveness.
-We consider multi-token joint decoding (MTJD), which generates multiple tokens from their joint distribution at each iteration, theoretically reducing perplexity and raising task performance.
-However, MTJD suffers from the high cost of sampling from the joint distribution of multiple tokens.
-Inspired by speculative decoding, we introduce multi-token assisted decoding (MTAD), a novel framework designed to accelerate MTJD.
-MTAD leverages a smaller auxiliary model to approximate the joint distribution of a larger model, incorporating a verification mechanism that not only ensures the accuracy of this approximation, but also increases the decoding efficiency over conventional speculative decoding.
-Theoretically, we demonstrate that MTAD closely approximates exact MTJD with a bounded error.
-Empirical evaluations across various tasks reveal that MTAD improves downstream performance by 25% compared to standard single-token sampling.
-Furthermore, MTAD achieves a 1.42$\times$ speed-up and consumes 1.54$\times$ less energy than vanilla speculative decoding methods.
-These results highlight MTAD’s ability to make multi-token joint decoding both effective and efficient, promoting more productive and high-performance deployment of LLMs.
-[^1] We release our code at https://github.com/ZongyueQin/MTAD
+[…]
 1 Introduction
@@
   4. 4.
-    We conducted comprehensive evaluations across various tasks, demonstrating that MTAD improves downstream performance by 25% compared to standard single-token sampling, while, at the same time, realizing a 1.42$\times$ speed-up and paring energy consumption by 1.54$\times$ compared to vanilla speculative decoding methods.
+[…]
 These advancements position MTAD as a robust solution for making multi-token joint decoding both effective and efficient, thereby facilitating more sustainable and high-performance deployment of large-scale language models.
@@
 For example, beam sampling (explained below) aims to return output with lower perplexity and is empirically proven to have better downstream performance in general (Shi et al., 2024).
-To further demonstrate the relationship between perplexity and downstream performance, we evaluate GPT-3.5-turbo on the spider (Yu et al., 2018) dataset.
-Employing a temperature of 2, the model generated 10 outputs for each input.
-We measured the average perplexities and execution accuracies for the outputs with the highest, lowest, and median (the 5th lowest) perplexity.
-As seen in Table 1, lower perplexity correlates with improved downstream performance, even in one of today’s largest models.
 Table 1: Relationship between perplexity and execution accuracy (EA, higher the better) for GPT-3.5-turbo.
@@
 Lowest PPL 1.07 62
+To further demonstrate the relationship between perplexity and downstream performance, we evaluate GPT-3.5-turbo on the spider (Yu et al., 2018) dataset.
+Employing a temperature of 2, the model generated 10 outputs for each input.
+We measured the average perplexities and execution accuracies for the outputs with the highest, lowest, and median (the 5th lowest) perplexity.
+As seen in Table 1, lower perplexity correlates with improved downstream performance, even in one of today’s largest models.
 Now we introduce commonly used decoding approaches.
@@
 Meanwhile, when the beam width for the auxiliary model is large enough, $\epsilon\rightarrow 0$, and the ratio bound in Theorem 3.5 converges to 1, This implies that MTAD converges to MTJD under these limiting conditions.
-Similar to Spectr (Sun et al., 2023) and SpecInfer (Miao et al., 2023), it is possible to enhance the number of accepted tokens in MTAD by allowing the draft model to produce multiple draft sequences and by applying tree-based attention (Miao et al., 2023) for simultaneous verification.
-However, our preliminary experiment results suggest that since MTAD already selects the longest accepted prefix sub-sequence, the advantage of generating multiple draft tokens is less significant.
-Moreover, this approach increases the memory cost during inference and may affect the error bounds derived above.
-Therefore, we leave a more detailed exploration of this extension as future work.
+3.3 Multi Candidate Verification
+As illustrated in Figure 3, the intermediate results of beam sampling naturally form a tree structure, where each layer contains $b$ nodes corresponding to the $b$ intermediate beams generated at that step.
+In vanilla MTAD, only the final output sequence is retained and verified by the target model, while all other intermediate beams are discarded.
+However, these discarded beams may in fact have higher likelihood under the target model.
+To address this limitation, we propose an enhanced version of MTAD that leverages all intermediate beams during verification.
+This optimization introduces two key benefits: (1) Improved efficiency.
+By incorporating more candidates into the verification process, the probability of accepting tokens at each step increases, which leads to faster decoding.
+(2) Enhanced output quality.
+Among the accepted candidates, we can select the one with the highest target likelihood, potentially yielding better generations than vanilla MTAD.
+To efficiently verify all intermediate beams, we adopt the tree attention mechanism introduced by Miao et al.
+(2023), which allows the target model to compute the conditional likelihood of every token in the draft tree in a single forward pass.
+Specifically, at each layer $i$, we evaluate the target likelihood $p(x_{t+1:t+i}^{(j)}\mid x_{1:t})$ for each beam $j\in{1,\dots,b}$.
+In vanilla MTAD, a beam sequence $x_{t+1:t+i}$ is accepted if its likelihood ratio $\frac{p(x_{t+1:t+i}\mid x_{1:t})}{q(x_{t+1:t+i}\mid x_{1:t})}\geq\tau$, which is a reasonable criterion when $x_{1:t}$ corresponds to a prefix of the final output.
+However, when verifying all intermediate beams—particularly with a large beam width $b$, many beams may have low draft likelihood $q(x_{t+1:t+i}^{(j)}\mid x_{1:t})$, making the denominator small and potentially resulting in the acceptance of low-quality candidates.
+To mitigate this issue, we revise the acceptance criterion: instead of comparing each beam against its own draft likelihood, we normalize all beams against the draft likelihood of the highest-likelihood final beam.
+Specifically, we accept a candidate $x_{t+1:t+i}^{(j)}$ if
+$\frac{p(x_{t+1:t+i}^{(j)}|x_{1:t})}{q(x_{t+1:t+i}^{**}|x_{1:t})}\geq\tau$ (9)
+where $x_{t+1:t+\gamma}^{**}$ is the output sequence of the beam sampling, which is the beam with the highest draft likelihood at the last step.
+Once all beams in the tree are verified, we select the longest accepted sequence and generate one additional token from the target model $p$ to be the output tokens at this iteration.
+If there are multiple accepted sequences with same lengths, the sequence with the highest target likelihood is chosen.
+This variant, which we call multi-candidate MTAD (MMTAD), improves both the robustness and effectiveness of decoding.
+The following theorem demonstrates that MMTAD achieves a higher acceptance rate at each decoding step compared to vanilla MTAD, which directly translates to better decoding efficiency.
+Theorem 3.7.
+MMTAD has a higher expected accepted sequence length than vanilla MTAD.
+Proof.
+Let $x^{**}_{t+1:t+\gamma}$ denote the output sequence of beam sampling.
+The accepted length of vanilla MTAD is $\max\{i:\frac{p(x^{**}_{t+1:t+i}|x_{1:t})}{q(x^{**}_{t+1:t+i}|x_{1:t})}\}\geq\tau\}$.
+Notice that MMTAD also verifies $x^{**}$ and the acceptance criterion does not change for any $x^{**}_{t+1:t+i}$.
+Meanwhile, MMTAD verifies more candidates which might pass the verification when $x^{**}_{t+1:t+i}$ fails.
+Therefore, MMTAD has a higher expected accepted length than MTAD.
+∎
+Additionally, the next theorem confirms that MMTAD also has bounded error on the perplexity ratio.
+Theorem 3.8.
+Let $x_{1:\Gamma_{N}}^{\text{multi}}$ be the sequence generated by multi-candidate MTAD (MMTAD), and $x_{1:\Gamma_{N}}^{*}$ be the sequence generated by deterministic exact MTJD.
+Under the same assumption as Theorem 3.5, i.e., $\forall x_{1:t}\in\mathcal{X}$, $|\log p(x|x_{1:t})-\log q(x|x_{1:t})|_{\infty}\leq U$, the perplexity ratio of MMTAD is bounded as:
+$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}}^{\text{multi}})}{PPL_{p}(x_{1:\Gamma_{N}}^{*})}\leq\tau^{-\frac{1}{\bar{\gamma}}}(1-\epsilon)^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{(1-\epsilon\bar{\gamma})H(p)+(1-\epsilon+\bar{\gamma})U}{\bar{\gamma}}\right)$ (10)
+The proof is provided in the Appendix A.
+Although Theorem 3.8 indicates that MMTAD has a slightly looser perplexity ratio bound than vanilla MTAD, due to the additional $(1-\epsilon)^{-\frac{1}{\bar{\gamma}}}$ term, in practice, MMTAD often yields higher-quality outputs.
+This is because it selects the longest verified sequence at each step, and in the case of ties, chooses the one with the highest target model likelihood.
+This strategy prioritizes more promising candidates and typically results in better generations than vanilla MTAD.
+(a) The draft tree of beam sampling
+(b) tree attention mask
+(c) MMTAD verification
+Figure 3: Illustration of MMTAD: (a) All intermediate beams of beam sampling naturally form a tree.
+Vanilla MTAD only verify the output beam (yellow blocks), MMTAD verify all the beams.
+(b) MMTAD utilizes tree attention to efficiently compute the target likelihood of each beam.
+(c) MMTAD selects the longest accepted sequence with the highest target likelihood to return.
 4 Energy Efficiency Analysis
@@
 Efficiency Analysis.
-We first observe that MTAD is the most efficient among all baselines in terms of both energy and time.
-On average, MTAD is 17.2% faster than the most efficient lossless baseline, MCSS, while consuming 6.1% less energy.
-Compared to lossy decoding algorithms, it attains 27.2% (15.1%) higher speed and 17.9% (14.1%) lower energy consumption than BiLD (typical decoding), respectively.
+We first observe that MMTAD is the most efficient among all baselines in terms of both energy and time.
+On average, MMTAD is 21.1% faster than the most efficient lossless baseline, MCSS, while consuming 23.4% less energy.
+Compared to lossy decoding algorithms, it attains 46.8% (31.5%) higher speed and 30.4% (28.3%) lower energy consumption than BiLD (typical decoding), respectively.
 Interestingly, despite utilizing only a single draft sequence, MTAD outperforms baselines that employ multiple draft sequences, such as Spectr, MCSS, and SpecInfer.
@@
 This suggests that all lossy decoding methods can trade off efficiency for performance by adjusting verification strictness.
-Most notably, MTAD consistently achieves the highest downstream performance.
-On average, it surpasses lossless decoding algorithms by 25.0%, BiLD by 11.8%, and typical decoding by 9.9%.
-These results confirm our claim that MTAD offers superior effectiveness compared to conventional decoding methods that rely solely on single-token distributions.
+Most notably, MMTAD consistently achieves the highest downstream performance.
+On average, it surpasses lossless decoding algorithms by 42.7%, BiLD by 27.3%, and typical decoding by 24.8%.
+These results confirm our claim that MMTAD offers superior effectiveness compared to conventional decoding methods that rely solely on single-token distributions.
 Table 5: Comparison of different speculative decoding methods across various models and metrics. Bold indicates best values, underline indicates second-best.
 Lossy Decoding Lossless Decoding Ours
-BiLD Typical SpD Spectr SpecInfer MCSS MTAD
+BiLD Typical SpD Spectr SpecInfer MCSS MTAD MMTAD
 HumanEval
 Llama-3-Instruct
-tokens/s $\uparrow$ 17.4 21.7 22.2 23.8 22.8 23.7 24.8
-J/token $\downarrow$ 10.0 8.1 7.8 7.8 7.9 7.8 7.6
-pass@1 $\uparrow$ 37.8 35.9 32.9 32.9 31.0 32.0 38.4
+tokens/s $\uparrow$ 17.4 21.7 22.2 23.8 22.8 23.7 28.2 29.7
+J/token $\downarrow$ 10.0 8.1 7.8 7.8 7.9 7.8 5.6 5.5
+pass@1 $\uparrow$ 37.8 35.9 32.9 32.9 31.0 32.0 43.2 45.1
 Llama-3
-tokens/s $\uparrow$ 19.6 22.5 22.2 24.4 22.5 23.8 25.6
-J/token $\downarrow$ 9.7 8.9 8.9 8.9 8.1 7.9 7.6
-pass@1 $\uparrow$ 19.5 20.0 15.9 16.0 17.7 17.0 22.0
+tokens/s $\uparrow$ 19.6 22.5 22.2 24.4 22.5 23.8 27.5 29.1
+J/token $\downarrow$ 9.7 8.9 8.9 8.9 8.1 7.9 6.1 6.1
+pass@1 $\uparrow$ 19.5 20.0 15.9 16.0 17.7 17.0 26.8 28.0
 Spider
 Llama-3-Instruct
-tokens/s $\uparrow$ 20.1 22.3 19.6 22.4 21.1 21.7 23.5
-J/token $\downarrow$ 10.2 9.5 10.5 9.6 10.2 10.0 9.2
-Acc $\uparrow$ 35.0 42.0 36.0 35.5 37.0 35.0 44.0
+tokens/s $\uparrow$ 20.1 22.3 19.6 22.4 21.1 21.7 22.8 25.5
+J/token $\downarrow$ 10.2 9.5 10.5 9.6 10.2 10.0 8.1 7.7
+Acc $\uparrow$ 35.0 42.0 36.0 35.5 37.0 35.0 44.0 54.0
 Llama-3
-tokens/s $\uparrow$ 23.3 32.3 31.1 32.1 32.6 32.7 $\downarrow$ 33.3
-J/token 8.2 7.9 7.5 7.1 8.1 8.0 7.8
-Acc $\uparrow$ 30.5 29.5 21.5 23.0 21.5 24.0 35.0
+tokens/s $\uparrow$ 23.3 32.3 31.1 32.1 32.6 32.7 37.9 43.4
+J/$\downarrow$ 8.2 7.9 7.5 7.1 8.1 8.0 6.0 5.5
+Acc $\uparrow$ 30.5 29.5 21.5 23.0 21.5 24.0 39.0 40.0
 MT-Bench
 Llama-3-Instruct
-tokens/s $\uparrow$ 25.9 23.4 26.0 26.2 26.3 26.8 $\downarrow$ 29.8
-J/token $\downarrow$ 10.8 12.2 10.0 9.9 10.0 9.9 9.2
-score $\uparrow$ 4.15 4.26 4.10 4.11 4.01 4.02 4.40
+tokens/s $\uparrow$ 25.9 23.4 26.0 26.2 26.3 26.8 29.8 32.9
+J/token $\downarrow$ 10.8 12.2 10.0 9.9 10.0 9.9 9.2 7.5
+score $\uparrow$ 4.15 4.26 4.10 4.11 4.01 4.02 4.40 4.39
 Llama-3
-tokens/s $\uparrow$ 24.5 22.3 24.1 24.5 24.5 25.7 28.2
-J/token $\downarrow$ 11.5 12.4 11.0 11.6 11.7 11.1 10.0
-score $\uparrow$ 3.41 3.24 3.39 3.41 3.35 3.36 3.75
+tokens/s $\uparrow$ 24.5 22.3 24.1 24.5 24.5 25.7 28.2 29.8
+J/token $\downarrow$ 11.5 12.4 11.0 11.6 11.7 11.1 10.0 10.1
+score $\uparrow$ 3.41 3.24 3.39 3.41 3.35 3.36 3.75 3.75
+5.3 Ablation Studies
+5.3.1 Draft Sequence Length
+We investigate the impact of the draft sequence length $\gamma$ on the performance of MMTAD.
+Figure 4 presents results for decoding speed, block efficiency (i.e., the average number of tokens generated per iteration), and output perplexity using the Llama-3-8B-Instruct model on the Spider dataset, with $\gamma\in\{3,4,5,6,7,8,9,10\}$.
+As $\gamma$ increases, block efficiency consistently improves.
+However, decoding speed saturates once $\gamma$ reaches 7.
+This is due to the growing computational overhead associated with generating and verifying longer draft sequences, which offsets the gains from improved block efficiency.
+Meanwhile, as shown in Figure 4(c), output perplexity remains stable across different values of $\gamma$.
+(a) $\gamma$ vs. Speed
+(b) $\gamma$ vs. Block efficiency
+(c) $\gamma$ vs. Perplexity
+Figure 4: Performance of MMTAD when draft length $\gamma\in\{3,4,5,6,7,8,9,10\}$.
+5.3.2 Beam Sampling Width
+Next, we study the effect of beam width $b$ used by the draft model when generating draft sequences for MMTAD.
+As shown in Figure 5, increasing $b$ leads to a slight improvement in block efficiency.
+This is because having more candidate beams increases the likelihood that more tokens will be accepted during verification.
+Additionally, output perplexity shows a slight decrease as $b$ increases, since MMTAD selects the longest accepted sequence with the highest likelihood under the target model.
+We also observe that decoding speed initially improves with larger $b$, owing to gains in block efficiency, but plateaus once $b$ reaches 5.
+This is due to the growing cost of verifying a larger number of candidate beams, which offsets the speedup from increased acceptance.
+(a) $b$ vs. Speed
+(b) $b$ vs. Block efficiency
+(c) $b$ vs. Perplexity
+Figure 5: Performance of MMTAD when beam width $b\in\{2,3,4,5,6\}$.
+5.3.3 Acceptance Threshold
+Finally, we examine the impact of the acceptance threshold $\tau$.
+As shown in Figure 6, increasing $\tau$ imposes a stricter acceptance criterion, leading to lower block efficiency and reduced decoding speed.
+This is expected, as fewer draft sequences satisfy the higher acceptance threshold.
+On the other hand, output perplexity decreases as $\tau$ increases, since stricter acceptance favors higher-confidence predictions.
+However, this improvement saturates as $\tau$ approaches 1, with diminishing returns in perplexity reduction.
+(a) $\tau$ vs. Speed
+(b) $\tau$ vs. Block efficiency
+(c) $\tau$ vs. Perplexity
+Figure 6: Performance of MMTAD when acceptance threshold $\tau\in\{0.1,0.3,0.5,0.7,0.9\}$.
 6 Related Work
@@
 Therefore $x_{t+1:t+j}$ is always accepted if $j\leq\frac{|\log\tau|}{U}$. So $\bar{\gamma}\geq\frac{|\log\tau|}{U}$
+A.6 Proof of Theorem 3.8
+Proof.
+Let $x^{**}$ be the output sequence of beam sampling. We have
+$\frac{p(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}{q(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}\geq\frac{p(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}{\max_{x^{{}^{\prime}}_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}}q(x^{{}^{\prime}}_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}$ (34)
+$\geq$ $\frac{(1-\epsilon)p(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}{q(x^{**}_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}\quad(\text{assumption of beam sampling error})$
+$\geq$ $(1-\epsilon)\tau$
+With the same procedure in the proof of Lemma A.1, we have
+$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\gamma_{N}})}{PPL_{q}(x_{1:\gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}(1-\epsilon)^{-\frac{1}{\bar{\gamma}}}$ (35)
+Then, with the same procedure to prove Theorem 3.5, we have
+$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{p}(x^{*}_{1:\Gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}(1-\epsilon)^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{(1-\epsilon\bar{\gamma})H(p)+(1-\epsilon+\bar{\gamma})U}{\bar{\gamma}}\right)$ (36)
 Appendix B Pseudocode of MTAD
@@
 The experiment results demonstrate that MTAD achieves better efficiency, better perplexity, as well as better downstream performance.
-Table 10: Number of Generated Tokens Per Iteration Across Thresholds
-Threshold
-0.1 0.3 0.5 0.7 0.9
-Llama-3 4.72 4.43 4.12 3.93 3.64
-Llama-3-Instruct 4.74 4.51 4.28 4.06 3.79
-C.2 Ablation Studies on Acceptance Threshold
-Figure 3: Effect of acceptance threshold on output perplexity and decoding speed.
-Next, we evaluate the effect of acceptance threshold $\tau$.
-Intuitively, when we increase $\tau$ from 0 to 1, the acceptance criterion becomes stricter, the efficiency drops while the output perplexity increases.
-Surprisingly, this expectation is only partially correct.
-As shown in Figure 3, the efficiency indeed drops when $\tau$ increases.
-However, the perplexity rises when $\tau$ is close to 1.
-When $\tau=1$, all the draft tokens are rejected, which makes MTAD equivalent to multinomial sampling.
-Similarly, when $\tau$ is close to 1, the advantage of multi-token joint decoding on effectiveness disappears, hence MTAD’s perplexity is similar to the perplexity of multinomial sampling.
-Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MTAD is relatively stable, suggesting that MTAD is not sensitive to the acceptance threshold.
-In addition, Table 10 exhibits how the number of accepted tokens change with thresholds.
-We can see that the number of accepted tokens grows smaller when the threshold increases, which explains the speed degradation.
-C.3 Ablation Studies on Number of Beams
-In Table 11, we investigate how the number of beams used in the beam decoding of the small model affects the inference performance.
-Increasing the number of beams boosts the quality of the draft tokens, which not only improves the output perplexity but also increases the average acceptance length and hence leads to better efficiency.
-But we can see that the increment slows down when the number of beams is too large.
-In addition, when the number of beams is too large, the inference cost of the small model will become too high.
-Table 11: Effect of beam count on inference performance (ChatGPT-Prompts dataset). Arrows indicate optimal direction ($\uparrow$=higher better, $\downarrow$=lower better).
-Number of Beams
-2 4 6 8
-Llama-2 Speed (tok/s) $\uparrow$ 55.9 59.9 60.2 61.3
-Energy (J/tok) $\downarrow$ 2.43 2.25 2.22 2.20
-Perplexity $\downarrow$ 2.44 2.12 2.14 2.10
-OPT Speed (tok/s) $\uparrow$ 51.0 54.1 54.3 55.9
-Energy (J/tok) $\downarrow$ 2.50 2.32 2.36 2.30
-Perplexity $\downarrow$ 3.63 3.16 3.42 3.19
-C.4 Ablation study of top-k and top-p sampling
+C.2 Ablation study of top-k and top-p sampling
 Table 10 demonstrates how the value of k and p in top-k and top-p warping affects our method. We can see that by changing the value of $k$ and $p$, MTAD consistently performs significantly better.
@@
 10 0.8 2.56 22.7 2.53 40.0 1.80 62.5
-C.5 Resutlts with OPT-30B and Llama-2-70B
+C.3 Resutlts with OPT-30B and Llama-2-70B
 Here we report the performances of different methods for OPT (350M and 30B) and Llama-2-Chat (7B and 70B).
@@
 […]
-C.6 Additional Experiments on CNNDM and Spider
-Table 14 and Table 15 reveal the downstream effectiveness of our method on CNNDM and Spider when the small model and large model are fine-tuned on the dataset.
-We can see that MTAD is consistently more effective and has a faster decoding speed.
-Table 14: Comparison of ROUGE-L Scores and Tokens per Second under Different Fine-Tuning Conditions on CNNDM
-No Fine-Tune Fine-Tune 68M Fine-Tune Both
-Method ROUGE-L Tokens/sec ROUGE-L Tokens/sec ROUGE-L Tokens/sec
-SpeD 0.114 37.7 0.114 20.4 0.164 24.3
-MTAD 0.118 44.2 0.121 25.0 0.168 27.1
-Table 15: Comparison of Execution Accuracy (EA) and Tokens per Second under Different Fine-Tuning Conditions on Spider
-No Fine-Tune Fine-Tune 68M Fine-Tune Both
-Method EA Tokens/sec EA Tokens/sec EA Tokens/sec
-SpD 11.5 28.5 11.5 27.8 16.3 25.6
-MTAD 13.0 30.3 14.8 32.3 18.3 29.4
-C.7 Visualization of perplexity and output quality
+C.4 Visualization of perplexity and output quality
 To further illustrate the relationship between perplexity and downstream performance, we present a scatter plot in Figure 7.
@@
 Figure 7: Relationship between relative perplexity (normalized by multinomial sampling’s perplexity) and relative performance score (normalized by multinomial sampling’s score).
-C.8 Correlation between Energy and Speed
-Figure 5 shows there is a correlation between speed and energy, whether we consider the entire table or focus on a specific dataset and model.
-For fairness, all methods for a given dataset and model were run on the same machine nodes.
-However, for a fixed method (e.g., Spectr), experiments on different datasets and models might be conducted on different nodes (all equipped with L40 GPUs).
-We did notice that the same configuration run on different machines may have varied energy consumption.
-This variation introduces some randomness, which could make the correlation appear less consistent across datasets and models.
-(a) group by dataset and model
-(b) group by methods
-Figure 5: Correlation between speed and energy
 Appendix D Energy Consumption Measurement
@@
 For MTAD, we choose the beam width from $\{4,8\}$, the number of draft tokens from $\{3,4\}$, and the acceptance threshold from $\{0.1,0.3,0.5,0.7,0.9\}$.
-Appendix G License of Datasets and Models
-Datasets:
-  • ChatGPT-Prompts: Non (https://huggingface.co/datasets/MohamedRashad/ChatGPT-prompts)
-  • ChatAlpaca: Apache-2.0 License
-  • CNN Dailymail: Apache-2.0 License
-  • MTBench: Apache-2.0 License
-  • Spider: CC BY-SA 4.0 License
-  • HumanEval: MIT License
-Models
-Codes